Jie Wu

Chair and Laura H. Carnell Professor Director of the Center for Networked Computing Department of Computer and Information Sciences Temple University Philadelphia, PA 19122 Office: (215)204-8450 Fax: (215)204-5082 Email: jiewu@temple.edu http://www.cis.temple.edu/˜ wu

1 Employment History

• Chairman and Laura H. Carnell Professor, Temple University 7/12 - present

• Chairman and Professor, Temple University 7/09 - 6/12

• Director of the Center for Networked Computing, Temple University 7/09 - present

• Program Director, National Science Foundation (NSF) 01/07 - 01/09

• Distinguished Professor, Florida Atlantic University 09/06 - 07/09

• Professor, Florida Atlantic University 08/98 - 08/06

• Associate Professor, Florida Atlantic University 08/94 - 07/98

• Assistant Professor, Florida Atlantic University 08/89 - 07/94

• Visiting Professor, Georgia Institute of Technology Summer 1991

• Lecturer, Shanghai University of Science and Technology 07/85 - 01/87

2 Statement of Professional Interests

Mobile Computing and Wireless Networks Routing Protocols Network Trust and Security Parallel Processing Distributed Algorithms Cloud and Green Computing Health-Care IT Social Network Applications

3 Educational Background

• Ph.D. in Computer Engineering, Florida Atlantic University, Boca Raton, FL, 08/89

• M.S. in Computer Science, Shanghai University of Science and Technology (now Shanghai University), Shanghai, P. R. of China, 07/85

• B.S. in Computer Engineering, Shanghai University of Science and Technology, Shanghai, P. R. of China, 07/82

4 Scholarly Achievements

4.1 Publications

4.1.1 Refereed Journals (277)

1. P. Ostovari, J. Wu, A. Khreishah, and N. Shroff, “Scalable Video Streaming with Helper Nodes using Random Linear Network Coding ,” accepted to appear in IEEE/ACM Transactions on Networking.

2. J. Zhang, X. Chen, Y. Xiang, W. Zhou, and J. Wu, “Robust Network Traffic Classification,” accepted to appear in IEEE/ACM Transactions on Networking.

3. Md. Z. Alam Bhuiyan, G. Wang, J. Cao, and J. Wu, “Deploying Wireless Sensor Networks with Fault-Tolerance for Structural Health Monitoring,” IEEE Transactions on Computers, Vol. 64, No. 2, 2015, 382-395.

4. S. Zhang, J. Wu, and S. Lu, “Collaborative Mobile Charging,” IEEE Transactions on Computers, Vol. 64, No. 3, 2015, 654-667.

5. D. Li and J. Wu, “On Data Center Network Architectures for Interconnecting Dual-Port Servers,” accepted to appear in IEEE Transactions on Computers.

6. G. Wang, W. Jiang, J. Wu, F. Li, and H. Zheng, “Trust Evaluation in Online Social Networks Using Generalized Network Flow,” accepted to appear in IEEE Transactions on Computers.

7. W. Zhang,Y. Lin, S. Xiao, J. Wu, and S.Zhou, “Privacy Preserving Ranked Multi-Keyword Search for Multiple Data Owners in Cloud Computing,” accepted to appear in IEEE Transactions on Computers.

8. H. Wang, Yong Li, P. Hui, D. Jin, and J. Wu, “Saving Energy in Partially Deployed Software Defined Networks ,” accepted to appear in IEEE Transactions on Computers.

9. D. Li and J. Wu, “Minimizing Energy Consumption for Frame-Based Tasks on Heterogeneous Multiprocessor Platforms,” IEEE Transactions on Parallel and Distributed Systems, Vol. 26, No. 3, 2015, 810-823.

10. Y. Dai, J. Wu, and C. Xin, “Efficient Virtual Backbone Construction without a Common Control Channel in Cognitive Radio Networks,” accepted to appear in IEEE Transactions on Parallel and Distributed Systems.

11. D. Tang, T. Li, J. Ren, and J. Wu, “Cost-Aware SEcure Routing (CASER) Protocol Design for Wireless Sensor Networks,” accepted to appear in IEEE Transactions on Parallel and Distributed Systems.

12. M. Xiao, J. Wu, and L. Huang, “Time-sensitive Utility-based Single-copy Routing in Low-duty-cycle Wireless Sensor Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 26, No. 5, 2015, 1452-1465.

13. A. Khreishah, I. M. Khalil, and J. Wu, “Universal Network Coding-based Opportunistic Routing for Unicast,” IEEE Transactions on Parallel and Distributed Systems, Vol. 26, No. 6, 2015, 1765-1773.

14. W. Chang and J. Wu, “Progressive or Conservative: Rationally Allocate Cooperative Work in Mobile Social Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 26, No. 7, 2015, 2020-2035.

15. S. Zhang, Z. Qian, Z. Luo, J. Wu, and S. Lu, “Burstiness-Aware Resource Reservation for Server Consolidation in Computing Clouds,” accepted to appear in IEEE Transactions on Parallel and Distributed Systems.

16. C. H. Liu, J. Fan, P. Hui, J. Wu, and K. K. Leung, “Towards QoI and Energy-Efficiency in Participatory Crowdsourcing,” accepted to appear in IEEE Transactions on Vehicular Technology.

17. Y. Dai, J. Wu, and Y. Zhao, “Boundary Helps: Reliable Route Selection With Directional Antennas in Cognitive Radio Networks,” accepted to appear in IEEE Transactions on Vehicular Technology.

18. H. Zhou, J. Chen, H. Zheng, and J. Wu, “Energy Efficiency and Contact Opportunities Trade-off in Opportunistic Mobile Networks,” accepted to appear in IEEE Transactions on Vehicular Technology.

19. L. Wu, X. Du, and J. Wu, “Effective Defense Schemes for Phishing Attacks on Mobile Computing Platforms,” accepted to appear in IEEE Transactions on Vehicular Technology.

20. M. Z. A Bhuiyan, G. Wang, J. Wu, J. Cao, X. Liu, and T. Wang, “ Dependable Structural Health Monitoring Using Wireless Sensor Networks,” accepted to appear in IEEE Transactions on Dependable and Secure Computing.

21. W. Jiang, J. Wu, and G. Wang, “On Selecting Recommenders for Trust Evaluation in Online Social Networks,” accepted to appear in ACM Transactions on Internet Technology.

22. E. Wang, Y. Yang, and J. Wu, “A Knapsack-based Buffer Management Strategy for Delay-tolerant Networks,” accepted to appear in Journal of Parallel and Distributed Computing.

23. Y. Shi, F. Zhang, J. Wu, and Z. Liu, “Randomized Oblivious Integral Routing for Minimizing Power Cost ,” accepted to appear in Theoretical Computer Science.

24. J. Medsger, A. Srinivasan, and J. Wu, “Information Theoretic and Statistical Drive Sanitization Models,” accepted to appear in the special issue of Journal of Privacy and Security.

25. S. Zhang, J. Wu, Z. Qian, and S. Lu, “MobiCache: Cellular traffic offloading leveraging cooperative caching in mobile social networks ,” accepted to appear in ELSEVIER Computer Networks.

26. S. Zhang, Z. Qian, J. Wu, and S. Lu, “Service-Oriented Resource Allocation in Clouds: Pursuing Flexibility and Efficiency, ” accepted to appear in Journal of Computer Science and Technology.

27. P. Ostovari, A. Khreishah, J. Wu, and W. Yang, “Trade-off between Redundancy and Feedback in Wireless Network Communication,” accepted to appear in Ad Hoc & Sensor Wireless Networks.

28. K. Li, H. Zheng, J. Wu, and X. Du, “Virtual Machine Placement in Cloud Systems through Migration Process,” accepted to appear in International Journal of Parallel, Emergent and Distributed Systems.

29. W. Wang, X. Chen, M. Lu, J. Wang, X. Zhang, and J. Wu, “Code Pruning in Opportunistic Routing through Bidirectional Coding Traffic Comparison,” accepted to appear in Wireless Communications and Mobile Computing.

30. Y. Dai and J. Wu, “Cooperation Scheme For Distributed Spectrum Sensing In Cognitive Radio Networks,” accepted to appear in Mobile Communications and Applications.

31. Y. Dai and J. Wu, “Pre-phase Improvement For Distributed Spectrum Sensing in Cognitive Radio Networks,” accepted to appear in Mobile Communications and Applications.

32. P. Ostovari, A. Khreishah, and J. Wu, “Broadcasting with Hard Deadlines in Wireless Multi-hop Networks Using Network Coding,” Wireless Communications and Mobile Computing Vol. 15, No. 5, 2015, 983-999.

33. K. Chen, G. Tan, M. Lu, and J. Wu, “CRSM: a practical crowdsourcing-based road surface monitoring system ,” accepted to appear in Wireless Networks.

34. J. Wu and Y. Wang, “Hypercube-based Multi-path Social Feature Routing in Human Contact Networks,” IEEE Transactions on Computers, Vol. 63, No. 2, 2014, 383-396.

35. W. Peng, F. Li, X. Zou, and J. Wu, “A Two-stage Deanonymization Attack Against Anonymized Social Networks,” IEEE Transactions on Computers, Vol. 63, No. 2, 2014, 290-303.

36. M. Xiao, J. Wu, and L. Huang, “Community-Aware Opportunistic Routing in Mobile Social Networks,” IEEE Transactions on Computers, Vol. 63, No. 7, 2014, 1682-1695.

37. C. Liu, Y. Pan, A. Chen, and J. Wu, “A DFA with Extended Character-set for Fast Deep Packet Inspection,” IEEE Transactions on Computers, Vol. 63, No. 8, 2014, 1925-1937.

38. Md. Z. Alam Bhuiyan, G. Wang, J. Cao, and J. Wu, “Deploying Wireless Sensor Networks with Fault-Tolerance for Structural Health Monitoring,” accepted to appear in IEEE Transactions on Computers.

39. S. Zhang, J. Wu, and S. Lu, “Collaborative Mobile Charging,” accepted to appear in IEEE Transactions on Computers.

40. S. Zhang, Z. Qian, J. Wu, S. Lu, and L. Epstein, “Virtual Network Embedding with Opportunistic Resource Sharing,” IEEE Transactions on Parallel and Distributed Systems, Vol. 25, No. 3, 2014, 816-827.

41. Y. Zhao and J. Wu, “The Design and Evaluation of An Information Sharing System for Human Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 25, No. 3, 2014, 796-805.

42. W. Peng, F. Li, X. Zou, and J. Wu, “Behavioral Malware Detection in Delay Tolerant Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 25, No. 1, 2014, 53-63.

43. G. Wang, Md. Z. Alam Bhuiyan, J. Cao, and J. Wu, “Detecting Movements of a Target Using Face Tracking in Wireless Sensor Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 25, No. 4, 2014, 939-949.

44. Y. Dai, J. Wu, and C. Xin, “Efficient Virtual Backbone Construction without a Common Control Channel in Cognitive Radio Networks,” accepted to appear in IEEE Transactions on Parallel and Distributed Systems.

45. J. Li, Y. Li, J. Ren, and J. Wu, “Hop-by-Hop Message Authentication and Source Privacy in Wireless Sensor Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 25, No. 5, 2014, 1223-1232.

46. Q. Liu, C. C. Tan, J. Wu, and G. Wang, “Towards Differential Query Services in Cost-Efficient Clouds,” IEEE Transactions on Parallel and Distributed Systems, Vol. 25, No. 6, 2014, 1648-1658.

47. G. Wang, W. Jiang, J. Wu, and Z. Xiong, “Fine-grained Feature-based Social Influence Evaluation in Online Social Networks,” accepted to appear in IEEE Transactions on Parallel and Distributed Systems.

48. G. Tan, M. Lu, F. Jiang, K. Chen, X. Huang, and J. Wu, “Bumping: A Bump-Aided Inertial Navigation Method for Indoor Vehicles Using Smartphones,” IEEE Transactions on Parallel and Distributed Systems, Vol. 25, No. 7, 2014, 1670-1680.

49. D. Tang, T. Li, J. Ren, and J. Wu, “Cost-Aware SEcure Routing (CASER) Protocol Design for Wireless Sensor Networks,” accepted to appear in IEEE Transactions on Parallel and Distributed Systems.

50. M. Xiao, J. Wu, and L. Huang, “Time-sensitive Utility-based Single-copy Routing in Low-duty-cycle Wireless Sensor Networks,” accepted to appear in IEEE Transactions on Parallel and Distributed Systems.

51. A. Khreishah, I. M. Khalil, and J. Wu, “Universal Network Coding-based Opportunistic Routing for Unicast,” accepted to appear in IEEE Transactions on Parallel and Distributed Systems.

52. W. Chang and J. Wu, “Progressive or Conservative: Rationally Allocate Cooperative Work in Mobile Social Networks,” accepted to appear in IEEE Transactions on Parallel and Distributed Systems.

53. Md. Z. Alam Bhuiyan, G. Wang, J. Cao, and J. Wu, “Sensor Placement with Multiple Objectives for Structural Health Monitoring,” ACM Transactions on Sensor Networks, Vol. 10, No. 4, 2014, 68:1-68:45.

54. J. Zhang, X. Chen, Y. Xiang, W. Zhou, and J. Wu, “Robust Network Traffic Classification,” accepted to appear in IEEE/ACM Transactions on Networking.

55. Q. Liu, G. Wang, and J. Wu, “Consistency as a Service: Auditing Cloud Consistency,” IEEE Transactions on Network and Service Management, Vol. 11, No. 1, 2014, 25-35.

56. Z. Song, C. H. Liu, J. Wu, J. Ma, and W. Wang, “QoI-Aware Multi-Task-Oriented Dynamic Participant Selection with Budget Constraints,” accepted to appear in IEEE Transactions on Vehicular Technology.

57. Y. Dai, J. Wu, and Y. Zhao, “Boundary Helps: Reliable Route Selection With Directional Antennas in Cognitive Radio Networks,” accepted to appear in IEEE Transactions on Vehicular Technology.

58. C. H. Liu, J. Fan, P. Hui, J. Wu, and K. K. Leung, “Towards QoI and Energy-Efficiency in Participatory Crowdsourcing,” accepted to appear in IEEE Transactions on Vehicular Technology.

59. P. Ostovari, J. Wu, and A. Khreishah, “Symbol-Level Reliable Broadcasting of Sensitive Data in Error-Prone Wireless Networks,” Journal of Parallel and Distributed Computing, Vol. 74, No. 7, 2014, 2673-2685.

60. Z. Qian, S. Zhang, Z. Luo, J. Wu, and S. Lu, “Burstiness-Aware Resource Reservation for Server Consolidation in Computing Clouds,” accepted to appear in IEEE Transactions on Parallel and Distributed Systems.

61. J. Wu, “Collaborative Mobile Charging and Coverage,” Journal of Computer Science and Technology, Vol. 29, No. 4, 2014, 550-561.

62. W. Jiang, G. Wang, and J. Wu, “Generating Trusted Graphs for Trust Evaluation in Online Social Networks,” Future Generation Computer Systems, Vol. 31, 2014, 48-58.

63. W. Chang, J. Wu, and C. C. Tan, “Improving Cooperative Trajectory Mapping Applications with Encounter-based Error Correction,” International Journal of Parallel, Emergent and Distributed Systems, Vol. 29, No. 1, 2014, 68-89.

64. Q. Liu, G. Wang, and J. Wu, “Time-Based Proxy Re-encryption Scheme for Secure Data Sharing in a Cloud Environment,” Information Sciences, Vol. 258, 2014, 355-370.

65. Y. Mao and J. Wu, “GFG Assisted Human Tracking using Smart Phones,” Ad Hoc & Sensor Wireless Networks, Vol. 21, No. 3-4, 2014, 259-281.

66. K. Li, J. Wu, and A. Blaisse, “Elasticity-aware Virtual Machine Placement in K-ary Data Centers,” the Parallel & Cloud Computing (PCC), Vol. 3 Iss. 2, 2014.

67. H. Zheng, K. Li, A. Blaisse, CC. Tan, and J. Wu, “Cloud CPU Verification Scheme for Individual End Users”, International Journal of Cloud Computing and Services, Vol. 6, Iss. 2, 2014, 377-390.

68. Y. Zhao, J. Wu, and C. Liu, “On Peer-Assisted Data Dissemination in Data Center Networks: Analysis and Implementation,” Tsinghua Science And Technology, Vol. 19, No. 1, 2014, 51-64.

69. Y. Zhao, J. Wu, and C. Liu, “Dache: A Data Aware Caching for Big-Data Applications Using the MapReduce Framework,” Tsinghua Science And Technology, Vol. 19, No. 1, 2014, 39-50.

70. C. Liu and J. Wu, “Fast Deep Packet Inspection with a Dual Finite Automata,” IEEE Transactions on Computers, Vol. 62, No. 2, 2013, 310-321.

71. H. Lai, Y. Pan, C. Liu, L. Lin, and J. Wu, “Sparse Learning-to-Rank via an Efficient Primal-Dual Algorithm,” IEEE Transactions on Computers, Vol. 62, No. 6, 2013, 1221-1233.

72. Y. Wang, W. -S. Yang, and J. Wu, “Analysis of a Hypercube-based Social Feature Multi-Path Routing in Delay Tolerant Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 24, No. 9, 2013, 1706-1716.

73. J. Fan, J. Chen, Y. Du, W. Gao, J. Wu, and Y. Sun, “Geo-Community-Based Broadcasting for Data Dissemination in Mobile Social Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 24, No. 4, 2013, 734-743.

74. D. Guo, H. Jin, T. Chen, J. Wu, L. Lu, D. Li, X. Zhou, “Partial Probing for Scaling Overlay Routing,” IEEE Transactions on Parallel and Distributed Systems, Vol. 24, No. 11, 2013, 2261-2272.

75. A. Khreishah, I. M. Khalil, and J. Wu, “Low Complexity and Provably Efficient Algorithm for Joint Inter and Intrasession Network Coding in Wireless Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 24, No. 10, 2013, 2015-2024.

76. F. Xu, X. Zhu, C. C. Tan, Q. Li, G. Yan, and J. Wu, “SmartAssoc: Decentralized Access Point Selection Algorithm to Improve Throughput,” IEEE Transactions on Parallel and Distributed Systems, Vol. 24, No. 12, 2013, 2482-2491.

77. X. Chen, Z. Dai, W. Li, Y. Hu, J. Wu, H. Shi, and S. Lu, “ProHet: A Probabilistic Routing Protocol with Assured Delivery Rate in Wireless Heterogeneous Sensor Networks,” IEEE Transactions on Wireless Communications, Vol. 12, No. 4, 2013, 1524-1531.

78. Y. Wang, J. Wu, and W. -S. Yang, “Cloud-based Multicasting with Feedback in Mobile Social Networks,” IEEE Transactions on Wireless Communications, Vol. 12, No. 12, 2013, 6043-6053.

79. X. Wang, Z. Li, and J. Wu, “Joint TCP Congestion Control and CSMA Scheduling without Message Passing,” IEEE Transactions on Wireless Communications, Vol. 12, No. 12, 2013, 6194-6204.

80. Y. Zhao and J. Wu, “Building a Reliable and High-performance Content-based Publish/subscribe System,” Journal of Parallel and Distributed Computing, Vol. 73, No. 4, 2013, 371-382.

81. B. Zhou, J. Wang, S. Li, Y. Cheng, and J. Wu, “A Continuous Secure Scheme in Static Heterogeneous Sensor Networks,” IEEE Communications Letters, Vol. 17, No. 9, 2013, 1868-1871.

82. J. Wu, Y. Dai, and Y. Zhao, “Effective Channel Assignments in Cognitive Radio Networks,” Computer Communications, Vol. 36, No. 4, 2013, 411-420.

83. G. Chen, Y. Wu, J. Wu, and W. Zheng, “TopCluster: A hybrid cluster model to support dynamic deployment in grid,” Journal of Computer and System Sciences, Vol. 79, No. 2, 2013, 201-215.

84. Y. Wang and J. Wu, “Ticket-based Multiple Packet Broadcasting in Delay Tolerant Networks,” Ad Hoc & Sensor Wireless Networks, Vol. 19, No. 3-4, 2013, 171-188.

85. X. Li, Z. Qian, S. Lu, and J. Wu, “Energy Efficient Virtual Machine Placement Algorithm with Balanced and Improved Resource Utilization in Data Center,” Mathematical and Computer Modeling, Vol. 58, No. 5-6, 2013, 1222-1235.

86. D. Li and J. Wu, “Energy-Aware Scheduling for Acyclic Synchronous Data Flows on Multiprocessors,” Journal of Interconnection Networks, Vol. 14, No. 3, 2013.

87. C. C. Tan, M. Korostelev, L. Bai, D. S. Mastrogiannis, and J. Wu, “Securing Remote Obstetrics Monitoring Systems,” International Journal on E-Health and Medical Communications, Vol. 4, Iss. 4, 2013, 34-47.

88. A. Srinivasan, S. Kolli, and J. Wu, “Steganographic Information Hiding that Exploits a Novel File System Vulnerability,” IJSN, Vol. 8, No. 2, 2013, 82-93.

89. F. Qi, Z. Tang, G. Wang, and J. Wu, “User Requirements-Based Security Ranking in SSL Protocol,” The Journal of Supercomputing, Vol. 64, No. 3, 2013, 762-776.

90. Q. Liu, C. C. Tan, J. Wu, and G. Wang, “Cooperative Private Searching in Clouds,” Journal of Parallel and Distributed Computing, Vol. 72, No. 8, 2012, 1019-1031. (Top-25 most downloaded papers at JPDC)

91. Q. Yuan, I. Cardei, and J. Wu, “An Efficient Prediction-based Routing in Disruption-Tolerant Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 23, No. 1, 2012, 19-31.

92. Y. Li, J. Ren, and J. Wu, “Quantitative Measurement and Design of Source-Location Privacy Schemes for Wireless Sensor Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 23, No. 7, 2012, 1302-1311.

93. C. Liu and J. Wu, “On Multi-copy Opportunistic Forwarding Protocols in Nondeterministic Delay Tolerant Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 23, No. 6, 2012, 1121-1128.

94. Y. Zhao, J. Wu, F. Li, and S. Lu, “On Maximizing the Lifetime of Wireless Sensor Networks Using Virtual Backbone Scheduling,” IEEE Transactions on Parallel and Distributed Systems, Vol. 23, No. 8, 2012, 1528-1535.

95. Y. Wang and J. Wu, “A Dynamic Multicast Tree based Routing Scheme without Replication in Delay Tolerant Networks,” Journal of Parallel and Distributed Computing, Vol. 72, No. 3, 2012, 424-436.

96. J. Y. Shi, M. Taifi, A. Khreishah, and J. Wu, “Tuple Switching Network – When Slower May Be Better,” Journal of Parallel and Distributed Computing, Vol. 72, No. 11, 2012, 1521-1534.

97. A. Khreishah, I. M. Khalil, P. Ostovari, and J. Wu, “Flow-based XOR Network Coding for Lossy Wireless Networks,” IEEE Transactions on Wireless Communications, Vol. 11, No. 6, 2012, 2321-2329.

98. X. Chen, N. C. Rowe, J. Wu and K. Xiong, “Improving the Localization Accuracy of Targets by Using Their Spatial-Temporal Relationships in Wireless Sensor Networks,” Journal of Parallel and Distributed Computing, Vol. 72, No. 8, 2012, 1008-1018.

99. X. Li, J. Wu, S. Lin, and X. Du, “Channel Switching Control Policy for Wireless Mesh Networks,” Journal of Parallel and Distributed Computing, Vol. 72, No. 10, 2012, 1295-1305.

100. Q. Liu, G. Wang, and J. Wu, “Secure and Privacy Preserving Keyword Search for Cloud Storage,” Journal of Network and Computer Applications, Vol. 35, No. 3, 2012, 927-933. (Top-25 most downloaded paper in JNCA)

101. F. Li, A. Srinivasan, and J. Wu, “Uncertainty Mitigation for Trustworthiness-Oriented Applications in Wireless Ad Hoc Networks,” IEICE Transactions, Vol. E95-D, No. 1, 2012, 12-19.

102. Z. Cheng, L. Bai, D. Choudhury, S. Biswas, and J. Wu, “Mobile Sensor Deployment and Coverage Using Multi-Agent-based Collective Formation Schemes,” International Journal of Performability Engineering, Vol. 8, No. 2, 2012, 141-150.

103. W. Chang, J. Wu, and C. C. Tan, “Friendship-based Location Privacy in Mobile Social Networks,” International Journal of Security and Networks, Vol. 6, No. 4, 2012, 226-236.

104. W. Chang, J. Wu, and C. C. Tan, “Wormhole Defense for Cooperative Trajectory Mapping,” International Journal of Parallel, Emergent and Distributed Systems (IJPEDS), Vol. 27, No. 5, 2012, 459-480.

105. J. C. Huang, X. Q. Li, and J. Wu, “A Semantic Searching Scheme in Heterogeneous Unstructured P2P Networks,” Journal of Computer Science and Technology, Vol. 26, No. 6, 2011, 925-941.

106. Z. Jiang, Z. Li, J. Wu, and N. Xiao, “Capability Information: A Cost-Effective Information Model for Multi-Hop Routing of Wireless Ad Hoc Networks in the Real Environment,” Journal of Parallel and Distributed Computing, Vol. 71, No. 8, 2011, 1085-1097.

107. C. Liu and J. Wu, “Practical Routing in a Cyclic MobiSpace,” IEEE/ACM Transactions on Networking, Vol. 19, No 2, 2011, 369-382.

108. Z. Li, J. Wu, J. Xie, T. Zhang, G. Chen, and Y. Dai, “Stability-Optimal Grouping Strategy of Peerto-Peer Systems,” IEEE Transactions on Parallel and Distributed Systems, Vol. 22, No. 12, 2011, 2079-2087.

109. L. Xie, Q. Li, W. Mao, J. Wu, and D. Chen, “Association Control for Vehicular WiFi Access: Pursuing Efficiency and Fairness,” IEEE Transactions on Parallel and Distributed Systems, Vol. 22, No. 8, 2011, 1323-1331.

110. D. Guo, J. Wu, Y. Liu, H. Jin, H. Chen, and T. Chen, “Quasi-Kautz Digraph for Peer-to-Peer Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 22, No. 6, 2011, 1042-1055.

111. G. Wang, Q. Liu, J. Wu, and M. Guo, “Hierarchical Attribute-Based Encryption and Scalable User Revocation for Sharing Data in Cloud Servers,” Computers & Security, Vol. 30, No. 5, 2011, 320-331.

112. Y. Xiao, M. Peng, H. Chen, S. Ozdemir, A.V. Vasilakos, and J. Wu, “Impacts of Sensor Node Distributions on Coverage in Sensor Networks,” Journal of Parallel and Distributed Computing, Vol. 71, No. 12, 2011, 1578-1591.

113. G. Wang and J. Wu, “FlowTrust: Trust Inference with Network Flows,” Frontiers of Computer Science in China, Vol. 5, No. 2, 2011, 181-194.

114. G. Wang, Q. Liu, and J. Wu, “Achieving Fine-Grained Access Control for Secure Data Sharing on Cloud Servers,” Concurrency and Computation: Practice and Experience, Vol. 23, No. 12, 2011, 1443-1464.

115. X. Chen, Q. Gu, Z. Jiang, J. Wu, “Fast Mobility Control Protocols with Sink Location Protection in Wireless Sensor Networks,” International Journal of Internet Protocol Technology (IJIPT), Vol. 6, No. 1/2, 2011, 3-17.

116. M. Lu and J. Wu, “Localized access point selection in infrastructure wireless LANs with performance guarantee,” Wireless Communications and Mobile Computing, Vol. 11. No. 8, 2011, 994-1010.

117. Y. Wang, X. Li, and J. Wu, “Delegation Forwarding in Delay Tolerant Networks Multicasting,” Journal of Communications (JCM), Vol. 6, No. 5, 2011, 384-392.

118. G. Wang and J. Wu, “Multi-Dimensional Evidence-Based Trust Management with Multi-Trusted Paths,” Future Generation Computer Systems, Vol. 27, No. 5, 2011, 529-538.

119. A. Srinivasan , F. Li, and J. Wu, “Secure Reputation Monitoring System - A Novel Connected Dominating Set-based Framework for WSNs,” Security and Communication Networks, Vol. 4, No. 1, 2011, 95-108.

120. I. Jawhar, J. Wu, and D. P. Agrawal, “Resource Scheduling in Wireless Networks Using Directional Antennas,” IEEE Transactions on Parallel and Distributed Systems, Vol. 21, No. 9, 2010, 1240-1253.

121. F. Li and J. Wu, “Uncertainty Modeling and Reduction in MANETs,” IEEE Transactions on Mobile Computing, Vol. 9, No. 7, July 2010, 1035 - 1048.

122. F. Li and J. Wu, “Attack and Flee: Game Theory-based Analysis On Interactions Among Nodes in MANETs,” IEEE Transactions on Systems, Man, and Cybernetics, Vol. 40, No. 3, June 2010, 612 - 622.

123. S. Yang, and J. Wu, “Efficient Broadcasting Using Network Coding and Directional Antennas in MANETs,” IEEE Transactions on Parallel and Distributed Systems, Vol. 21, No. 2, February 2010, 148-161.

124. D. Guo, J. Wu, H. Chen, Y. Yuan, and X. Luo, “The Dynamic Bloom Filters,” IEEE Transactions on Knowledge and Data Engineering, Vol. 22, No. 1, January 2010, 120-133.

125. S. Peng, W. Jia, G. Wang, J. Wu, and M. Guo, “Trusted Routing Based on Dynamic Trust Mechanism in Mobile Ad-Hoc Networks,” IEICE Transactions on Information and Systems, Vol 93-D, No. 3, pp. 510-517, March 2010.

126. J. Ren and J. Wu, “Survey on Anonymous Communications in Computer Networks,” Computer Communications, Vol. 33, No. 4, March 2010, 420-431.

127. Z. Jiang, J. Wu, and R. Kline, “Mobility Control for Achieving Optimal Configuration in Wireless Sensor Networks,” Wireless Networks, Vol. 16, No. 1, 2010, 141-152.

128. X. Chen, J. Shen, and J. Wu, “Improving Routing Protocol Performance in Delay Tolerant Networks Using Extended Information,” Journal of Systems and Software, Vol, 83, No. 8, 2010, 1301-1309.

129. X. Li, J. Xu, and J. Wu, “HR-SDBF: An Approach To Data-centric Routing in WSNs,” International Journal of High Performance Computing and Networking (IJHPCN), Vol. 6, No. 3/4, 2010, 181-196.

130. C. Liu and J. Wu, “Scalable Routing in Cyclic Mobile Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 20, No. 9, September 2009, 1325-1338.

131. C. Liu and J.Wu, “Virtual-Force-Based Geometric Routing Protocol in MANETs,” IEEE Transactions on Parallel and Distributed Systems, Vol. 20, No. 4, April 2009, 433-445.

132. M. Lu, F. Li, and J. Wu, “Efficient Opportunistic Routing in Utility-Based Ad Hoc Networks,” IEEE Transactions on Reliability, Vol. 58, No. 3, September 2009, 485-495.

133. I. Jawhar and J. Wu, “Resource Allocation for Real-Time and Multimedia Communication in TDMA-Based Wireless Networks,” International Journal of Ad Hoc and Ubiquitous Computing (IJAHUC), Vol. 4, No. 5, July 2009, 304-319.

134. M. Lu, J. Wu, M. Cardei, and M. Li, “Energy-Efficient Connected Coverage of Discrete Targets in Wireless Sensor Networks,” International Journal of Ad Hoc and Ubiquitous Computing, Vol. 4, No. 3/4, April 2009, 137-147.

135. B. Wu, J. Wu, and Y. Dong, “An Efficient Group Key Management Scheme for Mobile Ad Hoc Networks,” International Journal of Security and Networks (IJSN), Vol. 4, No. 1/2, February 2009, 125-134.

136. Q. Yuan, J. Wu, and I. Cardei, “SMRS: A Scalable Multi-Path Routing Scheme,” International Journal of Parallel, Emergent and Distributed Systems, Vol. 24, No. 1, February 2009, 69-84.

137. F. Qi, Z. Tang, G. Wang, and J. Wu, “SSL-Enabled Trusted Communication: Spoofing and Protecting the Non-Cautious Users,” Security and Communication Networks, Vol. 4, No. 4, 2009, 372-383.

138. S. Yang, J. Wu, and F. Dai, “Efficient Directional Network Backbone Construction in Mobile Ad Hoc Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 19, No. 12, Dec. 2008, 1601-1613.

139. J. Wu, S. Yang, and F. Dai, “Iterative Local Solutions for Connected Dominating Set in Ad Hoc Wireless Networks,” IEEE Transactions on Computers, Vol. 57, No. 5, May 2008, 702-715.

140. M. Cardei, S. Yang, and J. Wu, “Algorithms for Fault-Tolerant Topology in Heterogeneous Wireless Sensor Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 19, No. 4, April 2008, 545-558.

141. M. Lu, J. Yu and J. Wu, “Free-Riding on BitTorrent-like Peer-to-Peer File Sharing Systems: Modeling Analysis and Improvement,” IEEE Transactions on Parallel and Distributed Systems, Vol. 19, No. 7, July 2008, 954-966.

142. S. Yang, M. Li, and J. Wu, “Scan-Based Movement-Assisted Sensor Deployment Methods in Wireless Sensor Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 18, No. 8, August 2007, 1108-1121.

143. Z. Jiang, J. Wu, and D. Wang, “A New Fault-Information Model for Adaptive & Minimal Routing in 3-D Meshes,” IEEE Transactions on Reliability, Vol. 57, No. 1, 2008, 954-966.

144. V. Ponce, J. Wu, and X. Li, “Improve Peer Cooperation Using Social Networks,” International Journal of Parallel, Emergent and Distributed Systems (IJPEDS), Vol. 24, No. 3, 2009, 189-204.

145. A. Srinivasan, F. Li, J.Wu, and M. Li, “Clique-based Group Key Assignment in Wireless Sensor Networks,” International Journal of Security and Networks (IJSN), Vol. 3, No. 4, October 2008, 226-239.

146. F. Li, A. Srinivasan, and J. Wu, “PVFS: A Probabilistic Voting-based Filtering Scheme in Wireless Sensor Networks,” International Journal of Security and Networks (IJSN), Vol. 3, No. 3, August 2008, 173-182.

147. J. Wu and S. Yang, “Small World Model-based Polylogarithmic Routing using Mobile Nodes,” Journal of Computer Science and Technology, Vol. 23, No. 3, June 2008, 327-342.

148. X. Li, S. Zhong, and J. Wu, “ISRL: Intelligent Search by Reinforcement Learning in Unstructured Peer-to-Peer Networks,” International Journal of Parallel, Emergent, and Distributed Systems (I-JPEDS), Vol. 23, No. 1, February 2008, 17-44.

149. K. Ho, J. Wu, and J. Sum, “On the Session Lifetime Distribution of Gnutella,” International Journal of Parallel, Emergent and Distributed Systems (IJPEDS), Vol. 23, No. 1, February 2008, 1-15.

150. F. Dai and J. Wu, “Control Frame Shaping in Power Controlled and Directional MAC Protocols,” Ad Hoc & Sensor Wireless Networks: An International Journal, Vol. 5, No. 1-2, 2008, 1-26.

151. J. Wu, S. Yang, and F. Dai, “Logarithmic Store-Carry-Forward Routing in Mobile Ad Hoc Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 18, No. 6, June 2007, 735-748.

152. W. Lou and J. Wu, “Towards Broadcast Reliability in Mobile Ad Hoc Networks With Double Coverage,” IEEE Transactions on Mobile Computing, Vol. 6, No. 2, February 2007, 148-163.

153. A. Srinivasan and J. Wu, “Secure and Reliable Broadcasting in Wireless Sensor Networks using Multi-Parent Trees,” Security and Communication Networks, Vol. 2, No. 3, 2008, 239-253.

154. C. Li, G. Chen, M. Ye and Jie Wu, ”An Uneven Cluster-Based Routing Protocols for Wireless Sensor Networks,” Chinese Journal of Computers, Vol. 30, No. 1, January 2007, 27-36.

155. J. Sum, J. Wu, and C. -S. Leung, “On Profit Density-Based Greedy Algorithms for a Resource Allocation Problem in Web Services,” International Journal of Computer Applications, Vol. 29, No. 2, 2007, 155-163.

156. M. Ye, C. Li, G. H. Chen, and J. Wu, “An Energy Efficient Clustering Scheme in Wireless Sensor Networks,” International Journal of Ad Hoc & Sensor Wireless Networks, Vol. 3, No. 2-3, 2007, 99-199.

157. B. Wu, J. Wu, E. Fernandez, M. Ilyas, and S. Magliveras, “Secure and Efficient Key Management in Mobile Ad Hoc Networks,” Journal of Network and Computer Applications, Vol. 30, No. 3, 2007, 937-954. 937-954.

158. X. Li and J Wu, “A Hybrid Searching Scheme in Unstructured Peer-to-Peer Networks,” International Journal of Parallel, Emergent and Distributed Systems, Vol. 22, No. 1, 2007, 15-38.

159. B. Roussev and J. Wu, “Transforming Use Case Models to Class Models and OCL-Specifications,” International Journal of Computers and Applications, Vol. 29, No. 1, 2007, 59-69.

160. A. Srinivasan, J. Wu, and J. Teitelbaum, “Distributed Reputation-Based Secure Localization in Sensor Networks,” a special issue of Journal of Autonomic and Trusted Computing, 2007, 1-13.

161. G. Chen, C. Li, M. Ye, and J. Wu, “An Unequal Cluster-Based Routing Protocol in Wireless Sensor Networks,” Wireless Networks, Vol. 15, No. 2, 2007, 193-207.

162. J Sum, J. Wu, and K. Ho, “Analysis on a Localized Pruning Method for Connected Dominating Sets,” Journal of Information Science and Engineering (JISE), Vol. 23, No. 4, 2007, 1073-1086.

163. J. Wu and F. Dai, “Virtual Backbone Construction in MANETs using Adjustable Transmission Ranges,” IEEE Transactions on Mobile Computing, Vol. 5, No. 9, September 2006, 1188-1200.

164. J. Wu, M. Cardei, F. Dai, and S. Yang, “Extended Dominating Set and Its Applications in Ad Hoc Networks Using Cooperative Communication,” IEEE Transactions on Parallel and Distributed Systems, Vol. 17, No. 8, August 2006, 851-864.

165. J. Wu and F. Dai, “Mobility-Sensitive Topology Control in Mobile Ad Hoc Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 17, No. 6, June 2006, 522-535.

166. M. Cardei, J. Wu, and S. Yang, “Topology Control in Ad Hoc Wireless Networks using Cooperative Communication,” IEEE Transactions on Mobile Computing, Vol. 5, No. 6, June 2006, 711-724.

167. F. Dai and J. Wu, “Efficient Broadcasting in Ad Hoc Networks Using Directional Antennas,” IEEE Transactions on Parallel and Distributed Systems, Vol. 17, No. 4, April 2006, 335-347.

168. J. Wu, W. Lou, and F. Dai, “Extended Multipoint Relays to Determine Connected Dominating Sets in MANETs,” IEEE Transactions on Computers, Vol. 55, No. 3, March 2006, 334-347.

169. F. Dai and J. Wu, “On Constructing k-Connected k-Dominating Set in Wireless Ad Hoc and Sensor Networks,” Journal of Parallel and Distributed Computing (JPDC), Vol. 66, No. 7, 2006, 947-958.

170. M. Cardei, J. Wu, and M. Lu, “Improving Network Lifetime using Sensors with Adjustable Sensing Ranges,” International Journal of Sensor Networks (JSNet), Vol. 1, No. 1/2, 2006, 41-49.

171. S. Yang, F. Dai, M. Cardei, J. Wu, and F. Patterson, “On Connected Multiple Point Coverage in Wireless Sensor Networks,” International Journal of Wireless Information Networks, Vol. 13, No. 4, October 2006, 289-301.

172. B. Roussev, J. Wu, and L. Liu, “Distributed Computing Using Java: A Comparison of Two Server Designs,” Journal of Systems Architecture, Vol. 52, No. 7, July 2006, 432-440.

173. Z. Jiang and J. Wu, “A Limited-Global Information Model for Fault-Tolerant Routing in Dual-Cubes,” International Journal of Parallel, Emergent and Distributed Systems, Vol. 21, No. 1, February 2006, 61-77.

174. M. Cardei and J. Wu, “Energy-Efficient Coverage Problems in Wireless Ad-Hoc Sensor Networks,” Computer Communications Journal, Vol. 29, No. 4, February 2006, 413-420.

175. Y. Wang and J. Wu, “Label Routing Protocol: A New Cross-layer Protocol for Multi-Hop Ad Hoc Wireless Networks,” Journal of Pervasive Computing and Communications. Vol. 2, No. 4, January 2006, 362-369.

176. C. Yang, X. Li, and J. Wu, “Dominating-Set-Based Searching in Peer-to-Peer Networks,” International Journal of High Performance Computing and Networking (IJHPCN), Vol. 3, No. 4, 2005, 205-210.

177. M. Agarwal, J. H. Cho, L. X. Gao, and J. Wu, “Energy Efficient Broadcast in Wireless Ad Hoc Networks with Hitch-Hiking,” ACM/Kluwer MONET, a special issue on Mobile Network and Applications, Vol. 10, No. 6, December 2005, 897-910.

178. I. Jawhar and J. Wu, “QoS Support in TDMA-Based Mobile Ad Hoc Networks,” Journal of Computer Science and Technology, Vol. 20, No. 6, November 2005, 797-810.

179. J. Wu and Z. Jiang, “On Constructing the Minimum Orthogonal Convex Polygon for the Fault-tolerant Routing in 2-D Faulty Meshes,” IEEE Transactions on Reliability, Vol. 54, No. 3, September 2005, 449-458.

180. J. Wu, F. Gao, Z. Li, and Y. Min, “Optimal and Reliable Communication in Hypercubes using Extended Safety Vectors,” IEEE Transactions on Reliability, Vol. 54, No. 3, September 2005, 402-411.

181. F. Dai and J. Wu, “A Highly Reliable Multi-Path Routing Scheme for Ad Hoc Wireless Networks,” International Journal of Parallel, Emergent and Distributed Systems, Vol. 20, No. 3-4, September 2005, 205-219.

182. Q. Dai and J. Wu, “Computation of Minimal Uniform Transmission Range in Ad-Hoc Wireless Networks,” Cluster Computing, Vol. 8, No. 2-3, July 2005, 127-133.

183. J. Wu and F. Dai, “Efficient Broadcasting with Guaranteed Coverage in Mobile Ad-Hoc Networks,” IEEE Transactions on Mobile Computing, Vol. 4, No. 3, May/June 2005, 259-270.

184. J. Wu and D. Wang, “Fault-Tolerant and Deadlock-Free Routing in 2-D Meshes Using Rectilinear-Monotone Polygonal Fault Blocks,” International Journal of Parallel, Emergent and Distributed Systems, Vol. 20, No. 2, June 2005, 99-111.

185. I. Jawhar and J. Wu, “Race-Free Resource Allocation for QoS Support in Wireless Networks,” Ad Hoc & Sensor Wireless Networks: An International Journal, Vol. 1, No. 3, May 2005, 179-206.

186. J. Wu and S. Yang, “Energy-Efficient Node Scheduling Models in Sensor Networks with Adjustable Ranges,” International Journal of Foundations of Computer Science, Vol. 16, No. 1, February 2005, 3-17.

187. J. Wu and L. Sheng, “Deadlock-Free Multicasting in Irregular Networks Using Prefix Routing,” The Journal of Supercomputing, Vol. 31, No. 1, January 2005, 63-78.

188. F. Dai, Q. Dai, and J. Wu, “Power Efficient Routing Trees for Ad-Hoc Wireless Networks Using Directional Antenna,” International Journal of Ad-Hoc Networks, No. 3, 2005, 608-620.

189. F. Dai and J. Wu, “Performance Analysis of Broadcast Protocols in Ad Hoc Networks Based on Self-Pruning,” IEEE Transactions on Parallel and Distributed Systems, Vol. 15, No. 11, November 2004, 1027-1040.

190. J. Wu and F. Dai, “A Generic Distributed Broadcast Scheme in Ad Hoc Wireless Networks,” IEEE Transactions on Computers, Vol. 53, No. 10, October 2004, 1343-1354.

191. F. Dai and J. Wu, “An Extended Localized Algorithm for Connected Dominating Set Formation in Ad Hoc Wireless Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 15, No. 10, October 2004, 908-920.

192. J. Wu and F. Dai, “Mobility Control and Its Applications in Mobile Ad Hoc Networks,” IEEE Network, Vol. 18, No. 4, July/August 2004, 30-35.

193. C. -S. Leung, J. Sum, H. Shen, J. Wu, and G. H. Young, “Analysis and Design of an Agent Searching Algorithm for E-Marketplaces,” Journal of Cluster Computing, Vol. 7, No. 1, January 2004, 85-90.

194. J. Wu and Z. Jiang, “Extended Minimal Routing in 2D Meshes with Faulty Blocks,” International Journal of High Performance Computing and Networking (IJHPCN), Vol. 1, No. 1-3, Jan. 2004, 140-152.

195. J. Wu, “Optimal Broadcasting in Injured Hypercubes Using Directed Safety Levels,” Journal of Parallel and Distributed Computing, Vol. 63, No. 9, September 2003, 815-826.

196. J. Wu, “A Fault-Tolerant and Deadlock-Free Routing Protocol in 2-D Meshes Based on Odd-Even Turn Model,” IEEE Transactions on Computers, Vol 52, No. 9, September 2003, 1154-1169.

197. J. Wu, F. Dai, X. Lin, J. Cao, and W. Jia, “An Extended Fault-Tolerant Link-State Routing Protocol in the Internet,” IEEE Transactions on Computers. Vol. 52, No. 10, October 2003, 1298-1311.

198. Z. Jiang and J. Wu, “Fault-Tolerant Broadcasting in 2-D Wormhole-Routed Meshes,” The Journal of Supercomputing, Vol. 25, No.3, July 2003, 255-275.

199. J. Wu, B. Wu, and I. Stojmenovic, “Power-Aware Broadcasting and Activity Scheduling in Ad Hoc Wireless Networks Using Connected Dominating Sets,” Wireless Communications and Mobile Computing (WCMC), a special issue on Research in Ad Hoc Networking, Smart Sensing, and Pervasive Computing, Vol. 3, No. 4, June 2003, 425-438.

200. D. Xiang, A. Chen, and J. Wu, “Reliable Broadcasting in Wormhole-Routed Hypercube-Connected Networks using Local Safety Information,” IEEE Transactions on Reliability, Vol. 52, No. 2, June 2003, 245-256.

201. D. Xiang, A. Chen and J. Wu, “Local-Safety-Information-Based-Fault-Tolerant Broadcasting in Hypercubes,” Journal of Information Science and Engineering, Vol. 19, No. 3, May 2003, 467-478.

202. J. Wu and F. Dai, “Broadcasting in Ad Hoc Networks Based on Self-Pruning,” International Journal of Foundations of Computer Science, Vol. 14, No. 2, April 2003, 201-221.

203. J. Sum, H. Shen, J. Wu, C. -S. Leung, and G. Young, “Analysis on Extended Ant Routing Algorithms for Network Routing and Management,” The Journal of Supercomputing, Vol. 24, No. 3, March 2003, 327-340.

204. J. Wu and W. Lou, “Forward Node Set Based Broadcast in Clustered Mobile Ad Hoc Networks,” Wireless Communications and Mobile Computing (WCMC), a special issue on Algorithmic, Geometric, Graph, Combinatorial, and Vector Aspects of Wireless Networks and Mobile Computing, Vol. 3, No. 2, March 2003, 155-173.

205. J. Wu and F. Dai, “QoS Routing in Hypercube Multicomputers,” Parallel Processing Letters, Vol. 13, No. 1, March 2003, 43-52.

206. L. Sheng and J. Wu, “Maximum-Shortest-Path (MSP) is not Optimal for a General N × N Torus,” IEEE Transactions on Reliability, Vol. 52, No. 1, March 2003, 22-25.

207. J. Wu and S. Olariu, “On Cost-Optimal Merge of Two Intransitive Sorted Sequences,” International Journal of Foundations of Computer Science, Vol. 14. No. 1, February 2003, 99-106.

208. J. Wu, “A Simple Fault-Tolerant Adaptive and Minimal Routing Approach in 3-D Meshes,” Journal of Computer Science and Technology, Vol. 18, No. 1, January 2003, 1-13.

209. J. Wu, “An Extended Dynamic Source Routing Scheme in Ad Hoc Wireless Networks,” Telecommunication Systems, a special issue on Wireless Networks and Mobile Computing, Vol. 22, No. 1-4, January 2003, 61-75.

210. J. Wu and L. Sheng, “Deadlock-Free Routing in Regular Networks Using Prefix Routing,” Parallel Processing Letters, Vol. 13, No. 4, 2003, 705-720.

211. J. Shen, L. Sheng, and J. Wu, “Searching for Sorted Sequences of Kings in Tournaments,” SIAM Journal of Computing,Vol. 32, No. 5, 2003, 1201-1209.

212. J. Wu and F. Dai, “On Locality of Dominating Set in Ad Hoc Networks with Switch-On/Off Operations,” Journal of Interconnection Networks, a special issue on Algorithms for Networks, Vol. 3, No. 3 & 4, September & December 2002, 129-147.

213. J. Wu, “Extended Dominating-Set-Based Routing in Ad Hoc Wireless Networks with Unidirectional Links,” IEEE Transactions on Parallel and Distributed Systems, Vol. 13, No. 9, September 2002, 866-881.

214. J. Wu, “A Distributed Formulation of Smallest Faulty Orthogonal Convex Polygons in 2-D Meshes,” Journal of Parallel and Distributed Computing, Vol. 62, No. 7, July 2002, 1168-1185.

215. W. Lou and J. Wu, “On Reducing Broadcast Redundancy in Ad Hoc Wireless Networks,” IEEE Transactions on Mobile Computing, Vol. 1, No. 2, April-June 2002, 111-122.

216. S. Datta, I. Stojmenovic, and J. Wu, “Internal Node and Shortcut-Based Routing with Guaranteed Delivery in Wireless Networks,” Cluster Computing Journal, a special issue on Mobile Ad-hoc Networking, Vol. 5, No. 2, April 2002, 169-178.

217. J. Wu, F. Dai, M. Gao, and I. Stojmenovic, “On Calculating Power-Aware Connected Dominating Sets for Efficient Routing in Ad Hoc Wireless Networks,” Journal of Communications and Networks, a special issue on Innovations in Ad Hoc Mobile Pervasive Networks, Vol. 4, No. 1, March 2002, 59-70.

218. F. He and J. Wu, “An Efficient Parallel Implementation of the Everglades Landscape Fire Model Using Checkpointing,” Parallel Computing, Vol. 28, No. 1, January 2002, 65-82.

219. J. Wu, “A Fault-Tolerant Adaptive and Minimal Routing Scheme in n-D Meshes,” The Computer Journal, Vol. 45, No. 3, 2002, 349-363.

220. L. Sheng, Y. Wang, and J. Wu, “A Formula for the Optimal Shortest-Path Routing in 2-D Mesh-Connected Multicomputers,” Congressus Numerantium, Vol. 158, 2002, 131-141.

221. J. Wu and L. Sheng, “An Efficient Sorting Algorithm for a Sequence of Kings in a Tournament,” Information Processing Letters, Vol. 79, No. 6, September 2001, 297-299.

222. J. Wu and H. Li, “A Dominating-Set-Based Routing Scheme in Ad Hoc Wireless Networks,” Telecommunication Systems, a special issue on Wireless Networks, Vol. 18, No. 1-3, September 2001, 13-36.

223. J. Wu and Y. Yang “The Postal Network: A Recursive Network for Parameterized Communication Model,” Journal of Supercomputing, Vol. 19, No. 2, June 2001, 143-161.

224. D. Xiang, A. Chen, and J. Wu, “Local-Safety-Information-Based Broadcasting in Hypercube Multicomputers with Node and Link Faults,” Journal of Interconnection Networks, a special issue on Clustering Computing and Protocols, Vol. 2, No. 3, 2001, 365-378.

225. S. Olariu and J. Wu, “Guest Editors’ Introduction: Routing in Computer and Communication Networks,” Journal of Parallel and Distributed Computing, Vol. 61, No. 9, 2001, 1145-1147.

226. J. Wu and X. Chen, “A Fault-Tolerant Path-Based Multicasting in Mesh Networks,” Journal of Computer Science and Technology, Vol. 16, No. 4, 2001, 393-409.

227. B. Roussev and J. Wu, “Task Scheduling on NOWs using Lottery-Based Workstealing,” Annual Review of Scalable Computing , Vol. 3, 2001, 63-84.

228. L. Sheng and J. Wu, “A Note on ‘A Tight Lower Bound on the Number of Channels Required for Deadlock-Free Wormhole Routing’,” IEEE Transactions on Computers, Vol. 49, No. 9, September 2000, 1005.

229. S. Cang and J. Wu, “Time-Step Optimal Broadcasting in 3-D Meshes with Minimum Total Communication Distance,” Journal of Parallel and Distributed Computing, Vol. 60, No. 8, August 2000, 966-997.

230. S. Q. Zheng and J. Wu, “Dual of a Complete Graph as an Interconnection Network,” Journal of Parallel and Distributed Computing, Vol. 60, No. 8, August 2000, 1028-1046.

231. J. Wu, “Fault-Tolerant Adaptive and Minimal Routing in Mesh-Connected Multicomputers Using Extended Safety Levels,” IEEE Transactions on Parallel and Distributed Systems, Vol. 11, No. 2, February 2000, 149-159.

232. A. C. Wynn and J. Wu, “The Effect of Compression on Performance in a Demand Paging System,” Journal of Systems and Software, Vol. 50, No. 2, February 2000, 151-170.

233. E. B. Fernandez, J. Wu, and H. Qian, “A Combined Functional and Object-Oriented Approach to Software Design,” International Journal of Computers and Applications, Vol. 20, No. 2, 2000, 51-61.

234. J. Wu, “On Finding a Hamiltonian Path in a Tournament Using Semi-Heap,” Parallel Processing Letters, Vol. 10, No. 4, 2000, 279-294.

235. J. Wu and V. Rancov, “An Implementation of Context-Driven Parallel Computations,” International Journal of Computers and Applications, Vol. 22, No. 1, 2000, 29-37.

236. J. Wu and D. P. Agrawal, “Guest Editors’ Introduction: Challenges in Designing Fault-Tolerant Routing in Networks,” IEEE Transactions on Parallel and Distributed Systems, Vol. 10, No. 10, October 1999, 961-963.

237. J. Wu, “Maximum-Shortest-Path (MSP): An Optimal Routing Policy for Mesh-Connected Multicomputers,” IEEE Transactions on Reliability, Vol. 48, No. 3, September 1999, 247-255.

238. J. Wu, E. B. Fernandez, and Y. Luo, “Embedding of Binomial Trees in Hypercubes with Link Faults,” Journal of Parallel and Distributed Computing, Vol. 54, No. 1, October 1998, 59-74.

239. J. Wu and G. Guo, “Fault-Tolerance Measures for m-ary n-dimensional Hypercubes Based on Forbidden Faulty Sets,” IEEE Transactions on Computers, Vol. 47, No. 8, August 1998, 888-893.

240. J. Wu, “Adaptive Fault-Tolerant Routing in Cube-Based Multicomputers Using Safety Vectors,” IEEE Transactions on Parallel and Distributed Systems, Vol. 9, No. 4, April 1998, 321-334.

241. Y. Li, S. Q. Zheng, and J. Wu, “An Interconnection Network Based on the Dual of a Hypercube,” Informatica, a special issue on Optical Computing, Vol. 22, 1998, 499-508.

242. J. Wu, H. Huang, C. Fitz, Y. G. Wu, and F. Sklar, “A New Domain Partition Scheme for the Everglades Landscape Model (ELM),” International Journal of Computers and Applications, Vol. 20, No. 3, 1998, 137-146.

243. J. Wu, “Fault-Tolerant Communication in Cube-Based Multiple Bus Systems,” Parallel Processing Letters, a special issue on Computing on Bus-Based Architectures. Vol. 8, No. 3, 1998, 307-320.

244. J. Wu, E. B. Fernandez, and D. Dai, “Fault-Secure Scheduling,” The Computer Journal, Vol. 41, No. 4, 1998, 207-222.

245. J. Wu, “Extended Fibonacci Cubes,” IEEE Transactions on Parallel and Distributed Systems, Vol. 8, No. 12, December 1997, 1203-1210.

246. K. Huang and J. Wu, “Fault-Tolerant Resource Placement in Balanced Hypercubes,” Information Science, Vol. 99, No. 3-4, July 1997, 159-172.

247. J. Wu and K. Huang, “The Balanced Hypercube: A Cube-Based System for Fault-Tolerant Applications,” IEEE Transactions on Computers, Vol. 46, No. 4, April 1997, 484-490.

248. J. Wu, “Reliable Unicasting in Faulty Hypercubes Using Safety Levels,” IEEE Transactions on Computers, Vol. 46, No. 2, February 1997, 241-247.

249. J. Wu, “Dimension-Exchange-Based Global Load Balancing in Faulty Hypercubes,” Concurrency: Practice and Experience, Vol. 9, No. 1, January 1997, 41-61.

250. H. Zhuang, J. Wu, and W. Huang, “Optimal Planning of Robot Calibration Experiment Using Genetic Algorithms,” Journal of Robotic Systems, Vol. 14, No. 4, 1997, 741-752.

251. H. Qian and J. Wu, “Multitriangle: A New Planar-Type Interconnection Network,” International Journal of Foundations of Computer Science, a special issue on Interconnection Networks, Vol. 8, No. 2, 1997, 187-209.

252. J. Wu, E. B. Fernandez, and M. Zhang, “Design and Modeling of Hybrid Fault-Tolerant Software with Cost Constraints,” Journal of Systems and Software, Vol. 35, No. 2, November 1996, 141-149.

253. J. Wu, “Optimal Broadcasting in Hypercubes with Link Faults Using Limited Global Information,” Journal of Systems Architecture, Vol. 42, No. 5, November 1996, 367-380.

254. J. Wu “Reliable Communication on Cube-Based Multicomputers,” Journal of Computer Science and Technology, Vol. 11, No. 3, May 1996, 208-221.

255. H. Qian and J. Wu, “Enhanced Fibonacci Cubes,” The Computer Journal, Vol. 39, No. 4, 1996, 331-345.

256. J. Wu and K. Yao, “A Limited-Global-Information-Based Multicasting Scheme for Faulty Hypercube,” IEEE Transactions on Computers, Vol. 44, No. 9, September 1995, 1162-1166.

257. L. Brown and J. Wu, “Snooping in Fault-Tolerant Distributed Shared Memories,” Journal of Systems and Software, Vol. 29, No. 2, May 1995, 149-165.

258. J. Wu, “Safety levels-An Efficient Mechanism for Achieving Reliable Broadcasting in Hypercubes,” IEEE Transactions on Computers, Vol. 44, No. 5, May 1995, 702-706.

259. J. Wu and K. Huang, “The Balanced Hypercube Architecture for Fault-Tolerant Parallel Computation,” Programming and Computer Software, Vol. 1, 1995, 55-62.

260. J. Wu, “Tight Bounds on the Number of L-nodes in a Faulty Hypercube,” Parallel Processing Letters, Vol. 5, No. 2, 1995, 321-328.

261. J. Wu and K. Yao, “Reliability Optimization of Concurrent Software with Redundancy,” Informatica, Vol. 19, No. 3, 1995, 291-300.

262. K. Huang and J. Wu, “Area Efficient Layout of Balanced Hypercubes,” International Journal of High Speed Electronics and Systems, Vol. 6, No. 4, 1995.

263. J. Wu and E. B. Fernandez, “Using Petri Nets for the Design of Conversation Boundaries in Fault-Tolerant Software,” IEEE Transactions on Parallel and Distributed Systems, Vol. 5, No. 10, October 1994, 1106-1112.

264. J. Wu, Y. Wang, and E. B. Fernandez, “A Uniform Approach to Software and Hardware Fault Tolerance,” Journal of Systems and Software, Vol. 26, No. 2, August 1994, 117- 127.

265. J. Wu, “An Optimal Fault-Tolerant Nonredundant Broadcasting Scheme in Injured Hypercubes,” Journal of Parallel and Distributed Computing, Vol. 22, No. 2, August 1994, 295-313.

266. J. Wu and X. -H. Sun, “Optimal Cube-Connected Cube Multicomputers,” Journal on Microcomputer Applications, Vol. 17, No. 2, April 1994, 135-146.

267. J. Wu and T. Wu, “An Efficient Vector-Matrix-Vector Multiplication on Cube-Connected-Cubes Multicomputers,” International Journal of Mini and Microcomputers, Vol. 16, No. 3, 1994, 128-133.

268. J. Wu and E. B. Fernandez, “A Fault-Tolerant Distributed Broadcast Algorithm for Cube-Connected-Cycles,” Computer Systems and Engineering, Vol. 8, No. 4, October 1993, 224-233.

269. J. Wu and E. B. Fernandez, “Broadcasting in Faulty Hypercubes,” Microprocessing and Microprogramming, Vol. 39, 1993, 43-53.

270. J. Wu and M. Petrie, “Cube-Connected-Cubes Networks,” Microprocessing and Microprogramming, Vol. 33, No. 5, August 1992, 299-310.

271. J. Wu and E. B. Fernandez, “Comparison Connection Assignments for Diagnosis of Multiprocessor Systems Under a Two-Fault Assumption,” Computer Systems Science and Engineering, Vol. 7, No. 3, July 1992, 199-201.

272. J. Wu, E. B. Fernandez, and R. Zhang, “Some Extensions to the Lattice Model for Computer Security,” Computers and Security, Vol. 11, No. 4, 1992, 357-369.

273. J. Wu, “Fault-Tolerant Task Scheduling,” International Journal of Mini and Microcomputers, Vol. 13, No. 3, 1991, 135-139.

274. J. Wu and E. B. Fernandez, “A Simplification of A Conversation Design Scheme Using Petri Nets,” IEEE Transactions on Software Engineering, Vol. 15, No. 5, May 1989, 658-660.

275. J. Wu an G. Zhu, “A Program Synthesis System – SSS,” Journal of Computer Engineering, (in Chinese), Vol. 70, No. 3, August 1987, 18-28.

276. J. Wu an G. Zhu, “A Method of Achieving Several Goals Simultaneously,” Journal of Shanghai University of Science and Technology, (in Chinese), Vol. 20, No. 2, 1987, 29-35.

277. J. Wu, “A Necessary and Sufficient Condition for R = R^{+} and Its Generalization,” Journal of Shanghai University of Science and Technology, (in Chinese), Vol. 18, No. 2, 1985, 60-63.

4.1.2 Refereed Conference Proceedings (467)

1. M. Z. A Bhuiyan, G. Wang, J. Wu, X. Xing, and X. Liu, “Application-Oriented Sensor Network Architecture for Dependable Structural Health Monitoring,” Proc. of the 21st IEEE Pacific Rim International Symposium on Dependable Computing (IEEE PRDC 2015), November 18 - October 20, 2015.

2. Y. Dai, and J. Wu, “Integration of Spectrum Database and Sensing Results for Hybrid Spectrum Access System,” Proc. of the 12th International Conference on Mobile Ad hoc and Sensor Systems (MASS 2015), October 19 - October 22, 2015.

3. P. Ostovari, and J. Wu, “Reliable Broadcast with Joint Forward Error Correction and Erasure Codes In Wireless Communication Networks,” Proc. of the 12th International Conference on Mobile Ad hoc and Sensor Systems (MASS 2015), October 19 - October 22, 2015.

4. C. Song, J. Wu, L. Lu, and M. Liu, “RTS Assisted Mobile Localization: Mitigating Jigsaw Puzzle Problem of Fingerprint Space with Extra Mile,” Proc. of the 12th International Conference on Mobile Ad hoc and Sensor Systems (MASS 2015), October 19 - October 22, 2015.

5. H. Huang, S. Li, L. Chen, J. Gao, A. Mamat, and J. Wu, “Dynamic Mobile Charger Scheduling in Heterogeneous Wireless Sensor Networks,” Proc. of the 12th International Conference on Mobile Ad hoc and Sensor Systems (MASS 2015), October 19 - October 22, 2015.

6. D. Li, J. Wu, Z. Liu and F. Zhang, “Dual-Centric Data Center Network Architectures,” Proc. of the 44th International Conference on Parallel Processing (ICPP 2015), September 1 - September 4, 2015.

7. W. Chang and J. Wu, “SocialVoD: a Social Feature-based P2P System,” Proc. of the 44th International Conference on Parallel Processing (ICPP 2015), September 1 - September 4, 2015.

8. J. Li, J. Wu, X. Yang and S. Zhong, “Optimizing MapReduce based on Locality of K-V Pairs and Overlap between Shuffle and Local Reduce,” Proc. of the 44th International Conference on Parallel Processing (ICPP 2015), September 1 - September 4, 2015.

9. E. Wang, Y. Yang, J. Wu and W. Liu, “A Buffer Management Strategy on Spray and Wait Routing Protocol in DTNs,” Proc. of the 44th International Conference on Parallel Processing (ICPP 2015), September 1 - September 4, 2015.

10. H. Zheng and J. Wu, “Snowballing Effects in Preferential Attachment: The Impact of The Initial Links,” Proc. of the 24th IEEE International Conference on Computer Communications and Networks (IEEE ICCCN 2015), August 3 - August 6, 2015.

11. Z. Wan, J. Wu, and H. Zheng, “Utility-based Uploading Strategy in Cloud Scenarios,” Proc. of the 24th IEEE International Conference on Computer Communications and Networks (IEEE ICCCN 2015), August 3 - August 6, 2015.

12. D. Li and J. Wu, “FCell: Towards the Tradeoffs in Designing Data Center Network Architectures,” Proc. of the 24th IEEE International Conference on Computer Communications and Networks (IEEE ICCCN 2015), August 3 - August 6, 2015. (Best paper award)

13. P. Liu, B. Xu, Z. Jiang and J. Wu, “HAEP: Hospital Assignment for Emergency Patients in a Big City,” Proc. of the 24th IEEE International Conference on Computer Communications and Networks (IEEE ICCCN 2015), August 3 - August 6, 2015.

14. E. Wang, Y. Yang, and J. Wu, “Energy Efficient Phone-to-Phone Communication Based on WiFi Hotspot in PSN,” Proc. of the 24th IEEE International Conference on Computer Communications and Networks (IEEE ICCCN 2015), August 3 - August 6, 2015.

15. K. Sakai, M. Sun, W. Ku, J. Wu, and T. Lai, “Multi-Path-Based Avoidance Routing in Wireless Networks,” Proc. of the 35th IEEE International Conference on Distributed Computing Systems (IEEE ICDCS 2015), June 29 - July 2, 2015.

16. H. Zheng and J. Wu, “Optimizing Roadside Advertisement Dissemination in Vehicular Cyber-Physical Systems,” Proc. of the 35th IEEE International Conference on Distributed Computing Systems (IEEE ICDCS 2015), June 29 - July 2, 2015.

17. M. Xiao, H. Huang, L. Huang, W. Yang, and J. Wu, “Deadline-sensitive Opportunistic Utility-based Routing in Cyclic Mobile Social Networks,” Proc. of the 12th IEEE International Conference on Sensing, Communications, and Networking (IEEE SECON 2015), June 22 - June 25, 2015.

18. H. Zheng and J. Wu, “Data Collection and Event Detection in the Deep Sea with Delay Minimization,” Proc. of the 12th IEEE International Conference on Sensing, Communications, and Networking (IEEE SECON 2015), June 22 - June 25, 2015.

19. J. Chen, Q. Yuan, R. Du, and J. Wu, “MuCAR: A Greedy Multi-flow-based Coding-Aware Routing in Wireless Networks,” Proc. of the 12th IEEE International Conference on Sensing, Communications, and Networking (IEEE SECON 2015), June 22 - June 25, 2015.

20. P. Ostovari and J. Wu, “Reliable Videos Broadcast with Network Coding and Coordinated Multiple Access Points,” Proc. of the 14th IEEE International Symposium on a World of Wireless, Mobile and Multimedia Networks (IEEE WoWMoM 2015), June 14 - June 17, 2015.

21. S. Zhang, Z. Qian, F. Kong, J. Wu, and S. Lu, “P 3: Joint Optimization of Charger Placement and Power Allocation for Wireless Power Transfer ,” Proc. of the 34th IEEE International Conference on Computer Communications (IEEE INFOCOM 2015), April 26 - May 1, 2015.

22. Y. Zhao, W. Li, J. Wu, and S. Lu, “Quantized Conflict Graphs for Wireless Network Optimization,” Proc. of the 34th IEEE International Conference on Computer Communications (IEEE INFOCOM 2015), April 26 - May 1, 2015.

23. X. Liu, B. Xiao, K. Li, J. Wu, A. X. Liu, H. Qi, and X. Xie, “RFID Cardinality Estimation with Blocker Tags,” Proc. of the 34th IEEE International Conference on Computer Communications (IEEE INFOCOM 2015), April 26 - May 1, 2015.

24. M. Xiao, J. Wu, L. Huang, Y. Wang, and C. Liu, “Multi-Task Assignment for CrowdSensing in Mobile Social Networks,” Proc. of the 34th IEEE International Conference on Computer Communications (IEEE INFOCOM 2015), April 26 - May 1, 2015.

25. N. Wang, and J. Wu, “Trajectory Scheduling for Timely Data Report in Underwater Wireless Sensor Networks,” Proc. of the IEEE Global Communications Conference (GlobeCom 2015), December 6 - December 10, 2015.

26. W. Chang, and J. Wu, “Reliability Enhanced Social Crowdsourcing,” Proc. of the IEEE Global Communications Conference (GlobeCom 2015), December 6 - December 10, 2015.

27. Z. Wan, and J. Wu, “Utility-based VM Assignment in Cloud Scenarios,” Proc. of the IEEE Global Communications Conference (GlobeCom 2015), December 6 - December 10, 2015.

28. Q. Yuan, I. Cardei, J. Chen, and J. Wu, “Multi-copy Routing with Trajectory Prediction in Social Delay-Tolerant Networks,” Proc. of the IEEE Global Communications Conference (GlobeCom 2015), December 6 - December 10, 2015.

29. E. Wang, Y. Yang, J. Wu, W. Yang, and W. Liu, “A Lightweight Message Dissemination Strategy for Minimizing Delay in Online Social Networks,” Proc. of the IEEE Global Communications Conference (GlobeCom 2015), December 6 - December 10, 2015.

30. I. Jawhar, J. Wu, N. Mohamed, and S. Zhang, “An Efficient Graph Search Algorithm for Backbone Discovery In Wireless Linear Sensor Networks,” Proc. of the 4th IEEE Annual International Workshop on Mission-Oriented Wireless Sensor Networking (IEEE MASS MiSeNet 2015), October 19 - October 22, 2015.

31. C. Liu, J. Wu, L. Xie, S. Lu, and C. Wang, “FootStep-Tracker: An Anchor-Free Indoor Localization System via Sensing Foot Steps,” Proc. of the ACM International Joint Conference on Pervasive and Ubiquitous Computing (UbiComp 2015 Demo session), September 7 - September 11, 2015.

32. Q. Liu, G. Wang, J. Wu, and W. Chang, “User-Controlled Security Mechanism in Data-Centric Clouds,” Proc. of the 17th IEEE International Conference on High Performance Computing and Communications (HPCC 2015 Invited paper), August 24 - August 26, 2015.

33. H. Zheng and J. Wu, “DynFluid: Predicting Time-Evolving Rating in Recommendation Systems via Fluid Dynamics,” Proc. of the 14th IEEE International Conference on on Trust, Security and Privacy in Computing and Communications (IEEE TrustCom 2015), August 20 - August 22, 2015.

34. P. Ostovari, J. Wu, and A. Khreishah, “Cooperative Mobile Internet Access with Opportunistic Scheduling,” Proc. of the 3rd International Symposium on Resilient Communication Systems (ISRCS 2015), Augst 18 - Augst 20, 2015.

35. L. Wang, A. Anta, F. Zhang, J. Wu, and Z. Liu, “Multi-resource Energy-efficient Routing in Cloud Data Centers with Network-as-a-Service,” Proc. of the IEEE Symposium on Computers and Communications (ISCC 2015), July 6 - July 9, 2015.

36. E. Wang, Y. Yang, and J. Wu, “Dynamic Beaconing Control in Energy-constrained Delay Tolerant Networks,” Proc. of the International Conference on Communications (ICC 2015), June 8 - June 12, 2015.

37. A. Srinivasan, J. Wu, and J. Shi, “Android-Stego: A Novel Service Provider Imperceptible MMS Steganography Technique Robust to Message Loss,” Proc. of the 8th International Conference on Mobile Multimedia Communications (MOBIMEDIA 2015), May 25 - May 27, 2015.

38. B. Ji and J. Wu, “Node-based Scheduling with Provable Evacuation Time,” Proc. of the 49th Annual Conference on Information Sciences and Systems (CISS 2015), March, 2015.

39. P. Ostovari and J. Wu, “Robust Wireless Delivery of Scalable Videos using Inter-layer Network Coding,” Proc. of IEEE Wireless Communications and Networking Conference (IEEE WCNC 2015), March 9 - 12, 2015.

40. Y. Dai and J. Wu, “Opportunistic Routing Based Scheme with Multi-layer Relay Sets in Cognitive Radio Networks,” Proc. of IEEE Wireless Communications and Networking Conference (IEEE WCNC 2015), March 9 - 12, 2015.

41. E. Wang, Y. Yang, and J. Wu, “A Knapsack-based Message Scheduling and Drop Strategy for Delaytolerant Networks,” Proc. of the 12th European International Conference on Wireless Sensor Networks (EWSN 2015), February 9 - February 11, 2015.

42. D. Li, J. Wu, Z. Liu, and F. Zhang, “Joint Power Optimization Through VM Placement and Flow Scheduling in Data Centers,” Proc. of the 33rd International Performance Computing and Communications Conference (IPCCC 2014), December 5-7, 2014.

43. J. Wu, P. Sabatino, J. Tsan, and Z. Jiang, “An Infrastructure-less Vehicle Counting without Disruption,” Proc. of the 43rd International Conference on Parallel Processing (ICPP), September 9-12, 2014.

44. D. Li and J. Wu, “Energy-Aware Scheduling for Aperiodic Tasks on Multi-core Processors,” Proc. of the 43rd International Conference on Parallel Processing (ICPP), September 9-12, 2014.

45. D. Li and J. Wu, “On the Design and Analysis of Data Center Network architectures for Interconnecting Dual-Port Servers,” Proc. of the 33rd IEEE International Conference on Computer Communications (INFOCOM), April 27 - May 2, 2014.

46. X. Li, J. Wu, S. Tang, and S. Lu, “Let’s Stay Together: Towards Traffic Aware Virtual Machine Placement in Data Centers,” Proc. of the 33rd IEEE International Conference on Computer Communications (INFOCOM), April 27 - May 2, 2014.

47. W. Jiang, J. Wu, G. Wang, and H. Zheng, “FluidRating: A Time-Evolving Rating Scheme in Trustbased Recommendation Systems Using Fluid Dynamics,” Proc. of the 33rd IEEE International Conference on Computer Communications (INFOCOM), April 27 - May 2, 2014.

48. L. Xie, H. Han, Q. Li, J. Wu, and S. Lu, “Efficiently Collecting Histograms Over RFID Tags,” Proc. of the 33rd IEEE International Conference on Computer Communications (INFOCOM), April 27 - May 2, 2014.

49. C. Liu, Y. Pan, A. Chen, K. Bian, and J. Wu, “Scalable Opportunistic Forwarding Algorithms in Delay Tolerant Networks using Similarity Hashing,” Proc. of the 11th IEEE International Conference on Sensing, Communications and Networking (SECON), June 30 - July 3, 2014.

50. H. Zheng, Y. Wang, and J. Wu, “Optimizing Multi-copy Two-hop Routing in Mobile Social Networks,” Proc. of the 11th IEEE International Conference on Sensing, Communications and Networking (SECON), June 30 - July 3, 2014.

51. Y. Dai, J. Wu, and A. Daniels, “Channel Dynamics Matter: Forwarding Node Set Selection in Cognitive Radio Networks,” Proc. of the 15th International Symposium on a World of Wireless, Mobile and Multimedia Networks (WoWMoM), June 16-19, 2014.

52. H. Zheng and J. Wu, “Up-and-Down Routing in Mobile Opportunistic Social Networks with Bloom-Filter-Based Hints,” Proc. of IEEE/ACM International Symposium on Quality of Service (IWQoS), May 26-27, 2014.

53. Y. Wang and J. Wu, “Complex Contagions Models in Opportunistic Mobile Social Networks,” Proc. of the 23rd International Conference on Computer Communications and Networks (ICCCN), August 4-7, 2014.

54. Y. Dai and J. Wu, “Hitchhiking in Cognitive Radio Networks : Spectrum Sensing Assisted By Cores and Cluster,” Proc. of the 23rd International Conference on Computer Communications and Networks (ICCCN), August 4-7, 2014.

55. P. Ostovari and J. Wu, “The Benefits of Cooperation Between the Cloud and Private Data Centers for Multi-rate Video Streaming,” Proc. of the 23rd International Conference on Computer Communications and Networks (ICCCN), August 4-7, 2014.

56. S. Zhang, Z. Qian, J. Wu, and S. Lu, “Leveraging Tenant Flexibility in Resource Allocation for Virtual Networks,” Proc. of the 23rd International Conference on Computer Communications and Networks (ICCCN), August 4-7, 2014.

57. C. Liu, Y. Pan, J. Zhang, M. Xiao, A. Chen, and J. Wu, “A Utility-based Multi-copy Forwarding Algorithm in Delay Tolerant Networks,” Proc. of the 23rd International Conference on Computer Communications and Networks (ICCCN), August 4-7, 2014.

58. L. Wu, X. Du, and J. Wu, “MobiFish: A Lightweight Anti-Phishing Scheme for Mobile Phones,” Proc. of the 23rd International Conference on Computer Communications and Networks (ICCCN), August 4-7, 2014.

59. Z. Jiang, J. Wu, and P. Sabatino, “GUI: GPS-Less Traffic Congestion Avoidance in Urban Areas with Inter-Vehicular Communication,” Proc. of the 11th IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), October 28-30, 2014.

60. A. Gharaibeh, A. Khreishah, I. Khalil, and J. Wu, “Asymptotically-Optimal Incentive-Based En-Route Caching Scheme,” Proc. of the 11th IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), October 28-30, 2014.

61. H. Zheng and J. Wu, “Fast Information Cascade Prediction Through Spatiotemporal Decompositions,” Proc. of the 11th IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), October 28-30, 2014.

62. N. Wang and J. Wu, “A General Data and Acknowledgement Dissemination Scheme in Mobile Social Networks,” Proc. of the 11th IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), October 28-30, 2014.

63. J. Wu and H. Zheng, “On Efficient Data Collection and Event Detection with Delay Minimization in Deep Sea,” Proc. of the ACM MobiCom workshop on Challenged Networks (ACM CHANTS, in conjunction with ACM MobiCom), September 2014.

64. W. Chang and J. Wu, “Strengthening Data Privacy During Propagation,” Proc. of the ACM MobiCom workshop on Security and Privacy aspects of Mobile Environments (ACM SPME, in conjunction with ACM MobiCom), September 2014.

65. R. Beigel, J. Wu, and H. Zheng, “On Optimal Scheduling of Multiple Mobile Chargers in Wireless Sensor Networks,” Proc. of the 1st International Workshop on Mobile Sensing, Computing and Communication (MSCC), August 2014.

66. N. Wang and J. Wu, “InterestSpread: An Efficient Method for Content Transmission in Mobile Social Networks,” Proc. of the 1st International Workshop on Mobile Sensing, Computing and Communication (MSCC), August 2014.

67. S. Zhang, Z. Qian, J. Wu, and S. Lu, ”SEA: Stable Resource Allocation in Geographically Distributed Clouds,” Proc. of IEEE International Conference on Communications (ICC), June 2014.

68. R. Houlihan, X. Du, C. C. Tan, J. Wu, and M. Guizani, ”Auditing Cloud Service Level Agreement on VM CPU Speed,” Proc. of IEEE International Conference on Communications (ICC), June 2014.

69. K. Li, Z. Wan, J. Wu, and A. Blaisse, “Towards Makespan Minimization Task Allocation in Data Centers,” Proc. of the fifth International Conference on Cloud Computing, GRIDs, and Virtualization (CLOUD COMPUTING), May 25-29, 2014.

70. P. Ostovari, J. Wu, and Y. Dai, “Priority-Based Broadcasting of Sensitive Data in Error-Prone Wireless Networks,” Proc. of the 2nd International Symposium on Resilient Communication Systems (ISRCS), August 19-21, 2014.

71. Y. Wang, J. Wu, and M. Xiao, “Hierarchical Cooperative Caching in Mobile Opportunistic Social Networks,” Proc. of the IEEE Global Communications Conference (GLOBECOM), December 8-12, 2014.

72. C. Song, W. -S. Yang, J. Wu, and M. Liu, “Red or Green: Analyzing the Data Delivery with Traffic Lights in Vehicular Ad Hoc Networks,” Proc. of the IEEE Global Communications Conference (GLOBECOM), December 8-12, 2014.

73. S. Shao , A. Khreishah, M. B. Rahaim, H. Elgala, M. Ayyash, T. D. C. Little, and Jie Wu, “An Indoor Hybrid WiFi-VLC Internet Access System,” Proc. of Workshop on CellulAR Traffic Offloading to Opportunistic Networks (CARTOON), October 2014.

74. Y. He, C. Liu, Y. Pan, J. Zhang, J. Wu, Y. Zhao, S. Yang, and M. Lu, “A Localized Efficient Forwarding Algorithm in Large-scale Delay Tolerant Networks,” Proc. of the third Annual International Workshop on Mission-Oriented Wireless Sensor Networking (MiSeNet), October 2014.

75. Y. Wang, J. Wu, M. Xiao, and D. Zhang, “Heterogeneous Community-based Routing in Opportunistic Mobile Social Networks,” Proc. of the third Annual International Workshop on Mission-Oriented Wireless Sensor Networking (MiSeNet), October 2014.

76. I. Jawhar, X. Li, J. Wu, and N. Mohamed, “Backbone Discovery In Thick Wireless Linear Sensor Networks,” Proc. of the third Annual International Workshop on Mission-Oriented Wireless Sensor Networking (MiSeNet), October 2014.

77. W. Jiang and J. Wu, “Trust Models in Wireless Sensor Networks and Online Social Networks: A Comparative Study,” Proc. of the third Annual International Workshop on Mission-Oriented Wireless Sensor Networking (MiSeNet), October 2014.

78. J. Wu, M. Xiao, and L. Huang, “Homing Spread: Community Home-based Multi-copy Routing in Mobile Social Networks,” Proc. of the 32nd IEEE International Conference on Computer Communications (INFOCOM), April 2013.

79. Y. Dai, J. Wu, and C. Xin, “Virtual Backbone Construction for Cognitive Radio Networks without Common Control Channel,” Proc. of the 32nd IEEE International Conference on Computer Communications (INFOCOM), April 2013.

80. M. Xiao, J. Wu, C. Liu, and L. Huang, “TOUR: Time-sensitive Opportunistic Utility-based Routing in Delay Tolerant Networks,” Proc. of the 32nd IEEE International Conference on Computer Communications (INFOCOM), April 2013.

81. Q. Liu, G. Wang, F. Li, S. Yang, and J. Wu, “Outsourcing Privacy-Preserving Social Networks to a Cloud,” Proc. of the 32nd IEEE International Conference on Computer Communications (INFO-COM), April 2013.

82. Y. Zhao and J. Wu, “Dache: A Data Aware Caching for Big-Data Applications Using The MapReduce Framework,” Proc. of the 32nd IEEE International Conference on Computer Communications (INFOCOM) (Mini-Conference), April 2013.

83. S. Tang and J. Wu, “Qute: Quality-of-Monitoring Aware Sensing and Routing Strategy in Wireless Sensor Networks,” Proc. of the 14th ACM International Symposium on Mobile Ad Hoc Networking and Computing (MobiHoc), Jul 29 - Aug 1, 2013.

84. S. Zhang, J. Wu, and S. Lu, “Minimum Makespan Workload Dissemination in DTNs: Making Full Utilization of Computational Surplus Around,” Proc. of the 14th ACM International Symposium on Mobile Ad Hoc Networking and Computing (MobiHoc Poster), Jul 29 - Aug 1, 2013.

85. Md. Z. Alam Bhuiyan, G. Wang, J. Cao, and J. Wu, “Energy and Bandwidth-Efficient Wireless Sensor Networks for Monitoring High-Frequency Events,” Proc. of the 10th IEEE International Conference on Sensing, Communication, and Networking (SECON), June 2013.

86. P. Ostovari, J. Wu, and A. Khreishah, “Efficient Symbol-Level Transmission in Error-Prone Wireless Networks,” Proc. of the 14th International Symposium on a World of Wireless, Mobile and Multimedia Networks (WoWMoM), June 2013.

87. Y. Wang and J. Wu, “Social-Tie-Based Information Dissemination in Mobile Opportunistic Social Networks,” Proc. of the 14th International Symposium on a World of Wireless, Mobile and Multimedia Networks (WoWMoM), June 2013.

88. Y. Dai and J. Wu, “Whether and When to Share: Spectrum Sensing as An Evolutionary Game,” Proc. of the 8th International Conference on Cognitive Radio Oriented Wireless Networks (CROWNCOM), July 8-10, 2013.

89. Y. Dai and J. Wu, “Sense in Order: Channel Selection for Sensing in Cognitive Radio Networks,” Proc. of the 8th International Conference on Cognitive Radio Oriented Wireless Networks (CROWN-COM), July 8-10, 2013.

90. H. Zhou, H. Zheng, J. Wu, and J. Chen, “Energy-efficient Contact Probing in Opportunistic Mobile Networks,” Proc. of the 22nd International Conference on Computer Communications and Networks (ICCCN), July 30 - August 2, 2013.

91. H. Zheng, K. Li, C. C. Tan, and J. Wu, “User-based CPU Verification Scheme for Public Cloud Computing,” Proc. of the 6th IEEE International Conference on Cloud Computing (CLOUD), June 27-July 2, 2013.

92. W. Peng, F. Li, L. J. han, X. Zou, and J. Wu, “T-dominance: Prioritized Defense Deployment for BYOD Security,” Proc. of the First IEEE Conference on Communications and Network Security (CNS), October 14-16, 2013.

93. Y. Dai and J. Wu, “Boundary Helps: Efficient Routing Protocol using Directional Antennas in Cognitive Radio Networks,” Proc. of the 10h IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), October 14-16, 2013.

94. P. Ostovari, A. Khreishah, and J. Wu, “Multi-Layer Video Streaming with Helper Nodes using Network Coding,” Proc. of the 10h IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), October 14-16, 2013.

95. S. Yang and J. Wu, “Adaptive Backbone-based Routing in Delay Tolerant Networks,” Proc. of the 10h IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), October 14-16, 2013.

96. W. Peng, F. Li, X. Zou, and J. Wu, “The Virtue of Patience: Offloading Topical Cellular Content through Opportunistic Links,” Proc. of the 10h IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), October 14-16, 2013.

97. H. Zhou, J. Wu, H. Zhao, S. Tang, C. Chen, and J. Chen, “Incentive-Driven and Freshness-Aware Content Dissemination in Selfish Opportunistic Mobile Networks,” Proc. of the 10h IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), October 14-16, 2013.

98. Y. Yin, L. Xie, J. Wu, A. V. Vasilakos, and S. Lu, “Focus and Shoot: Efficient Identification over RFID Tags in the Specified Area,” Proc. of 10th International Conference on Mobile and Ubiquitous Systems: Computing, Networking and Services (MobiQuitous), December 2-14, 2013. (Received the Best Paper Award)

99. Md. Z. Alam Bhuiyan, G. Wang, J. Cao, and J. Wu, “Local Monitoring and Maintenance for Operational Wireless Sensor Networks,” Proc. of the 11th IEEE International Symposium on Parallel and Distributed Processing with Applications (ISPA), Melbourne, Australia, July 16-18, 2013. (Received the Best Paper Award)

100. D. Xiang, Z. Yu, and J. Wu, “Deadlock-Free Fully Adaptive Routing in Irregular Networks without Virtual Channels,” Proc. of the 11th IEEE International Symposium on Parallel and Distributed Processing with Applications (ISPA), Melbourne, Australia, July 16-18, 2013.

101. P. Ostovari, A. A. Khreishah, and J. Wu, “Cache Content Placement Using Triangular Network Coding,” Proc. of IEEE Wireless Communications and Networking Conference (WCNC), April 7-10, 2013.

102. Y. Zhao, J. Wu, C. Liu, and M. Lu, “A Hierarchical Naming System for Scalable Content Distribution in Large Networks,” Proc. of IEEE Wireless Communications and Networking Conference (WCNC), April 7-10, 2013.

103. W. Chang, J. Wu, C. C. Tan, and F. Li, “Sybil Defenses in Mobile Social etworks,” Proc. of the IEEE Global Telecommunications Conference (GLOBECOM), December 9-13, 2013.

104. C. Song, J. Wu, and M. Liu, “On Characterization of the Traffic Hole Problem in Vehicular Ad-hoc Networks,” Proc. of the IEEE Global Telecommunications Conference (GLOBECOM), December 9-13, 2013.

105. I. Jawhar, M. Ammar, S. Zhang, J. Wu, and N. Mohamed, “Ferry-Based Linear Wireless Sensor Networks,” Proc. of the IEEE Global Telecommunications Conference (GLOBECOM), December 9-13, 2013.

106. D. Li, J. Wu, K. Li, and K. Hwang, “Energy-aware Scheduling on Multiprocessor Platforms with Devices,” Proc. of the 3rd IEEE Conference on Cloud and Green Computing (CGC), September 30 - October 2, 2013.

107. P. Ostovari and J. Wu, “Throughput and Fairness Aware Dynamic Network Coding in Wireless Communication Networks,” Proc. of the 6th International Symposium on Resilient Control Systems (IS-RCS), August 13- 15, 2013.

108. X. Zhang, X. Dong, J. Wu, N. Xiong, and X. Li, “Fault-Aware Flow Control and Multi-path Routing in Wireless Sensor Networks,” Proc. of Workshop on Network Forensics, Security and Privacy (NFSP) (in conjunction with IEEE ICDCS 2013), July 2013.

109. H. Huang, S. Lin, A. Mamat, and J. Wu, “Predictive Scheduling for Spatial-dependent Tasks in Wireless Sensor Networks,” Proc. of the 19th IEEE Real-Time and Embedded Technology and Applications Symposium (RTAS, work in progress) (in conjunction with IEEE ICDCS 2013), April 9-11, 2013.

110. X. Li, J. Wu, and S. Lu, “QoS-Aware Service Selection in Geographically Distributed Clouds,” Proc. of the 7th International Workshop on Wireless Mesh and Ad Hoc Networks (WiMAN) (in conjunction with IEEE ICCCN 2013), July 30, 2013.

111. W. Jiang, J. Wu, and G. Wang, “RATE: Recommendation-aware Trust Evaluation in Online Social Networks,” Proc. of the 12th IEEE International Symposium on Network Computing and Applications (NCA, short paper), August 22 - 24, 2013.

112. M. Al-Hami, A. Khreishah, and J. Wu, “Video Streaming OverWireless LAN With Network Coding,” Proc. of the 12th IEEE International Symposium on Network Computing and Applications (NCA, short paper), August 22 - 24, 2013.

113. C. Song, J. Wu, W. Yang, and M. Liu, “Catching Up with Traffic Lights for Data Delivery in Vehicular Ad Hoc Networks,” Proc. of the 2nd ACM Annual International Workshop on Mission-Oriented Wireless Sensor Networking (MiSeNet, in conjunction with ACM MobiCom 2013), October 4, 2013.

114. H. Zheng and J. Wu, “Spectral Graph Multisection Through Orthogonality,” Proc. of 4th Workshop on Multiple Clusterings, Multi-view Data, and Multi-source Knowledge-driven Clustering (MultiClust) (in conjunction with ACM KDD 2013), August 11-14, 2013.

115. K. Li, H. Zheng, and J. Wu, “Migration-based Virtual Machine Placement in Cloud Systems,” Proc. of the 2nd IEEE International Conference on Cloud Networking (CloudNet), November 11-13, 2013.

116. K. Li, J. Wu, and A. Blaisse, “Elasticity-aware Virtual Machine Placement for Cloud Datacenters,” Proc. of the 2nd IEEE International Conference on Cloud Networking (CloudNet), November 11-13, 2013.

117. K. Chen, M. Lu, G. Tan, and J. Wu, “CRSM: Crowdsourcing based Road Surface Monitoring,” Proc. of the 11th IEEE/IFIP International Conference on Embedded and Ubiquitous Computing (EUC), November 13-15, 2013. (Received the Best Paper Award)

118. J. Wu and Y. Wang, “Social Feature-based Multi-Path Routing in Delay Tolerant Networks,” Proc. of the 31st IEEE International Conference on Computer Communications (INFOCOM), March 2012.

119. S. Zhang, Z. Qian, J. Wu, and S. Lu, “An Opportunistic Resource Sharing and Topology-Aware Mapping Framework for Virtual Networks,” Proc. of the 31st IEEE International Conference on Computer Communications (INFOCOM), March 2012.

120. Y. Li, J. Li, J. Ren, and J. Wu, “Providing Hop-by-Hop Authentication and Source Privacy in Wireless Sensor Networks,” Proc. of the 31st IEEE International Conference on Computer Communications (INFOCOM) (Mini-Conference), March 2012.

121. Q. Liu, C. C. Tan, J. Wu, and G. Wang, “Efficient Information Retrieval for Ranked Queries in Cost-Effective Cloud Environments,” Proc. of the 31st IEEE International Conference on Computer Communications (INFOCOM) (Mini-Conference), March 2012.

122. A. Khreishah, I. M. Khalil, and J. Wu, “Distributed Network Coding Based Opportunistic Routing for Multicast,” Proc.of 13th ACM International Symposium on Mobile Ad Hoc Networking and Computing (MobiHoc), June 2012.

123. W. Chang, J. Wu, and C. C. Tan, “Encounter-based Noise Cancelation For Cooperative Trajectory Mapping,” Proc. of the 10th IEEE Pervasive Computing and Communication Conference (PerCom), March 2012.

124. N. Xiong, A. V. Vasilakos, J. Wu, Y. R. Yang, A. Rindos, and Y. Pan. ”A Class of Practical Self-tuning Failure Detection Schemes for Cloud Computing Networks,” Proc. of the 26th IEEE International Parallel & Distributed Processing Symposium (IPDPS), May 2012.

125. W. Peng, F. Li, X. Zou, and J. Wu, “Seed and Grow: An Attack Against Anonymized Social Networks,” Proc. of the IEEE Communications Society Conference on Sensor, Mesh and Ad Hoc Communications and Networks (SECON), June 2012.

126. A. Khreishah, I. M. Khalil, and J. Wu, “Universal Opportunistic Routing Scheme using Network Coding,” Proc. of the IEEE Communications Society Conference on Sensor, Mesh and Ad Hoc Communications and Networks (SECON), June 2012.

127. W. Peng, F. Li, X. Zou, and J. Wu, “A Privacy-Preserving Social-Aware Incentive System for Word-of-Mouth Advertisement Dissemination on Smart Mobile Devices,” Proc. of the IEEE Communications Society Conference on Sensor, Mesh and Ad Hoc Communications and Networks (SECON), June 2012.

128. Z. Jiang, D. Wu, M. Guo, J. Wu and X. Wang, “Minimum Latency Broadcasting with Conflict Awareness in Wireless Sensor Networks,” Proc. of the 41st International Conference on Parallel Processing (ICPP), September 2012.

129. D. Li and J. Wu, “Energy-aware Scheduling for Frame-based Tasks on Heterogeneous Multiprocessor Platforms,” Proc. of the 41st International Conference on Parallel Processing (ICPP), September 2012.

130. M. Xiao, J. Wu, and L. Huang, “Time-Sensitive Utility-Based Routing in Duty-Cycle Wireless Sensor Networks with Unreliable Links,” Proc. of the 31st International Symposium on Reliable Distributed Systems (SRDS), October 2012.

131. Z. Jiang, J. Wu, M. Guo, Z. Zhao, and D. Wu, “A Quick and Reliable Routing for Infrastructure Surveillance with Wireless Sensor Networks,” Proc. of the 31st International Symposium on Reliable Distributed Systems (SRDS), October 2012.

132. S. Zhang, J. Wu, and S. Lu, “Collaborative Mobile Charging for Sensor Networks,” Proc. of the 9th IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), October 2012. (Received the Best Paper Runner-up Award)

133. S. Tang, J. Wu, G. Chen, C. Wang, X. Liu, T. Li, and X. Li, “On Minimum Delay Duty-Cycling Protocol in Sustainable Sensor Network,” Proc. of IEEE International Conference on Networked Protocols (ICNP) , October 2012.

134. Q. Liu, Y. Guo, J. Wu, and G. Wang, “Dynamic Grouping Strategy in Cloud Computing,” Proc. of the 2nd International Conference on Cloud and Green Computing (CGC), November 2012. (Received the Best Student Paper Award)

135. Q. Liu, G. Wang, and J. Wu, “Clock-Based Proxy Re-encryption Scheme in Unreliable Clouds,” Proc. of International Workshop on Security in Cloud Computing (CloudSec), September 2012.

136. Y. Dai and J. Wu, “Distributed Rerouting For Multiple Sessions in Cognitive Radio Networks,” Proc. of 11th IEEE International Conference on Ubiquitous Computing and Communications (IUCC), June 2012.

137. W. Chang, J. Wu, and C. C. Tan, “Cooperative Trajectory-based Map Construction,” Proc. of the 11th IEEE International Conference on Trust, Security and Privacy in Computing and Communications (TrustCom), June 2012.

138. P. Ostovari, A. A. Khreishah, and J. Wu, “Deadline-aware Broadcasting in Wireless Networks with Local Network Coding,” Proc. of the International Conference on Computing, Networking and Communications, Wireless Networks Symposium (ICNC), January 30 - February 2, 2012.

139. Y. Wang and J. Wu, “A Joint Replication-Migration-based Routing in Delay Tolerant Networks,” Proc. of IEEE International Conference on Communications (ICC), June, 2012.

140. H. Quan, A. Milicic, S. Vucetic, and J. Wu. “A Connectivity-Based Popularity Prediction Approach for Social Networks,” Proc. of IEEE International Conference on Communications (ICC), June, 2012.

141. P. Ostovari, A. Khreishah, and J. Wu, “Deadline-aware Broadcasting in Wireless Networks with Network Coding,” Proc. of the IEEE Global Telecommunications Conference (GLOBECOM), December 2012.

142. S. Zhang, J. Wu, and S. Lu, “Virtual Network Embedding with Substrate Support for Parallelization,” Proc. of the IEEE Global Telecommunications Conference (GLOBECOM), December 2012.

143. D. Lexie, S. Lin, and J. Wu, “Adaptive Battery Charge Scheduling with Bursty Workloads,” Proc. of the IEEE Global Telecommunications Conference (GLOBECOM), December 2012.

144. J. Robison, L. Bai, D. Mastrogiannis, C. C. Tan, and J. Wu, “A Survey on PHR Technology,” Proc. of the IEEE International Conference on e-Health Networking, Applications and Services (Healthcom), October 2012.

145. C. C. Tan, L. Bai, D. Mastrogiannis, and J. Wu, “Security Analysis of Emerging Remote Obstetrics Monitoring Systems’,” Proc. of the IEEE International Conference on e-Health Networking, Applications and Services (Healthcom), October 2012.

146. M. Korostelev, L. Bai, J. Wu, C. C. Tan, and D. Mastrogiannis, “Body Sensor Networks in Fetal Monitoring with NFC Enabled Android Devices,” Proc. of the 7th International Conference on Body Area Networks (Bodynets), September 2012.

147. Y. Zeng, J. Wu, N. Xiong, and D. Li, “Energy-Efficient Routing and Rate Allocation for Delay Tolerant Networks,” Proc. of the 2nd International Workshop on Cyber-physical Networking Systems (CPNS) (in conjunction with IEEE ICDCS), June 2012.

148. C. Song, J. Wu, M. Liu, H. Gong, and B. Gou, “RESen: Sensing and Evaluating the Riding Experience Based on Crowdsourcing by Smart Phones,” Proc. of the 8th IEEE International Conference on Mobile Ad-hoc and Sensor Networks (MSN), pp. 147-152, December 2012.

149. Y. Zhao and J. Wu, “Towards Approximate Event Processing in A Large-Scale Content-Based Network,” Proc. of the 31st International Conference on Distributed Computing Systems (ICDCS), June 2011.

150. Y. Zhao and J. Wu, “On the Construction of the Minimum Cost Content-Based Publish/Subscribe Overlays,” Proc. of the 8th Annual IEEE Communications Society Conference on Sensor, Mesh and Ad-Hoc Communications and Networks (SECON), June 2011.

151. C. C. Tan, Q. Liu , and J. Wu, “Secure Locking for Untrusted Clouds,” Proc. of the 4th International Conference on Cloud Computing (Cloud), July 2011.

152. Y. Zhao and J. Wu, “ZigZag: A Content-based Publish/Subscribe Architecture for Human Networks,” Proc. of the 20th IEEE International Conference on Computer Communications and Networks (ICCCN), July 2011.

153. J. Wu, Y. Dai, and Y. Zhao, “Local Spectrum Assignments in Cognitive Radio Networks,” Proc. of the 20th IEEE International Conference on Computer Communications and Networks (ICCCN), July 2011.

154. Y. Wang, Y. Guo, and J. Wu, “Making Many People Happy: Greedy Solutions for Content Distribution,” Proc. of 2011 International Conference on Parallel Processing (ICPP), September 2011.

155. C. Liu, A. Chen, D. Wu, J. Zhang, and J. Wu, “A DFA with Extended Character-set for Fast Deep Packet Inspection,” Proc. of 2011 International Conference on Parallel Processing (ICPP), September 2011.

156. X. Li and J. Wu, “Fine-grained Localization with Pairwise Nodes Coverage,” Proc. of SAHNS workshop (in conjunction with IEEE ICDCS), June 2011.

157. P. Ostovari, M. Dehghan, and J. Wu, “Connected Point Coverage in Wireless Sensor Networks using Robust Spanning Tree,” Proc. of WWASN workshop (in conjunction with IEEE ICDCS 2011), June 2011.

158. M. Taifi, A. Khreishah, J. Y. Shi, and J. Wu, “Sustainable GPU Computing at Scale,” Proc. of the 14th IEEE International Conference on Computational Science and Engineering (CSE), August 2011.

159. C. Liu, J. Wu, X. Guan, and L. Chen, “Cooperative File Sharing in Hybrid Delay Tolerant Networks,” Proc. of WWASN Workshops (in conjunction with IEEE ICDCS), June 2011.

160. Y. C. Zhao, J. Wu, and S. L. Lu, “Efficient SINR Estimating with Accuracy Control in Large Scale Cognitive Radio Networks,” Proc. of the 17th IEEE International Conference on Parallel and Distributed Systems (ICPADS), December 2011.

161. Y. Dai and J. Wu, “Efficient Channel Assignment Under Dynamic Source Routing in Cognitive Radio Networks,” Proc. of the 8th IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), October 2011.

162. A. Khreishah, I. M. Khalil, and J. Wu, “Polynomial Time and Provably Efficient Network Coding Scheme for Lossy Wireless Networks,” Proc. of the 8th IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), October 2011.

163. W. Peng, F. Li, X. Zou, and J. Wu, “Behavioral Detection and Containment of Proximity Malware in Delay Tolerant Networks,” Proc. of the 8th IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), October 2011.

164. A. Srinivasan and J. Wu, “Duplicate File Names- A Novel Steganographic Data Hiding Technique,” Proc. of 2011 International Workshop on Identity: Security, Management & Applications (ID), July 2011.

165. Q. Liu, C. C. Tan, J. Wu, and G. Wang, “Reliable Re-encryption in Unreliable Clouds,” Proc. of the IEEE Global Telecommunications Conference (GLOBECOM), December 2011.

166. W. Chang, J. Wu, and C. C. Tan, “Enhancing Mobile Social Network Privacy,” Proc. of the IEEE Global Telecommunications Conference (GLOBECOM), December 2011.

167. A. Khreishah, J. Wu, P. Ostovari, and I. M. Khalil, “Flow Based XOR Network Coding for Lossy Wireless Networks,” Proc. of the IEEE Global Telecommunications Conference (GLOBECOM), December 2011.

168. S. Zhang, Z. Qian, B. Tang, J. Wu, and S. Lu, “Opportunistic Bandwidth Sharing for Virtual Network Mapping,” Proc. of the IEEE Global Telecommunications Conference (GLOBECOM), December 2011.

169. Z. Guo, Z. Peng, B. Wang, J. Cui, and J. Wu, “Adaptive Routing in Underwater Delay Tolerant Sensor Networks,” Proc. of the 6th International ICST Conference on Communications and Networking in China (ChinaCom), April 2011.

170. Y. Zhao, J. Wu, and S. Lu, “Throughput Maximization in Cognitive Radio Based Wireless Mesh Networks,” Proc. of IEEE the 30th Military Communications Conference (MILCOM), November 2011.

171. M. Peng, Y. Xiao, H. Chen, Q. Hao, A. Vasilakos, and J. Wu, “Sensor Distribution on Coverage in Sensor Networks” Proc. of the 7th international ICST Conference on Heterogeneous Networking for Quality, Reliability, Security and Robustness (QShine), November 2010.

172. X. Zhang, H. Gong, M. Liu, S. Lu, and J. Wu, “Quantitative Analysis of the Effect of Transmitting Power on the Capacity of Wireless Ad Hoc Networks,” Proc. of the 11th ACM International Symposium on Mobile Ad Hoc Networking and Computing (MobiHoc), September 2010.

173. Y. Wang, K. Li, and J. Wu, “Distance Estimation by Constructing the Virtual Ruler in Anisotropic Sensor Networks,” Proc. of the 28th Annual Joint Conference of IEEE Communication and Computer Society (INFOCOM), March 2010.

174. F. Li, Y. Yang, and J. Wu, “CPMC: An Efficient Proximity Malware Coping Scheme in Smartphonebased Mobile Networks,” Proc. of the 28th Annual Joint Conference of IEEE Communication and Computer Society (INFOCOM), March 2010.

175. F. Xu, C. Tan, Q. Li, G. Yan, and J. Wu, “Designing a Better Access Point Association Protocol,” Proc. of the 28th Annual Joint Conference of IEEE Communication and Computer Society (INFOCOM), March 2010.

176. Y. Zhao, F. Li, J. Wu, and S. Lu, “VBS: Maximum Lifetime Sleep Scheduling for Wireless Sensor Networks Using Virtual Backbones,” Proc. of the 28th Annual Joint Conference of IEEE Communication and Computer Society (INFOCOM) (Mini-Conference), March 2010.

177. Y. Zhao and J. Wu, “B-Sub: A Practical Bloom-Filter-Based Publish-Subscribe System for Human Networks,” Proc. of the 30th International Conference on Distributed Computing Systems (ICDCS), June 2010.

178. Z. Xiao, Y. Xiao, and J. Wu, “A Quantitative Study of Accountability in Wireless Multi-hop Networks,” Proc. of the 39th International Conference on Parallel Processing (ICPP), September 2010.

179. Z. Jiang, J. Wu, and R. Ito, “A Metric for Routing in Delay-Sensitive Wireless Sensor Networks,” IEEE of the 7th IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), November 2010. (Received the Best Paper Award)

180. X. Li and J. Wu, “Channel On Demand: Optimal Capacity for Cooperative Multi-channel Multiinterface Wireless Mesh Networks,” IEEE of the 7th IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), November 2010.

181. J. Wu and Y. Wang, “A Non-Replication Multicasting Scheme in Delay Tolerant Networks,” IEEE of the 7th IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), November 2010.

182. W. Zhang, X. Du, J. Wu, S. Soysa, and Y. Liu, “Near-Minimum-Energy Routing in Heterogeneous Wireless Sensor Networks,” Proc. of IEEE Global Communications Conference (GLOBECOM), December 2010.

183. Y. Wang, X. Li, and J. Wu, “Multicasting in Delay Tolerant Networks: Delegation Forwarding,” Proc. of IEEE Global Communications Conference (GLOBECOM), December 2010.

184. Y. Dai, P. Yang, G. Chen, and J. Wu, “CFP: Integration of Fountain Codes and Optimal Probabilistic Forwarding in DTNs,” Proc. of IEEE Global Communications Conference (GLOBECOM), December 2010.

185. X. Hei, X. Du, J. Wu, and F. Hu, “Defending Resource Depletion Attacks on Implantable Medical Devices,” Proc. of IEEE Global Communications Conference (GLOBECOM), December 2010.

186. Y. Zhao and J. Wu, “Socially-Aware Publish/Subscribe System for Human Networks,” Proc. of IEEE Wireless Communications and Networking Conference (WCNC), April 2010.

187. J. Zheng, W. Jia, G. Wang, and J. Wu, “Target Trajectory Querying in Wireless Sensor Networks,” Proc. of IEEE International Conference on Communications (ICC), May 2010.

188. Y. Zhao and J. Wu, “Stochastic Sleep Scheduling for Large Scale Wireless Sensor Networks,” Proc. of IEEE International Conference on Communications (ICC), May 2010.

189. G. Wang, Q. Liu, and J. Wu, “Hierarchical Attribute-Based Encryption for Fine-Grained Access Control in Cloud Computing,” Proc. of the 17th ACM Conference on Computer and Communications Security (CCS) (poster abstract), October 2010.

190. F. Li, Y. Yang, J. Wu, and X. Zou, “Fuzzy Closeness-based Delegation Forwarding in Delay Tolerant Networks,” Proc. of the 5th IEEE International Conference on Networking, Architecture, and Storage (NAS), July 2010.

191. Y. Wang, Z. Zhang, and J. Wu, “A Distributed Approach for Hidden Wormhole Detection With Neighborhood Information,” Proc. of the 5th IEEE International Conference on Networking, Architecture, and Storage (NAS), July 2010.

192. Z. Jiang, Z. Li, N. Xiao, and J. Wu, “CR: Capability-Information-Based Routing in Wireless Ad Hoc Sensor Networks,” Proc. of the 5th IEEE International Conference on Networking, Architecture, and Storage (NAS), July 2010.

193. Y. Hu, W. Li, X. Chen, and J. Wu, “A Probabilistic Routing Protocol for Heterogeneous Sensor Networks,” Proc. of the 5th IEEE International Conference on Networking, Architecture, and Storage (NAS), July 2010.

194. Y. Ge, G. Wang, and J. Wu, “Node-Disjoint Multipath Routing with Group Mobility in MANETs,” Proc. of the 5th IEEE International Conference on Networking, Architecture, and Storage (NAS), July 2010.

195. X. Zhang, M. Liu, H. Gong, S. Lu, and J. Wu, “PCAR: A Power Controlled Routing Protocol for Wireless Ad Hoc Networks,” Proc. of IEEE International Symposium on a World of Wireless, Mobile and Multimedia Networks (WoWMoM), June 2010.

196. Q. Liu, G. Wang, and J. Wu, “Efficient Sharing of Secure Cloud Storage Services,” Proc. of the 3rd IEEE International Symposium on Trust, Security and Privacy for Emerging Applications (TSP) (in conjunction with CIT), July 2010.

197. X. Zhang, H. Gong, M. Liu, S. Lu, and J. Wu, “The Lower Bound on the Capacity of Wireless Ad Hoc Networks,” Proc. of IEEE International Conference on Wireless Communications, Networking and Information Security (WCNIS), June 2010.

198. Y. Xie, G. Wang, and J. Wu, “Fuzzy Location Service for Mobile Ad Hoc Networks,” Proc. of the 10th IEEE International Conference on Computer and Information Technology (CIT), July 2010.

199. G. Chen, Z. Li, J. Wu, and J. Xi, “On Maximum Stability with Enhanced Scalability in High-Churn D-HT Deployment,” Proc. of IEEE International Conference on Parallel Processing (ICPP), September 2009.

200. L. Xie, Q. Li, W. Mao, J. Wu, and D. Chen, “Achieving Efficiency and Fairness for Association Control in Vehicular Networks,” Proc. of the 17th IEEE International Conference on Network Protocols (ICNP), October 2009.

201. C. Liu and J. Wu, “An Optimal Probabilistically Forwarding Protocol in Delay Tolerant Networks,” Proc. of the 10th ACM International Symposium on Mobile Ad Hoc Networking and Computing (MobiHoc), May 2009.

202. Q. Yuan, I. Cardei, and J. Wu, “Predict and Relay: An Efficient Routing in Disruption-Tolerant Networks,” Proc. of 10th ACM International Symposium on Mobile Ad Hoc Networking and Computing (MobiHoc), May 2009.

203. F. Li and J. Wu, “MOPS: Providing Content-based Service in Disruption-tolerant Networks,” Proc. of the 29th International Conference on Distributed Computing Systems (ICDCS), June 2009.

204. M. Lu and J. Wu, “Erasure-Coding Based Utility Routing in Multi-hop Wireless Networks,” Proc. of the 6th IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), October 2009.

205. C. Liu, J. Wu, and I. Cardei, “Message Forwarding in Cyclic MobiSpace: The Multi-copy Case,” Proc. of the 6th IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), October 2009.

206. F. Li and J. Wu, “LocalCom: A Community-based Epidemic Forwarding Scheme in Disruptiontolerant Networks,” Proc. of IEEE Communications Society Conference on Sensor, Mesh and Ad Hoc Communications and Networks (SECON), June 2009.

207. Md. Z. Alam Bhuiyan, G. Wang, and J. Wu, “Target Tracking with Monitor and Backup Sensors in Wireless Sensor Networks,” Proc. of the 18th IEEE International Conference on Computer Communications and Networks (ICCCN), August 2009.

208. F. Li, A. Srinivasan, and J. Wu, “Thwarting Blackhole Attacks in Disruption-Tolerant Networks using Encounter Tickets,” Proc. of the 27th Annual Joint Conference of IEEE Communication and Computer Society (INFOCOM), April 2009.

209. M. Li and J. Wu, “Opportunistic Routing Algebra and Its Applications,” Proc. of the 27th Annual Joint Conference of IEEE Communication and Computer Society (INFOCOM), April 2009.

210. C. Liu and J. Wu, “Efficient Geometric Routing in Three Dimensional Ad Hoc Networks,” Proc. of the 27th Annual Joint Conference of IEEE Communication and Computer Society (INFOCOM) (Mini-Conference), April 2009.

211. X. Chen, Q. Gu, Z. Jiang, and J. Wu, “Sink-Anonymity Mobility Control in Wireless Sensor Networks,” Proc. of the 5th International Conference on Wireless and Mobile Computing, Networking and Communications (WiMob), October 2009.

212. R. Thanawala, J. Wu, and A. Srinivasan, “Efficient Resource Discovery in Mobile Ad Hoc Networks,” Proc. of IEEE International Conference on Communications (ICC), June 2009.

213. C. Liu and J. Wu, “Efficient Adaptive Routing in Delay Tolerant Networks,” Proc. of IEEE International Conference on Communications (ICC), June 2009.

214. F. Li and J. Wu, “FRAME: AN Innovative Incentive Scheme in Vehicular Networks,” Proc. of IEEE International Conference on Communications (ICC), June 2009.

215. X. Chen, J. Shen, T. Groves, and J. Wu, “Probability Delegation Forwarding in Delay Tolerant Networks,” Proc. of the 18th IEEE International Conference on Computer Communications and Networks (ICCCN), August 2009.

216. Md. Z. Alam Bhuiyan, G. Wang, and J. Wu, “Polygon-Based Tracking Framework in Surveillance Wireless Sensor Networks,” Proc. of the 5th International Conference on Parallel and Distributed Systems (ICPADS), December 2009.

217. P. Li, G. Wang, J. Wu, and H. C. Yang, “Hole Reshaping Routing in Large-Scale Mobile Ad-Hoc Networks,” Proc. of IEEE Global Communications Conference (GLOBECOM), November-December 2009.

218. X. Xing, G. Wang, J. Wu, and J. Li, “Square Region-Based Coverage and Connectivity Probability Model in Wireless Sensor Networks,” Proc. of the 5th International Conference on Collaborative Computing: Networking, Applications and Worksharing (CollaborateCom), November 2009.

219. F. Qi, Z. Tang, G. Wang, and J. Wu, “QoS-aware Optimization Strategy for Security Ranking in SSL Protocol,” Proc. of IEEE TSP-09 (in conjunction with IEEE MASS 2009), October 2009.

220. X. Xing, G. Wang, and J. Wu, “Herd-Based Target Tracking Protocol in Wireless Sensor Networks,” Proc. of the International Conference on Wireless Algorithms, Systems and Applications (WASA), LNCS 5682, August 2009, 135-148.

221. Q. Liu, G. Wang, and J. Wu, “An Efficient Privacy Preserving Keyword Search Scheme in Cloud Computing,” Proc. of IEEE/IFIP TrustCom-09 (in conjunction with IEEE/IFIP EUC), August 2009.

222. X. Chen, J. Shen, and J. Wu, “A Novel Information Model for Efficient Routing Protocols in Delay Tolerant Networks,” Proc. of the 8th International Workshop PMEO UCNS (in conjunction with IEEE IPDPS 2009), May 2009.

223. J. Chen and J. Wu, “Cooperative Anonymity Authentication in Vehicular Networks,” Proc. of the 5th IEEE International Workshop on Wireless and Sensor Networks Security (WSNS), August 2009.

224. E. B. Fernandez, J. Wu, M. M. Larrondo-Petrie, and Y. Shao, “On Building Secure SCADA Systems using Security Patterns,” Proc. of of the 5th Annual Workshop on Cyber Security and Information Infrastructure Research (CSIIRW), April 2009.

225. C. Liu and J. Wu, “Routing in a Cyclic MobiSpace,” Proc. of the 9th ACM International Symposium on Mobile Ad Hoc Networking and Computing (MobiHoc), 2008.

226. C. Liu and J. Wu, “Force-based Geometric Routing,” Proc. of the 9th ACM International Symposium on Mobile Ad Hoc Networking and Computing (MobiHoc) (poster Abstract), 2008.

227. F. Li and J. Wu, “A Winning-Probability-based Incentive Scheme in Vehicular Networks,” Proc. of IEEE International Conference on Network Protocols (ICNP) (poster abstract), 2008.

228. Z. Jiang and J. Wu, “A Hierarchical Structure based Coverage Repair in Wireless Sensor Networks,” Proc. of IEEE International Symposium on Personal, Indoor and Mobile Radio Communications (PIMRC), 2008.

229. I. Cardei, C. Liu, J. Wu, and Q. Yuan, ”DTN Routing with Probabilistic Trajectory Prediction,” Proc. of the International Conference on Wireless Algorithms, Systems and Applications (WASA), 2008.

230. J. Wu, M. Lu, and F. Li, “Utility-Based Opportunistic Routing in Multi-hop Wireless Networks,” Proc. of the 28th International Conference on Distributed Computing Systems (ICDCS), 2008.

231. Y. Wang and J. Wu, “A Nonblocking Approach for Reaching an Agreement on Request Total Orders,” Proc. of the 28th International Conference on Distributed Computing Systems (ICDCS), 2008.

232. J. Wu, S. Yang, and M. Cardei, “On Maintaining Sensor-Actor Connectivity in Wireless Sensor and Actor Networks,” Proc. of the 26th Annual Joint Conference of IEEE Communication and Computer Society (INFOCOM), 2008.

233. Z. Jiang, J. Ma, and W. Lou, and J. Wu, “An Information Model for Geographic Greedy Forwarding in Wireless Ad Hoc Sensor Networks,” Proc. of the 26th Annual Joint Conference of IEEE Communication and Computer Society (INFOCOM), 2008.

234. J. Wu, S. Yang, and M. Cardei, “Efficient Broadcast in MANETs Using Network Coding and Directional Antennas, ” Proc. of the 26th Annual Joint Conference of IEEE Communication and Computer Society (INFOCOM), 2008.

235. Q. Yuan and J. Wu, “DRIP: A Dynamic VoRonoi RegIons-Based Publish/Subscribe Protocol in Mobile Computing”, Proc. of the 26th Annual Joint Conference of IEEE Communication and Computer Society (INFOCOM) (Mini-Conference), 2008.

236. C. Liu and J. Wu, “Adaptive Routing in Dynamic Ad Hoc Networks” Proc. of IEEE Wireless Networks and Communication Conference (WCNC), 2008.

237. B. Qian and J. Wu, Multi-Path Intelligent Virtual Mobile Nodes for Ad Hoc Networks, Proc. of the 17th IEEE International Conference on Computer Communications and Networks (ICCCN), August 2008.

238. M. Lu and J. Wu, Localized Access Point Association in Wireless LANs with Performance Guarantee, Proc. of the 17th IEEE International Conference on Computer Communications and Networks (ICCCN), August 2008.

239. A. Srinivasan and J.Wu, A Novel Connected Dominating Set Based Sink Mobility Model for WSNs, Proc. of the 17th IEEE International Conference on Computer Communications and Networks (ICC-CN), August 2008.

240. X. Chen, Z. Jiang, and J. Wu, “Quick Convergence Mobility Control Schemes in Wireless Sensor Networks,” Proc. of the 10th Workshop on Advances on Parallel and Distributed Processing Symposium (APDCM) (in conjuection with IPDPS’08), 2008.

241. M. Lu and J. Wu, “Utility-Based Data-Gathering in Wireless Sensor Networks with Unstable Links,” Proc. of the 9th International Conference on Distributed Computing and Networking (ICDCN), 2008.

242. A. Srinivasan, F. Li, and J. Wu, “A Novel CDS-based Reputation Monitoring System for Wireless Sensor Networks,” Proc. of the 1st International Workshop on Wireless Security and Privacy (WiSP) (in conjunction with ICDDC), 2008.

243. Y. Yang, M. Cardei, and J. Wu, “Non-Uniform Sensor Deployment in Mobile Wireless Sensor Networks, Proc. of IEEE International Symposium on a World of Wireless Mobile and Multimedia Networks (WoWMoM), June 2008.

244. Z. Jiang and J. Wu, “Mobility Control for Complete Coverage in Wireless Sensor Networks,” Proc. of the Fifth Workshop on Wireless Ad hoc and Sensor Networks (WWASN) (in conjunction with ICDCS), 2008.

245. Y. Shao and J. Wu, “Understanding the Tolerance of Dynamic Networks: A Routing-Oriented Approach”, Proc. of the 3rd International Workshop on Delay/Disruption-Tolerant Mobile Networks (DTMN) (in conjunction with ICDCS), 2008.

246. C. Liu and J. Wu, “Scalable Routing in Delay Tolerant Networks,” Proc. of ACM International Symposium on Mobile Ad Hoc Networking and Computing (MobiHoc), 2007.

247. M. Lu and J. Wu, “Localized Access Point Selection in Infrastructure Wireless WLAN,” Proc. of IEEE Military Communications Conference (MILCOM), 2007.

248. F. Li and J. Wu, “Authentication Via Ambassadors: A Novel Authentication Mechanism in MANETs,” Proc. of IEEE Military Communications Conference (MILCOM), 2007.

249. J. Wu and S. Yang, “Polylogarithmic Store-Carry-Forward Routing using Mobile Nodes,” Proc. of IEEE Mobile Ad-Hoc and Sensor Systems (MASS), 2007.

250. M. Cardei, S. Yang, and J. Wu, “Fault-Tolerant Topology Control Algorithms for Heterogeneous Wireless Sensor Networks,” Proc. of IEEE Mobile Ad-Hoc and Sensor Systems (MASS), 2007.

251. C. Liu and J. Wu, “Destination-Region-Based Local Minimum Aware Geometric Routing,” Proc. of IEEE Mobile Ad-Hoc and Sensor Systems (MASS), 2007.

252. F. Li and J. Wu, “Uncertainty Mitigation for Utility-Oriented Routing in Wireless Ad Hoc Networks,” Proc. of IEEE Global Communications Conference (GLOBECOM), 2007.

253. M. Lu and J. Wu, “Utility-based Routing in Communication Networks with Unstable Links,” Proc. of IEEE International Conference on Computer Communications and Networks (ICCCN), 2007. (Nominated for the Best Paper Award)

254. W. Tang, H. Xiong, S. Zhong, and J. Wu, “Enhancing Semi-Supervised Clustering: A Feature Projection Perspective,” Proc. of ACM International Conference on Knowledge Discovery and Data Mining (SIGKDD), 2007.

255. D. Xiang, Y. Zhang, Y. Pan and J. Wu, “Deadlock-Free Adaptive Routing in Meshes Based on Cost-Effective Deadlock Avoidance Schemes,” Proc. of International Conference on Parallel Processing (ICPP), 2007.

256. J. Zhao, H. Yu, K. Zhang, W. Zheng, J. Wu, and J. Hu, “Achieving Reliability Through Replication in a Wide-Area Network DHT Storage System,” Proc. of International Conference on Parallel Processing (ICPP), 2007.

257. J. Wang, J. Cao, J. Li and J. Wu, “MHH: A Novel Protocol for Mobility Management in Publish/Subscribe System,” Proc. of International Conference on Parallel Processing (ICPP), 2007.

258. S. Yang, J. Wu, and F. Dai ”Efficient Backbone Construction Methods in MANETs Using Directional Antennas,” Proc. of IEEE International Conference on Distributed Computing Systems (ICDCS), June 2007.

259. J. Wu, “Dynamic Carrier-Assisted Routing in Mobile Networks,” Proc. of the 5th IEEE International Conference on Pervasive Computing and Communications (PerCom), Work-in-Progress, March 2007.

260. F. Li and J. Wu, “Mobility Reduces Uncertainty in MANETs,” Proc. of the 26th Annual Joint Conference of IEEE Communication and Computer Society (INFOCOM), May 2007.

261. D. Guo, J. Wu, H. Chen, and X. Luo, “MOORE: An Extendable Peer-to-Peer Network Based Incomplete Kautz Digraph With Constant Degree,” Proc. of the 26th Annual Joint Conference of IEEE Communication and Computer Society (INFOCOM), May 2007.

262. T. Anantvalee and J. Wu, “Reputation-Based System for Encouraging the Cooperation of Nodes in Mobile Ad Hoc Networks,” Proc. of IEEE International Conference on Communications (ICC), June 2007.

263. A. Srinivasan and J. Wu, “A Novel k-Parent Flooding Tree for Secure and Reliable Broadcasting in Sensor Networks,” Proc. of IEEE International Conference on Communications (ICC), June 2007.

264. X. Li, J. Wu, and J. Xu, “Searching using Probabilistic Hints in Unstructured P2Ps,” Proc. of the International Conference on Parallel and Distributed Computing and Systems (PDCS), 2007.

265. J. Huang, X. Li, and J. Wu, “A Class-Based Search System in Unstructured P2P Networks,” Proc. of the 21st IEEE International Conference on Advanced Information Networking and Applications (AINA), May 2007.

266. N. Delessy, E. B. Fernandez, M. M. Larrondo-Petrie, and J. Wu. “Patterns for Access Control in Distributed Systems,” Proc. of the 14th Conference on Pattern Languages of Programs (PLoP), September 2007.

267. Z. Jiang, J. Wu, A. Agah, and B. Lu. “Topology Control for Secured Coverage in Wireless Sensor Networks,” Proc. of the 3rd IEEE International Workshop on Wireless and Sensor Networks Security (WSNS), October 2007.

268. Z. Jiang and J. Wu, “On Achieving the Shortest-Path Routing in 2-D Meshes,” Proc. of the 9th Workshop on Advanced in Parallel and Distributed Computational Models (APDCM) (in conjunction with IEEE IPDPS), 2007.

269. T. Xu and J. Wu, “Quorum-Based IP Address Autoconfiguration in Mobile Ad Hoc Networks,” Proc. of the 5th Workshop on Mobile Distributed Computing (MCD) (in conjunction with IEEE ICDCS), 2007.

270. V. Ponce, J. Wu, and X. Li, “Improve Peer Cooperation Using Social Networks,” International Workshop on Hot Topics in Mobile Peer-to-Peer Networks (HOT-MP2P) (in conjunction with ICP-P), September 2007.

271. A. Srinivasan, J. Wu, F. Li, M. Li, “CAGE: Clique-based Assignment of Group kEy,” Proc. of the Second International Conference on Communications and Networking (ChinaCom), 2007.

272. M. Lu and J. Wu, “Social Welfare-Based Routing in Ad Hoc Networks,” Proc. of International Conference on Parallel Processing (ICPP), 2006.

273. J. Cao, J. -L. Wang, and J. Wu, “Achieving Bounded Delay on Message Delivery in Publish/Subscribe Systems,” Proc. of International Conference on Parallel Processing (ICPP), 2006.

274. J. Wu and S. Yang, “Optimal Movement-Assisted Sensor Deployment and Its Extensions in Wireless Sensor Networks,” Proc. of the 20th International Conference on Parallel and Distributed Systems (ICPADS), 2006.

275. J. Zhang and J. Wu, “XYZ: A Scalable, Partially Centralized Lookup Service for Large-Scale Peerto-Peer Systems,” Proc. of the 20th International Conference on Parallel and Distributed Systems (ICPADS), 2006.

276. M. Lu, F. Li, and J. Wu, “Incentive Compatible Cost- and Stability-Based Routing in Ad Hoc Networks,” Proc. of the 20th International Conference on Parallel and Distributed Systems (ICPADS), 2006.

277. C. Liu and J. Wu, “Position-Based Routing using Virtual Small World in MANETs,” Proc. of IEEE Global Telecommunications Conference (GLOBECOM), 2006.

278. D. Guo, J. Wu, H. Chen, and X. Luo, “Theory and Network Applications of Dynamic Bloom Filters,” Proc. of the 25th Annual Joint Conference of IEEE Communication and Computer Society (INFO-COM), April 2006.

279. B. Wu and J. Wu, “k-Anycast Routing Schemes for Mobile Ad Hoc Networks,” Proc. of the 20th International Parallel and Distributed Processing Symposium (IPDPS), April 2006.

280. I. Jawhar and J. Wu, “Resource Allocation in Wireless Networks Using Directional Antennas,” Proc. of the 4th IEEE International Conference on Pervasive Computing and Communications (PerCom), March 2006.

281. M. Ye, E. Chan, G. Chen, and J. Wu, “Energy Efficient Fractional Coverage Schemes for Low Cost Wireless Sensor Networks,” Proc. of WASN Workshop (in conjunction with IEEE ICDCS), 2006.

282. X. Li, S. Zhong, and J. Wu, “Improve Searching by Reinforcement Learning in Unstructured P2Ps,” Proc. of P2P/DAKS workshop (in conjunction with IEEE ICDCS), 2006.

283. Z. Jiang, R. Kline, J. Wu, and F. Dai, “A Practical Method to Form Energy Efficient Connected K-Coverage in Wireless Sensor Networks,” Proc. of WASN Workshop (in conjunction with IEEE ICDCS), 2006.

284. F. Li and J. Wu, “A Probabilistic Voting-Based Filtering Scheme in Wireless Sensor Networks,” Proc. of IWCMC 2006 Computer and Network Security Symposium, 2006.

285. Z. Jiang, J. Wu, and R. Kline, “Mobility Control with Local Views of Neighborhood in Mobile Networks,” Proc. of 2006 International Workshop on Networking, Architecture, and Storage, 2006, 9-14.

286. X. Li, J. Wu, and J. Xu, “Hint-Based Routing in WSN’s Using Scope Decay Bloom Filters,” Proc. of 2006 International Workshop on Networking, Architecture, and Storage, 2006, 111-118.

287. J. Sum, K. Ho, and J. Wu, “Some Analytical Results on a Localized Pruning Method for Connected Dominating Sets in MANETs,” Proc. of the 2nd International Workshop on Wireless, Ad Hoc, and Sensor Networks, 2006.

288. A. Srinivasan, J. Teitelbaum, and J. Wu, “DRBTS: Distributed Reputation-Based Beacon Trust System,” Proc. of the 2nd IEEE International Symposium on Dependable, Autonomic and Secure Computing (DASC), 2006.

289. C. Liu and J. Wu, “SWING: A Small World Iterative Navigation Greedy Routing Protocol in MANETs,” Proc. of IEEE Conference on Computer Communications and Networks (ICCCN), 2006.

290. S. Yang, J. Wu, and F. Dai, “Localized Movement-Assisted Sensor Deployment in Wireless Sensor Networks,” Proc. of the 1st International Workshop on Wireless Mesh Networks and Applications (WiMa) (in conjunction with IEEE MASS), 2006.

291. A. Agarwal, R. Shankar, J. Wu, and M. Mustafa, “A Deadlock Free Adaptive Routing Protocol for Network on Chip Architecture Design Methodology,” Proc. of the 4th IEEE Workshop on Embedded System and Real Time Multimedia, October 2006.

292. J. Wu and S. Yang, “SMART: A Scan-Based Movement Assisted Sensor Deployment Method in Wireless Sensor Networks,” Proc. of the 24th Annual Joint Conference of IEEE Communication and Computer Society (INFOCOM), March 2005.

293. D. Li, X. Lu, and J. Wu, “Fission E: A Scalable Constant Degree and Low Congestion DHT Scheme Based on Kautz Graph,” Proc. of the 24th Annual Joint Conference of IEEE Communication and Computer Society (INFOCOM), March 2005.

294. F. Dai and J. Wu, “Efficient Broadcasting in Ad Hoc Networks Using Directional Antennas,” Proc. of IFIP Networking 2005, May 2005.

295. F. Dai and J. Wu, “On Constructing k-Connected k-Dominating Set in Wireless Networks,” Proc. of the 19th International Parallel and Distributed Processing Symposium (IPDPS), April 2005.

296. F. Dai and J. Wu, “Proactive Route Maintenance in Ad Hoc Wireless Networks,” Proc. of the IEEE International Conference on Communications (ICC), May 2005.

297. S. Yang, J. Wu, and J. Cao, “Connected k-Hop Clustering in Ad Hoc Networks,” Proc. of International Conference on Parallel Processing (ICPP), June 2005.

298. D. Xiang, J. Sun, J. Wu, and K. Thulasiraman, “Fault-Tolerant Routing in Meshes/Tori Using Planarly Constructed Fault Blocks,” Proc. of International Conference on Parallel Processing (ICPP), June 2005.

299. X. Li and J. Wu, “Hybrid Searching Scheme in Unstructured P2P Networks,” Proc. of International Conference on Parallel Processing (ICPP), June 2005.

300. Z. Jiang, J. Wu, and D. Wang, “A New Fault Information Model for Fault-Tolerant Adaptive and Minimal Routing in 3-D Meshes,” Proc. of International Conference on Parallel Processing (ICPP), June 2005.

301. M. Lu, J. Wu, M. Cardei, and M. Li, “Energy-Efficient Connected Coverage of Discrete Targets in Wireless Sensor Networks,” Proc. of 2005 International Conference on Computer Networks and Mobile Computing (ICCNMC), 2005.

302. W. Jia, W. Tu, and J. Wu, “Hierarchical Multicast Tree Algorithms for Application Layer Mesh Networks,” Proc. of 2005 International Conference on Computer Networks and Mobile Computing (IC-CNMC), 2005. (Received the Best Paper Award)

303. H. Huang and J. Wu, “A Probabilistic Clustering Algorithm in Wireless Sensor Networks,” Proc. of the 62nd IEEE Semiannual Vehicular Technology Conference (VTC), Sept. 2005.

304. M. Cardei, J. Wu, M. Lu, and M. O. Pervaiz, “Maximum Network Lifetime in Wireless Sensor Networks with Adjustable Sensing Ranges,” Proc. of IEEE International Conference on Wireless and Mobile Computing, Networking and Communications (WiMob), Aug. 2005.

305. J. Wu, F. Dai, and S. Yang, “Iterative Local Solutions for Connected Dominating Set in Ad Hoc Wireless Networks,” Proc. of the 2nd IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), Nov. 2005.

306. S. Yang, F. Dai, M. Cardei, and J. Wu, “On Multiple Point Coverage in Wireless Sensor Networks,” Proc. of the 2nd IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), Nov. 2005. (Received the Best Paper Award)

307. C. Li, M. Ye, G. Chen, and J. Wu, “An Energy-Efficient Unequal Clustering Mechanism for Wireless Sensor Networks,” Proc. of the 2nd IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), Nov. 2005.

308. F. Dai and J. Wu, “View Consistency for Reliable Topology Control in Mobile Ad Hoc Networks,” Proc. of IEEE Global Telecommunications Conference (GLOBECOM), Nov. 2005.

309. Y. Chen, Y. He, J. Cao, and J. Wu, “A Greedy Algorithm for Capacity-Constrained Surrogate Placement in CDNs,” Proc. of IFIP International Conference on Network and Parallel Computing (NPC), Dec. 2005.

310. I. Jawhar and J. Wu, “A Dynamic Range Resource Reservation Protocol for QoS Support in Wireless Networks,” Proc. of the 3rd ACS/IEEE International Conference on Computer Systems and Applications (AICCSA), Jan. 2005.

311. X. Li and J. Wu, “Cluster-Based Intelligent Searching in Unstructured Peer-to-Peer Networks,” Proc. of the 3rd International Workshop on Mobile Distributed Computing (MDC’05) (in conjunction with ICDCS), June 2005.

312. Q. Fang, W. Jia, and J. Wu, “Available Bandwidth Detection with Improved Transport Control Algorithm for Heterogeneous Networks,” Proc. of the 3rd International Workshop on Mobile Distributed Computing (MDC) (in conjunction with ICDCS 2005), June 2005.

313. G. -J. Wang, J. Cao, L. F. Zhang, K. C. C. Chan, and J. Wu, “A Novel QoS Multicast Model in Mobile Ad Hoc Networks,” Proc. of the 7th Workshop on Advances in Parallel and Distributed Computational Models (APDCM), (in conjunction with IPDPS 2005), April 2005.

314. B. Wu, J. Wu, E. B. Fernandez, and S. Magliveras, “Secure and Efficient Key Management in Mobile Ad Hoc Networks,” Proc. of the 1st International Workshop on Systems and Network Security (SNS) (in conjunction with IPDPS), April 2005.

315. J. Wu, M. Cardei, F. Dai, and S. Yang, “Extended Dominating Set in Ad Hoc Networks Using Cooperative Communication,” Proc. of IFIP Networking 2005 (poster paper), June 2005.

316. Y. Wang and J. Wu, “Label Routing Protocol: A New Cross-Layer Protocol for Multi-Hop Wireless and Mobile Networks,” IEEE International Workshop on Heterogeneous Multi-Hop Wireless and Mobile Networks (MHWMN), 2005.

317. M. Cardei, J. Wu, and S. Yang, “Topology Control in Ad Hoc Wireless Networks with Hitch-Hiking,” Proc. of the 1st Annual IEEE Communications Society Conference on Sensor and Ad Hoc Communications and Networks (SECON), 2004.

318. Z. Jiang and J. Wu, “A Limited-Global Fault Information Model for Dynamic Routing in n-D Meshes,” Proc. of the 9th IEEE Workshop on Fault-Tolerant Parallel, Distributed and Network-Centric Systems (in conjunction with IPDPS), 2004.

319. I. Jawhar and J. Wu, “Quality of Service Routing in TDMA-Based Ad Hoc Networks,” Proc. of INFORMS Telecommunications Conference, 2004.

320. Q. Dai and J. Wu, “Construction of Power Efficient Routing Tree for Ad Hoc Wireless Networks using Directional Antenna,” Proc. of International Workshop on Wireless Ad Hoc Networking (WWAN 2004) (in conjunction with IEEE ICDCS), 2004.

321. J. Wu and S. Yang, “Coverage Issue in Sensor Networks with Adjustable Ranges,” Proc. of 2004 International Workshop on Mobile and Wireless Networking (MWN) (in conjunction with ICPP), Aug. 2004.

322. C. Huang, F. Dai, and J. Wu, “On-Demand Location-Aided QoS Routing in Ad Hoc Networks,” Proc. of International Conference on Parallel Processing (ICPP), Aug. 2004, 502-509.

323. I. Jawhar and J. Wu, “A Ticket-Based Expanding Ring Search Protocol for Peer-to-Peer Networks,” Proc. of the 8th World Multi-Conference on Systemic, Cybernetics and Informatics, July 2004.

324. M. Cardei, J. Wu, and S. Yang, “Low Power Hitch-Hiking Broadcast in Ad Hoc Wireless Networks,” Proc. of NSF Workshop on Theoretical and Algorithmic Aspect of Ad Hoc Wireless Networks, June 2004.

325. X. Li and J. Wu, “Hierarchical P2P Systems in a Small World,” Proc. of the 2nd International Latin American and Caribbean Conference for Engineering and Technology (LACCEI), June 2004.

326. F. Dai and J. Wu, “Efficient Broadcasting in Ad Hoc Wireless Networks Using Directional Antennas,” Proc. of NSF Workshop on Theoretical and Algorithmic Aspect of Ad Hoc Wireless Networks, June 2004.

327. M. Ye, C. F. Li, G. H. Chen, and J. Wu, “EECS: An Energy Efficient Clustering Scheme in Wireless Sensor Networks,” Proc. of IEEE International Workshop on Strategies for Energy Efficiency in Ad Hoc and Sensor Networks (IWSEEASN), April 2004.

328. W. Lou and J. Wu, “An Enhanced Message Exchange Mechanism in Cluster-Based Mobile Ad Hoc Networks,” Proc. of Parallel and Distributed Processing and Applications (ISPA), 2004, 223-232.

329. J. Wu and W. Lou, “Extended Multipoint Relays to Determine Connected Dominating Sets in MANET-S,” Proc. of the 1st Annual IEEE Communications Society Conference on Sensor and Ad Hoc Communications and Networks (SECON), 2004.

330. W. Lou and J. Wu, “A k-Hop Zone-Based Broadcast Protocol in Mobile Ad Hoc Networks,” Proc. of IEEE Global Telecommunications Conference (GLOBECOM), 2004.

331. J. Wu and F. Dai, “Mobility Management and Its Applications in Efficient Broadcasting in Mobile Ad Hoc Networks,” Proc. of the 23rd Annual Joint Conference of IEEE Communication and Computer Society (INFOCOM), 2004.

332. W. Lou and J. Wu, “Double-Covered Broadcast (DCS): A Simple Reliable Broadcast Algorithm in MANETs,” Proc. of the 23rd Annual Joint Conference of IEEE Communication and Computer Society (INFOCOM), 2004.

333. M. Agarwal, J. H. Cho, L. X. Gao, and J. Wu, “Energy Efficient Broadcast in Wireless Ad Hoc Networks with Hitch-Hiking,” Proc. of the 23rd Annual Joint Conference of IEEE Communication and Computer Society (INFOCOM), 2004.

334. J. Wu and F. Dai, “A Distributed Formation of a Virtual Backbone in Ad Hoc Networks using Adjustable Transmission Ranges,” Proc. of the 24th IEEE International Conference on Distributed Computing Systems (ICDCS), 2004, 372-379. (Nominated for the Best Paper Award)

335. I. Jawhar and J. Wu, “A Race-free Bandwidth Reservation Protocol for QoS Routing in Mobile Ad Hoc Networks,” Proc. of the 37th Hawaii International Conference on System Science (HICSS), Jan. 2004.

336. F. Dai and J. Wu, “Performance Analysis of Broadcast Protocols in Ad Hoc Networks Based on Self-Pruning,” Proc. of IEEE Wireless Communications and Networking Conference (WCNC), 2004.

337. J. Wu and Z. Jiang, “On Constructing the Minimum Orthogonal Convex Polygon in 2-D Faulty Meshes,” Proc. of IEEE International Parallel and Distributed Processing Symposium (IPDPS), 2004.

338. J. Wu and F. Dai, “Mobility-Sensitive Topology Control in Mobile Ad Hoc Networks,” Proc. of IEEE International Parallel and Distributed Processing Symposium (IPDPS), 2004.

339. J. Cao, J. Zhou, D. Chen, and J. Wu, “An Efficient Distributed Mutual Exclusion Algorithm Based on Relative Consensus Voting,” Proc. of IEEE International Parallel and Distributed Processing Symposium (IPDPS), 2004.

340. W. Lou and J. Wu, “A Reliable Broadcast Algorithm with Selected Acknowledgements in Mobile Ad Hoc Networks,” Proc. of IEEE Global Telecommunications Conference (GLOBECOM), 2003.

341. W. Lou and J. Wu, “A Double-Covered Broadcast Protocol for Mobile Ad Hoc Networks,” Proc. of IEEE Global Telecommunications Conference (GLOBECOM), 2003.

342. J. Shaikh, J. Solano, I. Stojmenovic, and J. Wu, “New Metrics for Dominating Set Based Energy Efficient Activity Scheduling in Ad Hoc Networks,” Proc. of WLN Workshop (in conjunction to IEEE Conference on Local Computer Networks), Oct. 2003, 726-735.

343. J. Wu, “An Enhanced Approach to Determine A Small Forward Node Set Based on Multipoint Relays,” Proc. of 2003 IEEE Semiannual Vehicular Technology Conference (VTC), Oct. 2003.

344. J. Wu and B. Wu, “A Transmission Range Reduction Scheme for Power-Aware Broadcasting in Ad Hoc Networks Using Connected Dominating Sets,” Proc. of 2003 IEEE Semiannual Vehicular Technology Conference (VTC), Oct. 2003.

345. Y. Dong and J. Wu, “Three Bluetree Formations for Constructing Efficient Scatternets in Bluetooth,” Proc. of the 7th Joint Conference on Information Sciences, Sept. 2003, 385-388.

346. Z. Jiang and J. Wu, “Fault-Tolerant Routing in Dual-Cube Networks,” Proc. of the 7th Joint Conference on Information Sciences, Sept. 2003, 389-392.

347. C. Yang and J. Wu, “A Dominating-Set-Based Routing in Peer-to-Peer Networks,” Proc. of the 2nd International Workshop on Grid and Cooperative Computing Workshop (GCC).

348. J. Wu, “Localized Algorithms and Their Applications in Ad Hoc Wireless Networks,” Proc. of International Symposium on Parallel and Distributed Processing and Applications, LNCS 2745, Springer, July 2003, 1-5.

349. J. Sum and J. Wu, “Asymptotic Analysis on An Unbounded Zero-One Knapsack with Discrete-Sized Objects,” Proc. of 2003 International Conference on Parallel and Distributed Processing Techniques and Applications (PAPTA), June 2003.

350. J. Sum and J. Wu, “On Two Allocation Methods For Web Services,” Proc. of 2003 International Conference on Parallel and Distributed Processing Techniques and Applications (PAPTA), June 2003.

351. J. Wu and F. Dai, “A Generic Distributed Broadcast Scheme in Ad Hoc Wireless Networks,” Proc. of the 23rd IEEE International Conference on Distributed Computing Systems (ICDCS), May 2003, 460-467. (Nominated for the Best Paper Award)

352. D. Zhou and J. Wu, “Survivable Multi-Level Ad-Hoc Group Operations,” Proc. of the International Workshop on Mobile and Wireless Networks, May 2003, 70-75.

353. Q. Dai and J. Wu, “Computation of Minimal Uniform Transmission Power in Ad Hoc Wireless Networks,” Proc. of the 2rd International Workshop on Assurance in Distributed Systems and Networks (in conjunction with ICDCS), May 2003, 680-684.

354. F. Dai and J. Wu, “Distributed Dominant Pruning in Ad Hoc Wireless Networks,” Proc. of IEEE International Conference on Communications (ICC), May 2003.

355. J. Wu and F. Dai, “An Extended Link Reversal Protocol in Dynamic Networks,” Proc. of the 17th IEEE International Parallel and Distributed Processing Symposium (IPDPS), April 2003.

356. J. Wu and F. Dai, “A Generic Broadcast Protocol in Ad Hoc Networks Based on Self-Pruning,” Proc. of the 17th IEEE International Parallel and Distributed Processing Symposium (IPDPS), April 2003.

357. Z. Jiang and J. Wu, “A Limited-Global Information Model for Fault-Tolerant Routing in Dual-Cubes,” Proc. of Workshop on Fault-Tolerant Parallel and Distributed Systems (in conjunction with IPDPS), April 2003.

358. W. Lou and J. Wu, “A Cluster-Based Backbone Infrastructure for Broadcasting in MANETS,” Proc. of Workshop on Wireless, Mobile, and Ad Hoc Networks (in conjunction with IPDPS), April 2003.

359. K. Du, J. Wu, and D. Zhou, “Chain-Based Protocols for Data Broadcasting and Gathering in Sensor Networks,” Proc. of Workshop on Parallel and Distributed Scientific and Engineering Computing with Applications (in conjunction with IPDPS), April 2003.

360. C. Wu and J. Wu, “On Self-Similarity and Hamiltonicity of Dual-Cubes,” Proc. of Workshop on Massively Parallel Processing (in conjunction with IPDPS), April 2003.

361. J. Wu and F. Dai, “Broadcasting in Ad Hoc Networks Based on Self-Pruning,” Proc. of the 22nd Annual Joint Conference of IEEE Communication and Computer Society (INFOCOM), March 2003.

362. W. Lou and J. Wu, “On Reducing Broadcast Redundancy in Mobile Ad Hoc Networks,” Proc. of the 36th Hawaii International Conference on System Science (HICSS), Jan. 2003.

363. F. Dai and J. Wu, “On Independently Finding Converging Paths in Internet,” Proc. of the 14th IASTED International Conference on Parallel and Distributed Computing and Systems, Nov. 2002.

364. D. Xiang, A. Chen, and J. Wu, “Fault-Tolerant Broadcasting in Hypercube Multicomputers Using Local Safety Information,” Proc. of the 9th International Conference on Parallel and Distributed Systems (ICPADS), Nov. 2002.

365. W. Lou and J. Wu, “Efficient Broadcast With Forward Node Set in Clustered Mobile Ad Hoc Networks,” Proc. of the 11th International Conference on Computer Communications and Networks (ICCCN), Oct. 2002.

366. J. Cao, X. Wang, and J. Wu, “A Mobile Agent Enabled Fully Distributed Mutual Exclusion Algorithm,” Proc. of the 6th IEEE International Conference on Mobile Agents (MA), Oct. 2002.

367. X. Chen, J. Wu, and X. Jia, “Prefix Routing in Wireless Ad-Hoc Networks of Mobile Stations,” Proc. of 2002 ISCA International Conference on Parallel and Distributed Computing Systems (PDCS), Oct. 2002, 66-71.

368. Z. Jiang and J. Wu, “A Limited Global Information Model for Dynamic Routing in 3-D Meshes,” Proc. of 2002 ISCA International Conference on Parallel and Distributed Computing Systems (PDCS), Oct. 2002, 343-348.

369. F. Dai, M. Dhawan, C. Wu, and J. Wu, “On Solutions to a Seating Problem,” Proc. of 2002 ISCA International Conference on Parallel and Distributed Computing Systems (PDCS), Oct. 2002, 398-403.

370. J. Wu and D. J. Wang, “Fault-Tolerant and Deadlock-Free Routing in 2-D Meshes Using Rectilinear-Monotone Polygonal Fault Blocks,” Proc. of International Conference on Parallel Processing (ICPP), 2002, 247-256.

371. J. Wu, “A Deterministic Fault-Tolerant and Deadlock-Free Routing Protocol in 2-D Meshes Based on Odd-Even Turn Model,” Proc. of the ACM International Conference on Supercomputing (ICS), June 2002, 67-76.

372. J. Wu and Z. Jiang. “Extended Minimal Routing in 2-D Meshes with Faulty Blocks,” Proc. of the 1st International Workshop on Assurance in Distributed Systems and Applications (in conjuction with ICDCS), 2002, 49-54.

373. J. Wu, B. Wu, and I. Stojmenovic, “Power-Aware Broadcasting and Activity Scheduling in Ad Hoc Wireless Networks Using Connected Dominating Sets,” Proc. of IASTED International Conference on Wireless and Optical Communication (WOC), 2002, 668-673.

374. J. Wu and F. Dai, “On Locality of Dominating Set in Ad Hoc Networks with Switch On/Off Operations,” Proc. of the 6th International Conference on Parallel Architectures, Algorithms, and Networks (I-SPAN), 2002, 85-90.

375. J. Wu and S. Olariu, “On Cost-Optimal Merge of Two Intransitive Sorted Sequences,” Proc. of Workshop on Advances in Parallel and Distributed Computational Models (in conjunction with IPDPS), 2002.

376. Z. Jiang and J. Wu, “A Limited-Global Fault Information Model for Dynamic Routing in 2-D Meshes,” Proc. of Workshop on Parallel and Distributed Scientific and Engineering Computing with Applications (in conjunction with IPDPS), 2002.

377. J. Wu. “An Enhanced Solution for Generating Loop-Free Routes in Dynamic Networks,” Proc. of the 6th Joint Conference on Information Sciences, March 2002, 261-265.

378. J. Wu, “An Extended Dynamic Source Routing Scheme in Ad Hoc Wireless Networks,” Proc. of the 35th Hawaii International Conference on System Sciences (HICSS), 2002.

379. J. Wu, M. Gao, and I. Stojmenovic, “On Calculating Power-Aware Connected Dominating Sets for Efficient Routing in Ad Hoc Wireless Networks,” Proc. of International Conference on Parallel Processing (ICPP), 2001, 346-356.

380. J. Cao, A. Chan, and J. Wu, “Achieving Replication Consistency Using Cooperating Mobile Agents,” Proc. of Workshop on Wireless Networks and Mobile Computing, 2001, 453-458.

381. J. Wu, “Broadcasting in Injured Hypercubes Using Directed Safety Levels,” Proc. of the 14th ISCA International Conference on Parallel and Distributed Computing Systems (PDCS), 2001, 7-12.

382. J. Wu and Z. Jiang, “Dynamic Fault-Tolerant Routing Cube-Based Multicomputers,” Proc. of IASTED International Conference on Parallel and Distributed Computing and Systems, 2001, 228-233.

383. J. Wu and L. Sheng “Deadlock-Free Prefix Multicasting in Irregular Networks,” Proc. of the 14th ISCA International Conference on Parallel and Distributed Computing Systems (PDCS), 2001, 25-30.

384. J. Sum, G. Young, H. Shen, and J. Wu, “Analysis on Extended Ant Routing Algorithms for Network Routing and Management,” Proc. of the 2rd International Conference on Parallel and Distributed Computing, Applications and Technologies (PDCAT), June 2001.

385. J. Wu, X. Lin, J. Cao, and W. Jia, “An Extended Fault-Tolerant Link-State Routing Protocol in the Internet,” Proc. of the 8th International Conference on Parallel and Distributed Systems (ICPADS), June 2001.

386. B. Roussev and J. Wu, “Client-Server Design Alternatives: Back to Pipes but with Threads,” Proc. of the 1st International Conference on Networking (ICN), Pascal Lorenz (Ed.), Lecture Notes in Computer Science 2094, July 2001, 854-866.

387. E. B. Fernandez, J. Wu, and D. Hancock, “A Design Method for Real-Time Object-Oriented Systems Using Communicating Real-Time State Machines,” Proc. of 2001 Information Resources Management Association Conference (IRMA), May 2001, 792-797.

388. J. Wu, “A Distributed Formation of Orthogonal Convex Polygons in Mesh-Connected Multicomputers,” Proc. of the 15th International Parallel and Distributed Processing Symposium (IPDPS), 2001.

389. S. Datta, I. Stojmenovic, and J. Wu “Internal Nodes and Shortcut Based Routing with Guaranteed Delivery in Wireless Networks,” International Workshop on Wireless Networks and Mobile Computing (in conjunction with ICDCS), 2001.

390. D. Xiang and J. Wu, “Reliable Unicasting in Faulty Hypercube Multicomputers Using Local Safety Information,” Proc. of the 4th International Conference on Algorithms and Architectures for Parallel Processing, Nov. 2000, 617-628.

391. J. Wu and H. Li, “Domination and Its Applications in Ad Hoc Wireless Networks with Unidirectional Links,” Proc. of International Conference on Parallel Processing (ICPP), Aug. 2000, 189-200.

392. J. Wu, “A Fault-Tolerant Adaptive and Minimal Routing Approach in n-D Meshes,” Proc. of International Conference on Parallel Processing (ICPP), Aug. 2000, 431-438.

393. D. Xiang and J. Wu, “Fault-Tolerant Multicasting in Hypercube Multicomputers Based on Local Safety Information,” Proc. of the 13th ISCA International Conference on Parallel and Distributed Computing Systems (PDCS), Aug. 2000, 529-534.

394. J. Wu, “A Simple Fault-Tolerant Adaptive and Minimal Routing Approach in 3-D Meshes,” Proc. of the 7th International Conference on Parallel and Distributed Systems (ICPADS), July 2000, 256-263.

395. J. Wu, “A Fault-Tolerant Adaptive and Minimal Routing Approach in 3-D Meshes,” Proc. of the 7th International Conference on Parallel and Distributed Systems (ICPADS), July 2000, 256-263.

396. J. Wu, F. Gao, Z. Li, and Y. Min, “Optimal Fault-Tolerant Routing in Hypercubes Using Extended Safety Vectors,” Proc. of the 7th International Conference on Parallel and Distributed Systems (ICPADS), July 2000, 264-271.

397. J. Wu, “On Sorting an Intransitive Total Ordered Set Using Semi-Heap,” Proc. of IEEE International Parallel and Distributed Processing Symposium (IPDPS), May 2000, 257-262.

398. B. Roussev and J. Wu, “A Java Runtime System for Parallel Execution of Multi-Threaded Java Applications on Networks on Workstations,” Proc. of the 11th IASTED International Conference on Parallel and Distributed Computing and Systems, Nov. 1999, 857-662.

399. X. Chen and J. Wu, “Path-Based Fault-Tolerant Multicasting in Mesh-Connected Multicomputers,” Proc. of the IASTED Parallel and Distributed Computing and Systems Conference, Nov. 1999, 72-77.

400. J. Wu and L. Sheng, “Deadlock-Free Routing in Irregular Networks Using Prefix Routing,” Proc. of the ISCA 12th International Conference on Parallel and Distributed Computing Systems, Aug. 1999, 424-430.

401. J. Wu and H. Li, “On Calculating Connected Dominating Set for Efficient Routing in Ad Hoc Wireless Networks,” Proc. of the 3rd International Workshop on Discrete Algorithms and Methods for Mobile Computing and Communications (Dial M) (in conjunction with MobiCom 1999), Aug. 1999, 7-14.

402. J. Wu and X. Chen, “A Fault-Tolerant and Time-Step Optimal Multicasting in 2-D Meshes,” Proc. of the 4th IEEE Workshop on Fault-Tolerant Parallel and Distributed Systems (in conjunction with IPPS/SPDP), April 1999.

403. F. He and J. Wu, “A Parallel Implementation of the Everglades Landscape Fire Model in Networks of Workstations,” Proc. of the 3rd Workshop on Communication, Architecture, and Applications for Network-based Parallel Computing (in conjunction with HPCA-5), Jan. 1999.

404. X. Chen and J. Wu, “Minimal Routing in 3-D Meshes using Extended Safety Levels,” Proc. of the 10th IASTED International Conference on Parallel and Distributed Computing and Systems, Nov. 1998.

405. F. He and J. Wu, “Parallel Simulation of a Forest Fire Model Using Express,” Proc. of the 4th International Conference on Computer Science & Informatics, Nov. 1998.

406. J. Wu and Y. Yang, “The Postal Network: A Versatile Interconnection Topology,” Proc. of International Conference on Parallel Processing (ICPP), Aug. 1998, 612-619.

407. J. Wu, “Fault Tolerant Adaptive and Minimal Routing in Mesh-Connected Multicomputers using Extended Safety Levels,” Proc. of the 18th International Conference on Distributed Computing Systems (ICDCS), May 1998, 428-435.

408. S. Cang and J. Wu, “Minimizing Total Communication Distance of a Broadcast on Mesh and Torus Networks,” Proc. of the 1st Merged Symposium IPPS/SPDP, April 1998, 11-17.

409. K. Huang, J. Wu, and E. B. Fernandez, “A Generalized Forward Recovery Checkpointing Scheme,” Proc. of 1998 Annual IEEE Workshop on Fault-Tolerant Parallel and Distributed Systems (in conjunction with IPPS/SPDP), April 1998.

410. J. Wu and E. B. Fernandez, “Circular Token: A Novel Synchronization Primitive,” Proc. of the 4th INFORMS Telecommunications Conference, March 1998.

411. V. Rancov and J. Wu, “Language Support for Context Driven Parallel Computations,” Proc. of the 10th IASTED International Conference on Parallel and Distributed Computing and Systems, Oct. 1997, 273-278,

412. Y. Li, J. Wu, and S. Q. Zheng, “An Interconnection Network Based on the Dual of a Hypercube,” Proc. of the 10th ISCA International Conference on Parallel and Distributed Computing Systems (PDCS), Oct. 1997, 263-268.

413. J. Wu, E. B. Fernandez, and Y. Luo, “Embedding of Binomial Trees in Hypercube Multiprocessors with Link Faults,” Proc. of International Conference on Parallel Processing (ICPP), August 1997, 96-99.

414. J. Wu and G. Guo, “Fault Tolerance Measures in Generalized Hypercubes Based on Forbidden Fault Sets,” Proc. of the 2nd Annual Workshop on Fault-Tolerant Parallel and Distributed Systems (in conjunction with IPPS), April 1997, 202-217.

415. J. Wu, V. Rancov, and A. Stoichev, “Parallel Computations on LAN-Connected Workstations,” Proc. of the 10th IASTED International Conference on Parallel and Distributed Computing and Systems, Oct. 1996, 193-197.

416. R. Yehia and J. Wu, “Extensions to Integrated-Snooper Fault-Tolerant Distributed Share Memory,” Proc. of ISCA International Conference on Computer Applications in Industry and Engineering, Dec. 1996, 23-27.

417. V. Rancov and J. Wu, “Implementing Context-Driven Parallel Computations,” Proc. of ISCA International Conference on Computer Applications in Industry and Engineering, Dec. 1996, 5-8.

418. J. Wu, “An Optimal Routing Policy for Mesh-Connected Topologies,” Proc. of International Conference on Parallel Processing (ICPP), August 1996, I-267-I-271.

419. J. Wu, “A Fault-Tolerant Distributed Unicasting Scheme for Cube-Based Multicomputers,” Proc. of the 16th International Conference on Distributed Computing Systems (ICDCS), May 1996, 166-174.

420. R. France, J. Wu, M. M. Larrondo-Petrie, and J. M. Bruel, “A Tale of Two Case Studies: Using Integrated Methods to Support Rigorous Requirements Specification,” Proc. of the International Workshop on Formal Methods Integration, March 1996.

421. S. Q. Zheng and J. Wu, “Dual of a Complete Graph as an Interconnection Network,” Proc. of the 7th IEEE Symposium on Parallel and Distributed Processing (SPDP), 1996, 433-442.

422. J. Wu, J. Pan, K. Gopu, F. Sklar, and Y. Wu, “Domain Partition Scheme for the Everglades Landscape Model (ELM),” Proc. of the 4th Software Engineering Research Forum, Nov. 1995, 53-59.

423. H. Qian, E. B. Fernandez, and J. Wu, “A Combined Functional and Object-Oriented Approach to Software Design,” Proc. of the 1st IEEE International Conference on Engineering of Complex Computer Systems, Nov. 1995, 167-174.

424. J. Wu, “Extended Fibonacci Cubes,” Proc. of the 7th IEEE Symposium on Parallel and Distributed Processing (SPDP), Oct. 1995, 248-251.

425. J. Wu, E. B. Fernandez, and D. Dai, “Optimal Fault-Secure Scheduling,” Proc. of the 8th ISCA International Conference on Parallel and Distributed Computing, Sept. 1995, 603-608.

426. H. Qian and J. Wu, “Enhanced Fibonacci Cubes,” Proc. of the 8th ISCA International Conference on Parallel and Distributed Computing (PDCS), Sept. 1995, 521-526.

427. H. Qian and J. Wu, “Unicast, Multicast, and Broadcast in Enhanced Fibonacci Cubes,” Proc. of the 4th International Conference on Computer Communications and Networks, Sept. 1995.

428. J. Wu, “Reliable Communication in Cube-Based Multicomputers,” Proc. of the International Workshop on High Performance Computing, July 1995.

429. J. Wu and H. Zhuang, “A Petri-Net-Based Collision and Deadlock Avoidance Scheme for FMS,” Proc. of the 1995 INRIA/IEEE Conference on Emerging Technologies and Factory Automation, Vol. 2, 1995, 511-520.

430. J. Wu, “Unicasting in Faulty Hypercubes Using Safety Levels,” Proc. of International Conference on Parallel Processing (ICPP), 1995, III-132-III-136.

431. R. France and J. Wu, “Towards Systematic Formalization of Informal Requirements Specifications: A Case Study on the Vending Machine Problem,” Proc. of ISCA International Conference on Computer Applications in Industry and Engineering, Dec. 1994.

432. H. Qian and J. Wu, “Multitriangle: A New Interconnection Network,” Proc. of the 7th ISCA International Conference on Parallel and Distributed Computing Systems (PDCS), October 1994, 755-760.

433. J. Wu, “Identification of Buddy Nodes in Faulty Hypercubes,” Proc. of the 6th ISATED-ISMM International Conference on Parallel and Distributed Computing and System, October 1994.

434. E. B. Fernandez, J. Wu and M. H. Fernandez, “User Group Structures in Object-Oriented Database Authorization,” Proc. of the 8th Annual IFIP Working Group 11.3 Working Conference on Database Security, August 1994, 57-76.

435. J. Wu “Dimension-Exchange-Based Load Balancing on Injured Hypercubes,” Proc. of International Conference on Parallel Processing (ICPP), August 1994, III-88-III-91.

436. L. Brown and J. Wu, “Dynamic Snooping in a Fault-Tolerant Distributed Shared Memory,” Proc. of the 14th International Conference on Distributed Computing Systems (ICDCS), June 1994, 218-226. (Nominated for the Best Paper Award)

437. L. Brown and J. Wu, “Optimal Triple Modular Redundancy Embeddings in the Hypercube,” Proc. of Euromicro Conference on Massively Parallel Computing Systems, May 1994, 600-610.

438. H. M. Pendyala and J. Wu, “Prefix Computation on Extended Fibonacci Cubes,” Congressus Numerantium, 103, 1994, 202-214.

439. H. Qian and J. Wu, “Embedding of Meshes of Trees into Hypercubes,” Congressus Numerantium, 105, 1994, 161-180.

440. J. Wu and K. Yao, “Multicasting in Injured Hypercubes Using Limited Global Information,” Proc. of the 5th IEEE Symposium on Parallel and Distributed Processing (SPDP), Dec. 1993, 548-555.

441. Y. Luo and J. Wu, “Gray-Code-Based Optimal Cube-Connected Cube Networks,” Congressus Numerantium, 95, Dec. 1993, 203-211.

442. L. Brown and J. Wu, “Recovery in Fault-Tolerant Distributed Shared Memories,” Proc. of the 4th Software Engineering Research Forum, Nov. 1993, 309-312.

443. R. France and J. Wu, “Combining SA/RT and Petri-Nets,” Proc. of Software Engineering Research Forum, Nov. 1993, 319-322.

444. J. Wu and T. Wu, “Vector-Matrix-Vector Multiplication on Cube-Connected-Cubes,” Proc. of the 3rd International Conference for Young Computer Scientists, July 1993, 1-4.

445. J. Wu, Y. Wang, and E. B. Fernandez, “A Uniform Approach to Software and Hardware Fault Tolerance,” Proc. of Euromicro Workshop on Parallel and Distributed Processing, Jan. 1993, 409-416.

446. K. Huang and J. Wu, “Optimal Fault-Tolerant Ring Networks,” Congressus Numerantium, 91, Nov. 1992, 43-54.

447. Z. Li and J. Wu, “A Novel Method of Function Decomposition,” Congressus Numerantium, 90, Nov. 1992, 161-179.

448. J. Wu and E. B. Fernandez, “Broadcasting in Faulty Hypercubes,” Proc. of the 11th IEEE Symposium on Reliable Distributed Systems, Oct. 1992, 122-129.

449. K. Huang and J. Wu, “Balanced Hypercubes,” Proc. of International Conference on Parallel Processing (ICPP), August 1992, I-80-I-84.

450. J. Wu, “Efficient Fault-Tolerant Nonredundant Broadcasting in Hypercube Multicomputers,” Proc. of International Conference on Parallel Processing (ICPP), August 1992, III-23-III-36.

451. R. Zhang, E. B. Fernandez, and J. Wu, “Parallel Processing in Robotics,” Proc. of the 2nd International Conference on Automation Technology, Vol. 3, July 1992, 47-53.

452. K. Venkatesh, O. Masory, and J. Wu, “Simulation and Robot Scheduling of a Just-in-Time Flexible Automated System Using Timed Petri Nets,” Proc. of the 2nd International Conference on Automation Technology, Vol. 4, July 1992, 73-79.

453. J. M. Krull, J. Wu, and A. M. Molina, “Evaluation of a Fault-Tolerant Distributed Broadcast Algorithm in Hypercube Multicomputers,” Proc. of the 20th ACM Annual Computer Science Conference, March 1992.

454. J. Wu and E. B. Fernandez, “Broadcasting in Faulty Cube-Connected Cycles with Minimum Recovery Time,” Proc. of Parallel Processing: CONPAR92, Lecture Notes in Comp. Science 634, Springer Verlag, 1992, 833-834.

455. M. Gangadhar, J. Wu, and E. B. Fernandez, “A Real Time Fault-Tolerant Scheduling Method Based on Comparisons,” Proc. of IFAC Workshop on Algorithms and Architectures for Real-Time Control, Dec. 1991.

456. M. M. Zhang and J. Wu, “Modeling and Analysis of a Fault-Tolerant System,” Proc. of International Conference for Young Computer Scientists, July 1991.

457. Z. Li and J. Wu, “A Fully Fault-Tolerant Multidestination Routing Scheme in Hypercube Multiprocessors,” Proc. of International Conference on Parallel Processing (ICPP), 1991, III-290-III-291.

458. J. Wu, “On the Decomposition of a Boolean Function into Two Unate Functions,” Congressus Numerantium, 77, December 1990, 205-211.

459. J. Wu, “A Fault-Tolerant Task Scheduling Method for Parallel Processing Systems,” Proc. of ISMM International Conference on Parallel and Distributed Computing and Systems, October 1990.

460. R. Zhang, E. B. Fernandez, and J. Wu, “A Parallel Implementation of Robot Control Equations on MS T414 Transputers,” Proc. of North American Transputer Users Group, 1990 fall meeting.

461. J. Wu and E. B. Fernandez, “Fault-Tolerant Distributed Broadcast Algorithms on Cube-Connected-Cycles Networks,” Proc. of the 1990 Canadian Conference on Electrics and Computer Engineering, September 1990, 51.4.1-51.4.4.

462. J. Wu, “Automatic Programming for Array Manipulation,” Proc. of the 3rd Florida AI Research Symposium, April 1990, 207-209.

463. J. Wu and E. B. Fernandez, “Fault-Tolerant Multiprocessor Networks Through an Extended G-Network,” Proc. of PARBASE-90, March 1990, 216-220.

464. J. Wu and E. B. Fernandez, “The Design of Decentralized Control Algorithms in Distributed Systems,” Proc. of ISMM International Conference on Intelligent Distributed Processing, December 1989, 94-97.

465. J. Wu and V. Chandra, “A Situation Adaptive Algorithm for Test Minimization for Applications in Safety of Robot Systems,” Proc. of Conference on Recent Advances in Robotics, May 1988, 36.

466. J. Wu and E. B. Fernandez, “A Weighted Tree Model for Algorithms with Boolean Vector Operations,” Proc. of the 20th Southeastern Symposium on System Theory, March 1988, 150-154.

467. G. Zhu, J. Wu, and H. Pan, “A Reproductive Prolog Realization System,” Proc. of Annual Shanghai Computer Society Conference, 1985.

4.1.3 Books and Book Chapters (4 books and 36 book chapters)

1. J. Wu and Y. Wang, Opportunistic Mobile Social Networks, Taylor & Francis, 2014.

2. J. Wu, Distributed Systems Design, CRC Press, 1999. (Chinese Edition, China Machine Press, 2001)

3. J. Wu, Handbook of Theoretical and Algorithmic Aspects of Sensor, Ad Hoc Wireless and Peer-to-Peer Networks, Auerbach Publications, 2006. (Flyer)(Preface) (Part I: Ad Hoc Wireless Networks) (Part II: Sensor Networks) (Part III: Peer-to-Peer Networks).

4. C. C. Tan and J. Wu, Security in RFID Networks and Communications, in Wireless Network Security: Theories and Applications, L. Chen and J. Ji (eds), Springer, 2013.

5. W. Chang and J. Wu, A New View of Privacy in Social Networks: Strengthening Privacy During Propagation, is accept to appear in Handbook of Research on Modern Cryptographic Solutions for Computer and Cyber Security, IGI Global, 2015.

6. W. Zhang, J. Wu, and Y. Lin, Secure and Privacy Preserving Keyword Search Over the Large Scale Cloud Data, is accept to appear in Handbook of Research on Modern Cryptographic Solutions for Computer and Cyber Security, IGI Global, 2015.

7. W. Chang and J. Wu, A Survey of Sybil Attacks in Networks, in Sensor Networks for Sustainable Development, M. Ilyas (ed), CRC Press, 2014.

8. P. Ostovari, J. Wu and A. Khreishah, Network Coding Techniques for Wireless and Sensor Networks, in The Art of Wireless Sensor Networks, Vol. 1, 129-162, H. M. Ammari (ed), Springer, 2014.

9. X. Li and J. Wu, Utility-Based Routing in Wireless Sensor Networks, in The Art of Wireless Sensor Networks, Vol. 1, 235-270, H. M. Ammari (ed), Springer, 2014.

10. B. Hurd, C. C. Tan, and J. Wu, The Future in Mobile Multicore Computing, in Scalable Computing: Theory and Practice, S. U. Khan, L. Wang, and A. Y. Zomaya (eds), John Wiley & Sons, 2014.

11. D. Li and J. Wu, Energy-Aware Scheduling on Multiprocessor Platforms, Springer Briefs in Computer Science, 2013, ISBN 978-1-4614-5223-2, pp. i-vii, 1-59.

12. H. Quan, J. Wu, and J. Shi, Online Social Networks and Social Network Services: A Technical Survey, in Handbook of Pervasive Communication, M. Ilyas and H. Mouftah (eds), CRC Press, 2010.

13. J. Chen and J. Wu, A Survey on Cryptograph Applied to Secure Mobile Ad Hoc Networks and Wireless Sensor Networks, in Handbook of Research on Developments and Trends in Wireless Sensor Networks: From Principle to Practice, H. Jin and W. Jiang (eds), IGI Global, 2008.

14. B. Wu, M. Cardei, and J. Wu, A Survey of Key Management in Mobile Ad Hoc Networks, Chapter in Handbook of Research on Wireless Security, J. Zheng, Y. Zhang, and M. Ma (eds.), Idea Group Inc, 2008.

15. Y. Shao, C. Liu, and J. Wu, Delay Tolerant Networks in VANETs, in the Handbook on Vehicular Networks, M. Weigle and S. Olariu (eds.), Taylor & Francis, March 2009.

16. M. O. Pervaiz, M. Cardei, and J. Wu, Security in Wireless Local Area Networks, Security in Distributed and Networking Systems, Y. Xiao and Y. Pan (eds.), World Scientific Publishing Co., 2008.

17. B. Wu, J. Chen, J. Wu, and M. Cardei, A Survey on Attacks and Countermeasures in Mobile Ad Hoc Networks, in Wireless/Mobile Network Security, Y. Xiao, X. Shen, and D. -Z. Du (eds.), Springer, 2007.

18. T. Anantvalee and J. Wu, A Survey on Intrusion Detection in Mobile Ad Hoc Networks, in Wireless/Mobile Network Security, Y. Xiao, X. Shen, and D. -Z. Du (eds.), Springer 2007.

19. A. Srinivasan and J. Wu, A Survey on Secure Localization in Wireless Sensor Networks, in Encyclopedia of Wireless and Mobile Communications, B. Furht (ed.), CRC Press, Taylor and Francis Group, 2007.

20. F. Li, A. Srinivasan, and J. Wu, Security Issues in Sensor Surveillance, in theHandbook on Mobile Ad Hoc and Pervasive Communications, M. Denko, American Scientific Publishers, 2007.

21. I. Cardei, C. Liu, and J. Wu, Routing in Wireless Networks with Intermittent Connectivity, in the Encyclopedia of Wireless and Mobile Communications, B. Furht (ed.), CRC Press, Taylor and Francis Group, 2007.

22. A. Srinivasan, J. Teitelbaum, H. Liang, J. Wu, and M. Cardei, Reputation and Trust-Based Systems for Ad Hoc and Sensor Networks, in Algorithms and Protocols for Wireless Ad Hoc and Sensor Networks, A. Boukerche (ed.), Wiley & Sons, 2006.

23. X. Li and J. Wu, Searching Techniques in Peer-to-Peer Networks, in Handbook of Theoretical and Algorithmic Aspects of Ad Hoc, Sensor, and Peer-to-Peer Networks, J. Wu (ed.), Auerbach Publications, 2006, 613-642.

24. J. Wu and F. Dai, Mobility Control and Its Applications in Mobile Ad Hoc Networks, in Handbook of Algorithms for Mobile and Wireless Networking and Computing, A. Boukerche and I. Chlamtac (eds.), CRC Press, 2005.

25. M. O. Pervaiz, M. Cardei, and J. Wu, Routing Security in Ad Hoc Wireless Networks, in Network Security, S. Huang, D. MacCallum, and D. -Z. Du (eds.), Springer, 2005.

26. D. Simplot-Ryl, I. Stojmenovic, and J. Wu, Energy Efficient Broadcasting, Activity Scheduling and Area Coverage in Sensor Networks, in Handbook of Sensor Networks: Algorithms and Architectures, I. Stojmenovic (ed.), Wiley, 2005, 343-379.

27. S. Yang and J. Wu, New Technologies of Multicasting in MANETs, in Design and Analysis of Wireless Networks, Y. Pan and Y. Xiao (eds), Nova Science Publishers, 2005.

28. I. Jawhar and J. Wu, Quality of Service Routing in Ad Hoc Networks, Resource Management in Wireless Networking, M. Cardei, and D. -Z. Du (eds.), Springer, 2005.

29. W. Lou and J. Wu, Localized Broadcasting in Mobile Ad Hoc Networks Using Neighbor Designation , Handbook of Mobile Computing, I. Maghoub and M. Ilyas (eds.), CRC Press, 2005, 663-686.

30. M. Cardei and J. Wu, Energy-Efficient Coverage Problems in Wireless Ad Hoc Sensor Networks, in Handbook of Sensor Networks, M. Ilyas (ed.), CRC Press, 2004, 19.1-19.20.

31. I. Stojmenovic and J. Wu, Broadcasting and Activity-Scheduling in Ad Hoc Networks, in Ad Hoc Networking, S. Basagni, M. Conti, S. Giordano, and I. Stojmenovic (eds.), IEEE Press, June 2004, 205-229.

32. I. Jawhar and J. Wu, QoS Routing in Ad Hoc Networks, in Resource Management in Wireless Networking, M. Cardei, I. Cardei, and D. -Z. Du (eds.), Springer, 2004, 365-400.

33. X. Chen and J. Wu, Multicasting Techniques in Mobile Ad Hoc Networks, in Handbook of Ad Hoc Wireless Networks, M. Ilyas and B. Furht (eds.), CRC Press, 2002.

34. J. Wu, Dominating-Set-Based Routing in Ad Hoc Wireless Networks, in Wireless Networks and Mobile Computing, I. Stojmenovic (ed.), John Wiley & Sons, 2002, 425-450.

35. E. B. Fernandez, J. Wu, and D. R. Hancock, A Design Method for Real-Time Object-Oriented Systems Using Communicating Real-Time State Machines, in Successful Software Reengineering, S. Valenti (ed.), IRM Press, 2002, 171-185.

36. Y. Pan, C. Qiao, Y. Yang, and J, Wu, Recent Developments in Optical Multistage Networks, in Optical Networks- Recent Advances, L. Ruan and D. -Z. Du (eds.), Kluwer Academic Publisher, 2001, 151-185.

37. B. Roussev and J. Wu, Task Scheduling on NOWs using Lottery-Based Workstealing, in Annual Review of Scalable Computing, Y. C. Kwong (ed.), World Scientific, 2001, 63-84.

38. J. Wu and X. Chen, A Fault-Tolerant and Time-Step Optimal Multicasting in 2-D Meshes, in Dependable Network Computing, Chapter 13, D. R. Avresky (ed.), Kluwer Academic Publishers, 2000, 279-312.

39. J. Wu, Broadcasting in Hypercubes Using Safety Levels, in Fault-Tolerant Parallel and Distributed Systems, D. Avresky (ed.), IEEE Computer Society Press, 1995, 70-76.

40. J. Wu and E. B. Fernandez, The Extended G-Network, A Fault-Tolerant Interconnection for Multiprocessors, in Parallel Architectures, N. Rishe, S. Navathe, and D. Tal (eds.) IEEE Computer Society Press, 1991, 250-259.

4.1.4 Magazine Articles (2)

1. J. Wu, “Some Reflections on C&I Education,” CCF China National Computer Conference, Invited Keynote, Communication of CCF, (in Chinese), Vol. 8, No. 1, 2012, 40-43.

2. J. Wu and F. Dai, “Mobility Control and Its Applications in Mobile Ad Hoc Networks,” IEEE Network, July-August 2004, 30-35.

4.1.5 Special Issues (8)

1. K. Kant, S. V. Pemmaraju, K. M. Sivalingam, and J. Wu, “Distributed Computing and Networking,” Communication Networks, Vol. 5935, 2010.

2. J. Wu and M. Cardei, “Theoretical and Algorithmic Aspects of Sensor, Ad Hoc Wireless, and Peerto-Peer Networks,” Journal of Parallel and Distributed Computing, Vol. 62, No. 2, 2005.

3. J. Wu and I. Stojmenovic, “Ad Hoc Networks,” IEEE Computer, February 2004.

4. K. Nakano and J. Wu, “Advances in Parallel and Distributed Computational Models,” a special issue in International Journal of Foundations of Computer Science, 2004.

5. K. Nakano and J. Wu, “Advances in Parallel and Distributed Computational Models,” a special issue in International Journal of Foundations of Computer Science, 2003.

6. J. Wu and S. Olariu, “Routing in Computer and Communication Networks,” Journal of Parallel and Distributed Computing, September 2001.

7. J. Wu and D. P. Agrawal, “Challenges in Designing Fault-Tolerant Routing in Networks,” IEEE Transactions on Parallel and Distributed Systems, November 1999.

8. Associate Editor, Modeling and Simulation, Technical Communication Services, Vol. 23, 1993.

4.1.6 Semi-Refereed Publications (43)

1. J. Wu, “A New Algorithmic and Graph Model for Dynamic Computer Networks,” Proc. of the 43th Southeastern International Conference on Combinatorics, Graph Theory, and Computing, March 2010, (abstract only).

2. L. Sheng, Y. Wang, and J. Wu, “A Formula for the Optimal Shortest-Path Routing in 2-D Mesh-Connected Multicomputers,” Proc. of the 33rd Southeastern International Conference on Combinatorics, Graph Theory, and Computing, March 2002, (abstract only).

3. W. Lou and J. Wu, “A Multipath Routing Protocol for Unidirectional Networks,” Proc. of 2001 International Conference on Parallel and Distributed Processing Techniques and Applications (PDP-TA’2001), June 2001, 2021-2027.

4. Z. Jiang and J. Wu, “A Fault-Tolerant Broadcasting in 2-D Wormhole-Routed Meshes,” Proc. of 2001 International Conference on Parallel and Distributed Processing Techniques and Applications (PDPTA’2001), June 2001, 2028-2034.

5. J. Sum, J. Wu, and H. Shen, “Communication Bottleneck in Agent-Mediated eMarketplace,” Proc. of 2001 International Conference on Parallel and Distributed Processing Techniques and Applications (PDPTA’2001), June 2001, 2015-2020.

6. C. Wu and J. Wu, “An Application of PVM in Parallel Computing,” Proc. of Workshop on Data Intensive Parallel and Distributed Computation (in conjunction with PDPTA’2001), 1604-1609.

7. J. Wu and L. Sheng, “An Efficient Sorting Algorithm for a Sequence of Kings in a Tournament,” Technical Report, DIMACS, 2000-2012.

8. W. Lou, J. Wu, and L. Sheng, “On the Existence of a Sorting Sequence of Kings in A Tournament,” Proc. of the 31st Southeastern International Conference on Combinatorics, Graph Theory, and Computing, March 2000, (abstract only).

9. J. Wu, “Dominating and Absorbent Set in Directed Graphs and Its Applications in Wireless Networks,” Proc. of the 31st Southeastern International Conference on Combinatorics, Graph Theory, and Computing, 2000.

10. J. Wu, “On Finding a Hamilton Path in a Tournament Using Semi-Heap,” Proc. of the 30th Southeastern International Conference on Combinatorics, Graph Theory, and Computing, March 1999, (abstract only).

11. J. Wu and H. Li, “On Calculating Connected Dominating Set for Efficient Routing in Ad Hoc Wireless Networks,” Proc. of the 30th Southeastern International Conference on Combinatorics, Graph Theory, and Computing, March 1999, (abstract only).

12. X. Chen and J. Wu, “Fault-Tolerant Multicasting in 3-D Meshes,” Proc. of the 30th Southeastern International Conference on Combinatorics, Graph Theory, and Computing, March 1999, (abstract only).

13. L. Sheng and J. Wu, “On Minimum Number of Channels Required for Deadlock-Free Wormhole Routing,” Proc. of the 30th Southeastern International Conference on Combinatorics, Graph Theory, and Computing, 1999.

14. J. Wu, “On Constructing Faulty Orthogonal Convex Polygons in 2-D meshes,” Proc. of the 29th Southeastern International Conference on Combinatorics, Graph Theory, and Computing, March 1998, (abstract only).

15. C. Yang and J. Wu, “Enhanced WK-Recursive Networks,” Proc. of the 29th Southeastern International Conference on Combinatorics, Graph Theory, and Computing, March 1998, (abstract only).

16. X. Chen and J. Wu, “Minimal Routing in 3-D Meshes Using Extended Safety Levels,” Proc. of the 29th Southeastern International Conference on Combinatorics, Graph Theory, and Computing, March 1998, (abstract only).

17. J. Wu, “Optimal Routing in 2-Dimensional Meshes with Faulty Blocks,” Proc. of the 27th Southeastern International Conference on Combinatorics, Graph Theory and Computing, March 1997, (abstract only).

18. R. Yehia and J. Wu, “Extensions To Integrated-Snooper Fault Tolerant Distributed Share Memory,” Proc. of ISCA International Conference on Computer Applications in Industry and Engineering, December 1996, 23-27.

19. V. Rancov and J. Wu, “Implementing Context-Driven Parallel Computations,” Proc. of ISCA International Conference on Computer Applications in Industry and Engineering, December 1996, 5-8.

20. J. Wu, J. Pan, K. Gopu, F. Sklar, and Y. Wu, “Domain Partition Scheme for the Everglades Landscape Model (ELM),” Proc. of the 4th Software Engineering Research Forum, November 1995, 53-59.

21. J. Wu, “Optimal Broadcasting in Hypercubes with Link Faults Using Limited Global Information,” Proc. of the 4th ORSA Telecommunication Conference, April 1995.

22. J. Wu, “A Trail-Based Deadlock-Free Routing Scheme in Wormhole-Routed Networks,” Proc. of IEEE Southcon’95, March 1995, 285-288.

23. C. Yang and J. Wu, “A General Deadlock-Free Routing Algorithm for Wormhole-Routed Torus Networks,” Proc. of the 26th Southeastern International Conference on Combinatorics, Graph Theory and Computing, March 1995, (abstract only).

24. J. Wu, “Broadcasting in Injured Hypercubes Using Limited Global Information,” Proc. of IEEE Southeastcon’94, April 1994, 343-347.

25. H. More and J. Wu, “Throughput Improvement Through Dynamic Load Balancing,” Proc. of IEEE Southeastcon’94, April 1994, 339-342.

26. K. Huang and J. Wu, “Consistent Recoverable Tree Embedding in Balanced Hypercubes,” Proc. of the 25th Southeastern International Conference on Combinatorics, Graph Theory and Computing, March 1994, (abstract only).

27. H. Pendyala and J. Wu, “Computing Prefix Sums in Extended Fibonacci Cubes,” Proc. of the 25th Southeastern International Conference on Combinatorics, Graph Theory and Computing, March 1994, (abstract only).

28. Y. Luo and J. Wu, “Gray-Code-Based Optimal Cube-Connected Cube Networks,” Proc. of the 24th Southeastern International Conference on Combinatorics, Graph Theory and Computing, February 1993 (abstract only).

29. K. Huang and J. Wu, “Hypercube Embedding in Balanced Hypercubes,” Proc. of the 24th Southeastern International Conference on Combinatorics, Graph Theory and Computing, February 1993, (abstract only).

30. J. Wu and H. More, “Load Balancing and Fault Tolerance in Hypercube Systems,” Proc. of the 24th Southeastern International Conference on Combinatorics, Graph Theory and Computing, February 1993, (abstract only).

31. H. Zhuang, J. Wu, and R. Sudhakar, “A Parallel Algorithm for Stereo Motion Estimation,” Proc. of the 5th Annual Conference on Recent Advances in Robotics, June 1992, 816-818.

32. J. Wu and X. -H. Sun, “Optimal Cube-Connected Cubes,” Proc. of the 3rd Annual Pittsburgh Conference on Modeling and Simulation, Vol. 23, No. 3, May 1992, 1463-1470.

33. L. Brown, J. Wu, and E. B. Fernandez, “Fault-Tolerant Task Scheduling for Hypercube Multiprocessors,” Proc. of the 23rd Annual Pittsburgh Conference on Modeling and Simulation, Vol. 23, No. 3, May 1992, 1449-1456.

34. K. Huang and J. Wu, “Optimal Fault-Tolerant Ring Networks,” Proc. of the 23rd Southeastern International Conference on Combinatorics, Graph Theory and Computing, February 1992, (abstract only).

35. Z. Li and J. Wu, “A Novel Method of Function Decomposition,” Proc. of the 23rd Southeastern International Conference on Combinatorics, Graph Theory and Computing, February 1992, (abstract only).

36. J. Wu, “Software Fault Tolerance Using Hierarchical N-Version Programming,” Proc. of IEEE Southeastcon ’91, 1991, 241-247.

37. J. Wu and E. B. Fernandez, “Petri Net Modeling Techniques,” Proc. of the 21st Pittsburgh Conference on Modeling and Simulation, May 1990, 1311-1317.

38. J. Wu and E. B. Fernandez, “A Scheduling Approach for Communicating Tasks,” Proc. of the 22nd Southeastern Symposium on System Theory, March 1990, 91-97.

39. J. Wu, “On the Decomposition of A Boolean Function into Two Unate Functions,” Proc. of the 21st Southeastern International Conference on Combinatorics, Graph Theory and Computing, February 1990, (abstract only).

40. J. Wu and E. B. Fernandez, “Scheduling of Real-Time Tasks,” Proc. of the 20th Southeastern International Conference on Combinatorics, Graph Theory and Computing, February 1989, (abstract only).

41. J. Wu and V. Chandra, “A Situation Adaptive Algorithm for Test Minimization for Applications in Safety of Robot Systems,” Proc. of Conference on Recent Advances in Robotics, May 1988, 36.

42. J. Wu, “A Method of Achieving Several Goals Simultaneously,” Proc. of the 1st Artificial Intelligence Research Symposium, May 1988, 35-39.

43. J. Wu and E. B. Fernandez, “A Weighted Tree Model for Algorithms with Boolean Vector Operations,” Proc. of the 20th Southeastern Symposium on System Theory, March 1988, 150-154.

4.2 Keynote Speeches (25)

1. Keynote Speaker: Algorithmic Crowdsourcing: Current Status and Future Perspective, IPCCC, 2014.

2. Keynote Speaker: Resource Allocation in Network Virtualization, WSDIF (in conjunction with IEEE MASS), 2014.

3. Keynote Speaker: Social-Aware Routing in Delay Tolerant Networks, DySON (in conjunction with IEEE INFOCOM), 2014.

4. Keynote Speaker: Algorithmic Crowdsourcing and Applications in Big Data, Intel-CSU Transparent Computing and Big Data Summit, 2014.

5. Keynote Speaker: Collaborative Mobile Charging and Coverage in Wireless Sensor Networks, THU-SJTU Joint Workshop on Mobile Computing and Networking, 2014.

6. Keynote Speaker: Collaborative Mobile Charging and Coverage in WSNs, MiSeNet (in conjunction with ACM MobiCom), 2013.

7. Plenary Speaker: Resilient Priority-Based Data Transmission Using Network Coding, Resilience Week, 2013.

8. Keynote Speaker: Algorithmic Crowdsourcing, The 5th International Workshop on Hop Topics in Peer-to-peer Computing and Online Social Networking (HotPOST), July 2013.

9. Keynote Speaker: Utility-based Routing and Its Application in Dynamic Wireless Networks, IEEE TrustCom, June 2012.

10. Keynote Speaker: Utility-based Routing and its Application in Dynamic Wireless Network, IEEE ChinaCom, August 2012.

11. Keynote Speaker: Some Reflections on C&I Education, CCF CNCC, November 2011.

12. Keynote Speaker: Some Challenges in Routing in Dynamic Networks, 2011 IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS), October 2011.

13. Keynote Speaker: Trust Mechanisms and Their Applications in Dynamic and Mobile Computer Systems, 2009 IEEE/IFIP International Symposium on Trusted Computing and Communications, August 2009.

14. Keynote Speaker: A Utility-based Routing Scheme and Its Applications. 2009 IEEE International Symposium on Parallel and Distributed Processing with Applications(ISPA), August 2009.

15. Keynote Speaker: On Self-Organization in MANETs, The 2nd ACM International Workshop on Foundations of Wireless Ad Hoc and Sensor Networking and Computing (FAWANC), May 2009.

16. Keynote Speaker: A Utility-based Routing Scheme in Ad Hoc Networks, 2008 International Conference on Wireless Algorithms, Systems and Applications (WASA), October 2008.

17. Keynote Speaker: Mobility in Ad Hoc Wireless Networks: Friend or Foe. The 8th IEEE International Symposium on Network Computing and Applications (NCA), July 2008.

18. Keynote Speaker: Mobility in Ad Hoc Wireless Networks: Friend or Foe. 2008 International Conference on Networking, Architecture, and Storage (NAS), July 2008.

19. Keynote Speaker: A Utility-based Routing Scheme in MANETs, The 4th IEEE International Conference on Mobile Ad-Hoc and Sensor Systems (MASS), October 2007.

20. Keynote Speaker: Mobility in Wireless Networks: Blessing or Curse. The 4th Annual International Conference on Mobile and Ubiquitous Systems: Computing, Networking and Services (MOBIQUI-TOUS 2007), August 2007.

21. Keynote Speaker: Controlled Store-Carry-Forward Routing in Mobile Ad Hoc Networks. The First International Workshop on Specialized Ad Hoc Networks and Systems (SAHNS 2007), in conjunction with IEEE ICDCS’07, June 2007.

22. Keynote Speaker: The 2nd International Conference on Mobile Ad-Hoc and Sensor Networks (MSN), 2006.

23. Keynote Speaker: The 6th International Workshop on Parallel and Distributed Scientific and Engineering Computing (PDSEC), 2005.

24. Keynote Speaker: The 3rd International Workshop on Mobile Distributed Computing (MDC), 2005.

25. Keynote Speaker: The 6th International Workshop on Advances in Parallel and Distributed Computational Models, 2004.

26. Keynote Speaker: International Conference on Parallel and Distributed Computing, Applications and Technologies (PDCAT), 2004.

27. Keynote Speaker: International Symposium on Parallel Processing and Applications, 2003.

4.3 Panels (9)

1. Panelist: Networking Cyber-physical Applications in a Data-centric World, IEEE ICCCN, 2015.

2. Panelist: Mobile Cloud Computing Challenges and Potential-A Communications and Networking Perspective, IEEE INFOCOM, 2014.

3. Panelist: NSF/TCPP Curriculum on Parallel/Distributed Computing for Undergraduates, IEEE Edu-Par, 2012.

4. Panelist: the Network ViewPoint of Internet of Things (IOT), IOT and Sensor Networks Forum, CNCC, 2011.

5. Panelist: Mobile and Wireless Networking, IEEE ICCCN, 2011.

6. Panel Organizer: Network Science: Hype or Reality? ACM MobiCom, 2009.

7. Panelist: Mobile Computing and Systems: Where is the Tofu?, ACM MobiCom/MobiHoc, 2007.

8. Panel Organizer: Network Systems in Challenging Environments: Hype or Reality, IEEE INFOCOM, 2008.

9. Panelist: Pervasive and Mobile Computing: Continued Challenges and Issues, IEEE ICCCN, 2008.

4.4 Distinguished Lectures and Invited Seminars (152)

1. “Optimizing Roadside Advertisement Dissemination in Vehicular CPS,” Shanghai Jiaotong University, June 2015.

2. “Collaborative Mobile Charging and Coverage,” University of Science and Technology Beijing, June 2015.

3. “Algorithmic Corwdsourcing: Current State and Future Perspective,” Tsinghua University, June 2015.

4. “Optimizing Roadside Advertisement Dissemination in Vehicular CPS,” Central South University, June 2015.

5. “Algorithmic Corwdsourcing: Current State and Future Perspective,” Huazhong University of Science and Technology, June 2015.

6. “Optimizing Roadside Advertisement Dissemination in Vehicular CPS,” Sat Yet Sen University, June 2015.

7. “Collaborative Mobile Charging and Coverage,” Zhejiang University of Technology, June 2015.

8. “Collaborative Mobile Charging and Coverage,” Hangzhou Electronic Technology University, June 2015.

9. “Algorithmic Crowdsourcing: Current Status and Future Perspective,” Texas State University, December 2014.

10. “Collaborative Mobile Charging and Coverage in Wireless Sensor Networks,” Syracuse University, October 2014.

11. “Algorithmic Crowdsourcing and Funding Opportunities,” Purdue University Calumet, September 2014.

12. “Algorithmic Crowdsourcing and Applications in Social Networking,” Illinois Institute of Technology, September 2014.

13. “Algorithmic Crowdsourcing and Applications in Social Networking,” Colorado State University IS-TeC, March 2014. (Distinguished Lecture)

14. “Collaborative Mobile Charging and Coverage in Wireless Sensor Networks,” Colorado State University, March 2014.

15. “Algorithmic Crowdsourcing and Applications in Big Data,” Shenzhen Institutes of Advanced Technology, May 2014.

16. “Collaborative Mobile Charging and Coverage in Wireless Sensor Networks,” Pontificia Universidad Catolica de Chile, June 2014.

17. “Algorithmic Crowdsourcing and Applications in Big Data,” Universidad Finis Terrae, June 2014.

18. “Algorithmic Crowdsourcing and Applications in Big Data,” Tsinghua University, July 2014.

19. “Collaborative Mobile Charging and Coverage in Wireless Sensor Networks,” Institute of Computer Technology, Chinese Academy of Science, July 2014. (Distinguished Lecture)

20. “Algorithmic Crowdsourcing and Applications in Big Data,” Suzhou University, July 2014.

21. “Collaborative Mobile Charging and Coverage in Wireless Sensor Networks,” University of Electronic Science and Technology of China, July 2014.

22. “Collaborative Mobile Charging and Coverage in Wireless Sensor Networks,” Jilin University, July 2014.

23. “Collaborative Mobile Charging and Coverage in Wireless Sensor Networks,” New Jersey Institute of Technology, November 2013.

24. “Algorithmic Crowdsourcing,” Hunan University, November 2013.

25. “Resilient Priority-Based Data Transmission Using Network Coding,” Resilience Week, August 2013.

26. “Social-Aware Routing in Delay Tolerant Networks,” Central South University, June 2013.

27. “Utility-Based Routing,” Dalian University of Technology, June 2013.

28. “Utility-based Routing and Its Applications,” Wayne State University, March 2013. (ACM Distinguished Lecture)

29. “Utility-based Routing and Its Applications in Dynamic Wireless Networks,” University of Michigan Dearborn, October 2012. (Distinguished Lecture)

30. “Utility-based Routing,” University of Belgrade, September 2012.

31. “Utility-based Routing,” Hunan University, August 2012.

32. “Utility-based Routing,” Hunan University of Humanities, Science and Technology, August 2012.

33. “Cooperative Private Searching in Clouds,” Tsinghua University, July 2012. (Intellectual Venture Chair Professor Group Distinguished Lecture)

34. “A Utility-based Routing in Dynamic Wireless Network,” Rutgers University, March 2012.

35. “Some Challenges in Routing in Dynamic Networks,” Ohio State University, Oct. 2011. (Distinguished Lecture)

36. “Security and Privacy in Multimedia Retrieval in Cloud,” The Polytechnic University of Hong Kong, June 2011.

37. “Challenges in Wireless and Mobile network Applications,” University of Electronic Science and Technology of China, June 2011.

38. “Challenges in Wireless and Mobile network Applications,” Tsinghua University, June 2011.

39. “Challenges in Wireless and Mobile network Applications,” UAE University, June 2011. (Distinguished Lecture)

40. “A Utility-based Routing in Wireless Networks,” City of University of Hong Kong, June 2011.

41. “Challenges in Wireless and Mobile network Applications,” The Polytechnic University of Hong Kong, May 2011.

42. “Mobility in Wireless Networks: Friends or Foes,” Pennsylvania State University, April 2011.

43. “Mobility in Wireless Networks: Friends or Foes,” University of Pennsylvania, March 2011.

44. “Mobility in Wireless Networks: Friends or Foes,” Indian Institute of Technology, Delhi, January 2011.

45. “Mobility in Wireless Networks: Friends or Foes,” Indian Institute of Technology, Bombay, January 2011.

46. “Mobility in Wireless Networks: Friends or Foes,” College of William and Mary, October 2010. (Distinguished Lecture)

47. “Taming Complex Dynamic Network Models: Slicing and Visualization,” Shanghai Jiaotong University, China, July 2010.

48. “The Wireless Networks in Challenged Environments: Challenges and Opportunities,” Central South University, China, July 2010.

49. “Taming Complex Dynamic Network Models: Slicing and Visualization,” The Polytechnic University of Hong Kong, July 2010.

50. “The Wireless Networks in Challenged Environments: Challenges and Opportunities,” Zhongshan University, China, July 2010.

51. “Computer Technology in a Changing World of Globalization,” University of Ricardo Palma, Lima, Peru, May 2010.

52. “Mobility in Wireless Networks: Friends or Foes,” Lehigh University, April 2010.

53. “Mobility in Wireless Networks: Friends or Foes,” Drexel University, March 2010.

54. “Mobility in Wireless Networks: Friends or Foes,” University of Memphis, March 2010.

55. “Mobility in Wireless Networks: Friends or Foes,” Stevens Institute of Technology, Dec. 2009.

56. “Mobility in Wireless Networks: Friends or Foes,” University of Illinois at Urbana-Champaign, Oct. 2009.

57. “Trust Mechanisms and Applications in Dynamic Network Systems,” Indiana University-Purdue University Indianapolis, Oct. 2009.

58. “Mobility in Wireless Networks: Friends or Foes,” Institute of Software at Chinese Academic of Sciences, Sept. 2009.

59. “Mobility in Wireless Networks: Friends or Foes,” Tsinghua University, Sept. 2009.

60. “Mobility in Wireless Networks: Friends or Foes,” Shanghai Jiaotong University, May 2009.

61. “Research Directions in Mobile Ad Hoc Networks,” Shanghai Jiaotong Univresity, Shanghai, China August 2006.

62. “Research Directions in Mobile Ad Hoc Networks,” Nanjing University, Nanjing, China, July 2006.

63. “Mobility in Mobile Ad Hoc Networks: Friend or Foe,” Zhejiang University, Zhejiang, China, July 2006.

64. “Research Directions in Mobile Ad Hoc Networks,” National University of Defense Technology, Changda, China, July 2006.

65. “Research Directions in Mobile Ad Hoc Networks,” Inner Mongolia University, Inner Mongolia, Huhot, China, June 2006.

66. “Mobility in Mobile Ad Hoc Networks: Friend or Foe,” Toyko Denki University, Toyko, Japan, June 2006.

67. “Mobility in Mobile Ad Hoc Networks: Friend or Foe,” Wayne State University, April 2006.

68. “Mobility in Mobile Ad Hoc Networks: Friend or Foe,” University of Alabama, March 2006.

69. “Self-Stabilization in Mobile Ad Hoc Networks,” Georgia Institute of Technology, November 2005.

70. “Mobility in Mobile Ad Hoc Networks: Friend or Foe,” Ohio State University, October 2005. (Distinguished Lecture)

71. “Localized Algorithms and Their Applications in Ad Hoc Wireless Networks,” Montclair State University, September 2005.

72. “How to Write a Successful Proposal: A Personal Reflection,” Montclair State University, September 2005.

73. “Research Directions in Mobile Ad Hoc Networks,” National Central University, Taiwan, July 2005.

74. “Scan-Based Load Distribution in Sensor Networks,” National Central University, Taiwan, July 2005.

75. “Dominating-Set-Based Routing in Ad Hoc Networks,” National Central University, Taiwan, July 2005.

76. “Dominating-Set-Based Routing in Ad Hoc Networks,” Hong Kong Baptist University, Hong Kong, July 2005.

77. “Mobility in Mobile Ad Hoc Networks: Friend or Foe,” City University of Kong Kong, Hong Kong, July 2005.

78. “Scan-Based Load Distribution in Sensor Networks,” Hong Kong Polytechnic University, Hong Kong, July 2005.

79. “Dominating-Set-Based Routing in Ad Hoc Networks,” University of Burnos Aires, Burnos Aires, Argentina, June 2005.

80. “Localized Algorithms and Their Applications in Ad Hoc Wireless Networks,” College of Computing, Wuhan University, Wuhan, China, May 2005.

81. “Ad Hoc Networks,” College of Computing, Huazhong University of Science and Technology, Wuhan, China, May 2005.

82. “Localized Algorithms and Their Applications in Ad Hoc Wireless Networks,” College of Computing, Shanghai University, Shanghai, China, May 2005.

83. “Dominating-Set-Based Routing in Ad Hoc Wireless Networks,” Department of Computer Science, American University of Beirut, Beirut, Lebanon, January 2005.

84. “Dominating-Set-Based Routing in Ad Hoc Wireless Networks,” Zhongshang University, Guangzhou, China, December 2004.

85. “Local-Information-Based Routing and Broadcasting in Ad Hoc Networks,” Department of Computer Engineering, George Washington University, November 2004.

86. “Efficient and Localized Broadcasting in MANETs,” NSF PI Meeting, Manchester, NH, November 2004.

87. “Localized Algorithms and Their Applications in Sensor Networks,” Department of Computer Science, Shanghai Jiaotong University, Shanghai, China, July 2004.

88. “Localized Algorithms and Their Applications in Ad Hoc (and Sensor) Networks,” International Conference on High Performance Computational Applications, Shanghai, China, July 2004.

89. “Localized Algorithms and Their Applications in Sensor Networks,” Tyco, Boca Raton, FL, May 2004.

90. “Secured Routing Protocols in Ad Hoc Wireless Networks,” Dept. of Computer Science, West Chester University, April 2004.

91. “Dominating-Set-Based Searching in Peer-to-Peer Networks,” University of Aziu, Aziu, Japan, March 2004.

92. “Dominating-Set-Based Routing in Ad Hoc Wireless Networks,” Dept. of Computer Science, Nanjing University, Nanjing, China, March 2004.

93. “Localized Algorithms and Their Applications in Ad Hoc Wireless Networks,” Dept. of Computer Science, Hong Kong University of Science and Technology, Hong Kong, March 2004.

94. “Distributed System Design: An Overview,” National University of Defense Technology, Changsa, China, December 2003.

95. “Dominating-Set-Based Routing in Ad Hoc Wireless Networks,” College of Engineering, Central Chinese University, Changsa, China, December 2003.

96. “Dominating-Set-Based Routing in Ad Hoc Wireless Networks,” Dept. of Computer Science, Michigan State University, November 2003.

97. “Dominating-Set-Based Routing in Ad Hoc Wireless Networks,” Dept. of Electrical and Computer Engineering, Old Dominion University, October 2003.

98. “Localized Algorithms and Their Applications in Ad Hoc Wireless Networks,” Chinese NSF Workshop on Advance in Computer Science, Fujian, China, December 2003.

99. “Dominating-Set-Based Routing in Ad Hoc Wireless Networks,” Dept. of Computer Science, University of Tsukuba, Tsukuba, Japan, July 2003.

100. “Dominating-Set-Based Routing in Ad Hoc Wireless Networks,” Dept. of Electrical and Computer Engineering, McGill University, Montreal, Canada, June 2003.

101. “Dominating-Set-Based Routing in Ad Hoc Wireless Networks,” Dept. of Computer Engineering, University of Massachusetts at Amherst, November 2002.

102. “Dominating-Set-Based Routing in Ad Hoc Wireless Networks,” Dept. of Computer Science, University of Nebraska at Lincoln, October 2002.

103. “Dominating-Set-Based Routing in Ad Hoc Wireless Networks,” Dept. of Computer Science, Northeastern University, Qinghuangdao, China, August 2002.

104. “Distributed System Design,” Dept. of Computer Science, Nanjing University, Nanjing, China, July 2002.

105. “Dominating-Set-Based Routing in Ad Hoc Wireless Networks,” Dept. of Computer Science, University of Connecticut, December 2001.

106. “Distributed System Design: An Overview,” College of Information Science, Shanghai Jiaotong University, Shanghai, China, December 2001.

107. “Fault-Tolerant Routing in Hypercubes,” Dept. of Mathematics, Polytechnic University of Milan, Milan, Italy, November 2001.

108. “Dominating-Set-Based Routing in Ad Hoc Wireless Networks,” Dept. of Computer Science, Polytechnic University of Milan, Milan, Italy, November, 2001.

109. “Dominating-Set-Based Routing in Ad Hoc Wireless Networks,” Dept. of Computer Science, Zhejiang University, Hangzhou, China, June 2001.

110. “Distributed System Design: An Overview,” Dept. of Computer Science, Tamkang University, Taiwan, May 2001.

111. “Dominating-Set-Based Routing in Ad Hoc Wireless Networks,” Dept. of Computer Science, National Jiaotong University, Taiwan, May 2001.

112. “Dominating-Set-Based Routing in Ad Hoc Wireless Networks,” Academy of Science, Taiwan, May 2001.

113. “On Sorting An Intransitive Total Ordered Set Using Semi-Heap,” Dept. of Computer Science, Kent State University, January 2001.

114. “Parallel Processing: Past, Present, and Future,” Dept. of Computing, The Hong Kong Polytechnic University, Hong Kong, December 2000.

115. “On Sorting An Intransitive Total Ordered Set Using Semi-Heap,” Dept. of Computing, The Hong Kong Polytechnic University, Hong Kong, December 2000.

116. “Routing Protocols for Ad Hoc Wireless Networks,” Dept. of Computer Science, University of Minnesota (sponsored by Cray Inc.), October 2000.

117. “Routing Protocols for Ad Hoc Wireless Networks,” Dept. of Electrical & Computer Engineering and Computer Science, University of Cincinnati (IEEE Distinguished Lecture), October 2000.

118. “Routing Protocols for Ad Hoc Wireless Networks,” IEEE Ottawa Chapter (IEEE Distinguished Lecture), August 2000.

119. “Routing Protocols for Ad Hoc Wireless Networks,” Dept. of Computer Science, City University of Hong Kong, Hong Kong, June 2000.

120. “Routing Protocols for Ad Hoc Wireless Networks,” Dept. of Computer Science, Shanghai Jiaotong University, Shanghai, China, May 2000.

121. “Routing Protocols for Ad Hoc Wireless Networks,” Dept. of Computer Science and Technology, Tsinghua University, Beijing, China, May 2000.

122. “Routing Protocols for Ad Hoc Wireless Networks,” Dept. of Computer Science, Shanghai University, Shanghai, China, May 2000.

123. “Routing Protocols for Ad Hoc Wireless Networks,” IEEE Central Texas Chapter (IEEE Distinguished Lecture), January 2000.

124. “Routing Protocols for Ad Hoc Wireless Networks,” Dept. of Mathematics and Computer Science, Southwest Texas State University, January 2000.

125. “Routing Protocols for Ad Hoc Wireless Networks,” Motorola at Plantation, Florida, November 1999.

126. “On Sorting An Intransitive Total Ordered Set Using Semi-Heap,” Dept. of Computer Science, Nanjing University, Nanjing, China, June 1999.

127. “Fault-Tolerant Communication in Direct Networks,” Dept. of Computer Science, University of Toledo, May 1999.

128. “Fault-Tolerant Communication in Direct Networks,” IEEE Philadelphia Chapter, January 1999. (IEEE Distinguished Lecture)

129. “Parallel Processing: Past, Present, and Future,” Dept. of Computer Science, Univ. of New South Wale, Sydney, Australia, December 1998.

130. “Fault-Tolerant Communication in Point-to-Point Networks,” Dept. of Computer Science, Old Dominion University, November 1998.

131. “Fault-Tolerant Adaptive and Minimal Routing in Mesh-Connected Multicomputers,” Dept. of Computer Science, National Central University, Taiwan, May 1998.

132. “Reliable Communication in Cube-Based Multicomputers,” Dept. of Computer Science, National Taiwan University, Taiwan, May 1998.

133. “Distributed System Design: An Overview,” Dept. of Computer Science, National Tsinghua University, Taiwan, May 1998.

134. “Parallel Processing: Past, Present, and Future,” Dept. of Engineering Science, National Cheng-Kung University, Taiwan, May 1998.

135. “Reliable Communication in Cube-Based Multicomputers,” Dept. of Computer Science, Tokyo Denki University, Tokyo, Japan, April 1998.

136. “Reliable Communication in Cube-Based Multicomputers,” Dept. of Computer Science, The University of Vermont, September 1997.

137. “Minimal Routing in 2-D Meshes Using Extended Safety Levels,” Dept. of Computer Science, Nanjing University, Nanjing, China, July 1997.

138. “Reliable Communication in Cube-Based Multicomputers Using Safety Vectors,” Dept. of Computer Science, Nanjing University, Nanjing, China, July 1997.

139. “Reliable Communication in Cube-Based Multicomputers,” Dept. of Computer Science, Xi’an Jiaotong University, Xi’an, China, July 1997.

140. “Parallel Processing: Past, Present, and Future,” Institute of Computing Science, Chinese Academic of Science, Beijing, China, June 1997.

141. “Reliable Communication in Cube-Based Multicomputers,” Dept. of Computer Science and Technology, Tsinghua University, Beijing, China, June 1997.

142. “Parallel Processing: Past, Present, and Future,” Dept. of Computer Science, Nanjing University, Nanjing, China, June 1996.

143. “Distributed System Design,” American University of Bulgaria, Blagoevgrad, Bulgaria, April 1996.

144. “Unicasting in Faulty Hypercube Using Safety Levels,” Dept. of Computer Science, University of Newcastle Upon Tyne, Newcastle, England, March 1996.

145. “Parallel Processing: Past, Present, and Future,” Motorola at Boynton Beach, FL, September 1995.

146. “Reliable Communication Using Safety Levels,” Dept. of Computer Science, Nanjing University, Nanjing, China, July 1995.

147. “Parallel Processing: Past, Present, and Future,” The 4th International Conference for Young Computer Scientists, Beijing, China, July 1995.

148. “Parallel Processing: Past, Present, and Future,” Dept. of Computer Science, Shanghai Railway University, Shanghai, China, July 1995.

149. “Unicasting in Faulty Hypercubes Using Safety Levels,” Dept. of Computer Science, Louisiana State University, February 1995.

150. “Reliable Broadcasting in Injured Hypercubes,” Institute of Computing Science, Polish Academy of Sciences, Warsaw, Poland, June 1994.

151. “Reliable Broadcasting in Injured Hypercubes,” Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China, July 1993.

152. “Reliable Broadcasting in Injured Hypercubes,” Dept. of Computer Science, Shanghai University of Science and Technology, Shanghai, China, July 1993.

153. “Optimal Binomial-Tree Based Broadcasting in Injured Hypercubes,” Dept. of Computer and Information Science, Polytechnic University of Madrid, Madrid, Spain, January 1993.

154. “Broadcasting in Faulty Hypercubes,” Dept. of Computer Science, University of Hong Kong, Hong Kong, July 1992.

155. “The Design of Fault-Tolerant Concurrent Software Using Petri Nets,” College of Computing, Georgia Institute of Technology, May 1991.

156. “The Design of Fault-Tolerant Concurrent Software Using Petri Nets,” Dept. of Computer Science, Old Dominion University, April 1991.

157. “Cube-Connected-Cube Networks,” Dept. of Computer Science, University of Rhode Island, October 1990.

158. “Hypercube Multicomputers,” IEEE Computer Society, Palm Beach Chapter, September 1990.

159. “The Design of Fault-Tolerant Concurrent Software Using Petri Nets,” Dept. of Computer Science, Utah State University, February 1990.

160. “The Design of Fault-Tolerant Concurrent Software Using Petri Nets,” Dept. of Computer Science, Xavier University, New Orleans, August 1989.

5 Research

5.1 Sponsored Research

• REU Site: Enhancing Undergraduate Experience in Mobile Cloud Computing (PI) (CNS 1460971) 05/2015 - 04/18 Supported by National Science Foundation (NSF), Arlington, VI. ($342,525)

• US Ignite: Mobility-Enhanced Public Safety Surveillance System using 3D Cameras and High Speed Broadband Networks (PI) (CNS 1449860) 09/14 - 08/16 Supported by National Science Foundation (NSF), Arlington, VI. ($199,995)

• I/UCRC Phase II: Center on Intelligent Storage (Co-PI) (IIP 1439672) 09/14-08/19 Supported by National Science Foundation (NSF), Arlington, VI. ($65,000)

• A Test-bed of Secure Mobile Cloud Computing for Military Applications (Co-PI) 09/14 - 08/15 Supported by Army Research Office. ($86,211)

• NSF CAREER Workshop (PI) (CCF 1301774) 12/12-11/13 Supported by National Science Foundation (NSF), Arlington, VI. ($49,999)

• Collaborative Research: CCSS: Cyber-Enabled Smart Systems for Seamless Secure Monitoring and Communications (PI) (ECCS 1231461) 09/12-08/15 Supported by National Science Foundation (NSF), Arlington, VI. ($149,868)

• REU Site: Enhancing Undergraduate Experience in Next Generation Networking Technologies (Co-PI) (CNS 1156574) 06/12 - 05/16 Supported by National Science Foundation (NSF), Arlington, VI. ($319,932)

• Urban Apps and Maps Ecosystem Studio (Co-PI) 10/11 - 09/15 Supported by Economic Development Adminstration (EDA), Department of Commerce. ($500,000)

• Eager: A Meso-Scale GENI WiMax Project (PI) (CNS 1138963) 09/11 - 08/12 Supported by National Science Foundation (NSF), Arlington, VI. ($199,716)

• Formula Grant: Body Sensor Networks and Their Applications in Maternal Fetal Monitoring (PI) 09/11 - 08/13 Supported by State of Pennsylvania. ($300,000)

• NeTS: Medium: Collaborative Research: Mobilie Content Sharing Networks: Theory to Implementation. (PI) (CNS 1065444) 07/11 - 06/15 Supported by National Science Foundation (NSF), Arlington, VI. ($457,000)

• CCSS: An Architecture for Joint Integration of Inter and Intrasession Network Coding in Lossy Wireless Multihop Networks. (Co-PI) (ECCS 1128209) 08/11 - 07/14 Supported by National Science Foundation (NSF), Arlington, VI. ($329,163)

• EAGER: Mobile Multicore Computing. (PI) (CCF 1028167) 09/10 - 08/12 Supported by National Science Foundation (NSF), Arlington, VI. ($144,409)

• MRI-R2: Acquisition: A Hybrid High-Performance GPU/CPU System. (PI) (CNS 0958854) 05/10 - 04/13 Supported by National Science Foundation (NSF), Arlington, VI. ($839,221)

• EAGER: A New Algorithmic and Graph Model for Networking in Challenged Environments (PI) (CNS 0948184) 09/09 - 09/11 Supported by National Science Foundation (NSF), Arlington, VI. ($199,391)

• Energy-Efficient Design in Wireless Networks Using Cooperative Communication (PI) (CCF 0830289, renewal of CCR 039741) 09/08 - 09/11 Supported by National Science Foundation (NSF), Arlington, VI. ($277,370)

• Mobility-Assisted Routing in Mobile Networks. (PI) 09/06-09/09 Supported by National Science Foundation (NSF), Arlington, VA ($284,943)

• IEEE MASS 2006 Travel Grant. (PI) 09/06-09/07 National Science Foundation (NSF), Arlington, VA ($15,000)

• Acquisition of a NUMA-Based Supercluster for High Performance Computing. (PI) 9/05 - 9/08 Supported by National Science Foundation (NSF), Arlington, VA (CNS 0521410) ($447,191)

• IEEE MASS 2006 Travel Grants. (PI) 07/06-08/07 Supported by a grant from Army Research Office (ARO), NC. ($8,000)

• Service Grant. (PI) 06/06-09/06 Supported by SIVA Corp., Boca Raton, FL ($4,445)

• Design of High-Performance Publish/Subscribe Systems for Mobile Computing Environments. (Co-PI with Prof. J. Cao) 2006-2007 Supported by Hong Kong RGC grant, Hong Kong (805,200 HK Yuan = $102,000)

• Outstanding Overseas Young Researcher Award. (PI) 08/06 - present Supported by National Nature Science Foundation of China (NSFC), China (400,000 Yuan = $50,000)

• Research Experiences for Undergraduates (REU) 1. (PI) 05/06-05/07 Supported by National Science Foundation (NSF), Arlington, VA ($11,875)

• Research Experiences for Undergraduates (REU) 2. (PI) 05/06-05/07 Supported by National Science Foundation (NSF), Arlington, VA ($15,000)

• Research Experiences for Undergraduates (REU) 3. (PI) 05/06-05/07 Supported by National Science Foundation (NSF), Arlington, VA ($15,000)

• Efficient and Localized Broadcasting in Ad Hoc Wireless Networks. (PI) 09/04 - 09/07 Supported by National Science Foundation, Arlington, VA (CNS 0422762) ($225,000)

• Wireless and Sensor Network Laboratory. (Co-PI) 08/04 - 08/07 Supported by National Science Foundation, Arlington, VA (CNS 0434533) ($57,263)

• Secure Telecommunication Networks. (Co-PI) 09/04 - 09/06 Supported by Defense Information Systems Agency (DISA) ($300,000 out of $3,000,000)

• Parallel Programming. (PI) 05/05-08/05 Supported EDSS, Inc., Port St. Lucie, FL ($6,000)

• Energy Efficient Communication in Sensor Networks. (PI) 09/03 - 09/06 Supported by National Science Foundation (NSF), Arlington, VA (CCR 039741) ($375,000)

• International Workshop on Theoretical and Algorithmic Aspects of Sensor, Ad Hoc Wireless, and Peerto-Peer Networks. (PI) 10/03 - 10/04 Supported by National Science Foundation (NSF), Arlington, VA (CCR 0335238) ($22,200)

• Research Experiences for Undergraduates (REU) 4. (PI) 05/03 - 05/04 Supported by National Science Foundation (NSF), Arlington, VA ($11,650)

• Self-Organization Cluster-Tree Design in ns-2 Platform. (PI) 08/03 - 07/04 Supported by a grant Florida Communication Research Lab., Motorola, Inc. Plantation, FL ($60,000)

• Research Experiences for Undergraduates (REU) 5. (PI) 05/02 - 05/03 Supported by National Science Foundation (NSF), Arlington, VA ($5,500)

• A Low Cost Parallel and Distributed Systems Laboratory. (PI) 09/01 - 09/04 Supported by National Science Foundation (NSF), Arlington, VA (EIA 0130806) ($79,039)

• A Dominating-Set-Based Routing Scheme in Ad Hoc Wireless Networks. (PI) 06/00 - 06/03 Supported by National Science Foundation (NSF), Arlington, VA (ANI 0073736) ($181,957)

• A Theory of Fault-Tolerant Adaptive and Minimal Routing in Mesh-Connected Multicomputers. (PI) 04/99 - 04/03 Supported by National Science Foundation (NSF), Arlington, VA (CCR 9900646) ($134,109)

• Implementation of Radio Service Software. (PI) 01/97 - present Supported by a grant from Motorola, Inc. Plantation, FL ($1,898.255)

• Research Experiences for Undergraduates (REU) 6. (PI) 04/02- 05/03 Supported by National Science Foundation (NSF), Arlington, VA ($5,825)

• Software Path Coverage. (PI) 01/03 - 12/03 Supported by a grant from Motorola, Inc. Plantation, FL ($55,487)

• Cell Phone Software Testing.(PI) 01/02 - 12/02 Supported by a grant from Motorola, Inc. Coconut Creek, FL ($12,505)

• Web Design Using Java. (PI) 08/01-12/01 Supported by a grant from SIVA Corp., Boca Raton, FL ($15,890)

• Network Administration. (PI) 04/01-12/01 Supported by a grant from Malema Corp., Boca Raton, FL ($19,028)

• Testing of Application Software. (PI) 10/22/99-04/21/00 Supported by a grant from Motorola, Inc., Boynton, FL ($9,287)

• Implementation of Application Software. (PI) 08/12/99-12/31/99 Supported by a grant from Alpine Engineered, Boca Raton, FL ($11,582)

• Functional Parallelization of the Everglades Landscape Fire Model. (PI) 04/97-03/99 Supported by a grant from South Florida Water Management District, West Palm Beach, FL ($34,219)

• Domain-Specific Software Engineering on Pager Development. (PI) 01/97-01/98 Supported by a grant from Motorola, Inc., Boynton Beach, FL ($40,025)

• Load Balancing of the Everglades Landscape Model. (PI) 10/94-04/97 Supported by a grant from South Florida Water Management District, West Palm Beach, FL ($62,556)

• Computer System Evaluation. (PI) 01/90-05/95 Supported by a grant from IBM, Boca Raton, FL ($1,030,000)

• A Privilege-Based Security Model for Distributed Systems. (PI) 07/91-07/92 Supported by a grant from Florida Atlantic University Foundation, Inc. ($500)

• Fault-Tolerant Concurrent Software. (Co-PI) 01/90-12/90 Supported by a grant from Florida High Tech. and Industry Council ($52,744)

6 Professional Service

6.1 Membership Professional Societies

• Institute of Electrical and Electronics Engineering (IEEE): Computer Society (Fellow)

• Association for Computing Machinery (ACM) (Senior Member since 2006)

• American Association for the Advancement of Science (AAAS)

• Society for Computers and Their Applications (ISCA)

• Upsilon Pi Epsilon: Computing Science Honor Society

6.2 Editorial Board

• IEEE Transactions on Service Computing (1/2013 - present)

• IEEE Transactions on Computers (2009 - 2014)

• Science China (Information Sciences) (2014 - present)

• Journal of Parallel and Distributed Computing (2009 - present)

• Journal of Computer Science and Technology (A publication of the Chinese Computer Federation, 9/2002 - present)

• Frontier of Computer Science in China (01/2010 - present)

• International Journal of Security and Networks (IJSN) (01/2008)

• Ad Hoc & Sensor Wireless Networks: An International Journal (AHSWN) (10/2004)

• International Journal of Parallel, Emergent, and Distributed Systems (IJPEDS) (10/2004)

• Journal of Pervasive Computing and Communications (JPCC) (09/2004)

• Journal of Interconnection Networks (1/2007 - present)

• International Journal of High Performance Computing and Networking (IJHPCN) (09/2003)

• IEEE Transaction on Mobile Computing (11/2008 - 2010)

• ACTA Press/ IASTED International Journal of Computers and Applications (Co-Editor-in-Chief, 09/2003 - 01/2007)

• Computer Networks (2009)

• Elsevier Computer Communications (1/2008 - 12/2009)

• IEEE Transactions on Parallel and Distributed Systems (2001 - 2005)

• ISCA International Journal of Computers and Their Applications (Associate Editor in Networks and Protocols, 9/1999 - 12/2002).

• International Advisory Board, High Education Publishing Inc., China.

6.3 Conference Organization (major role)

1. General Chair: The 45th International Conference on Parallel Processing (ICPP), 2016.

2. General Chair: IEEE Conference on Communications and Network Security (CNS), 2016.

3. Symposium chair: Resilient Communication Systems, Resilience week, August 2015.

4. Steering committee member: Resilience Week, August, 2015.

5. Program Area Chair: IEEE Conference on Computer Communications (INFOCOM), 2015.

6. Steering Committee: ACM Annual International Workshop on Mission-Oriented Wireless Sensor Networking (NISENET).

7. General Chair: The 15th ACM International Symposium on Mobile Ad Hoc Networking and Computing (MobiHoc), 2014.

8. Program Area Chair: IEEE International Conference on Network Protocols (ICNP), 2014.

9. Program Area Chair: IEEE Conference on Computer Communications (INFOCOM), 2014.

10. General Chair: The 33rd IEEE International Conference on Distributed Computing Systems (ICDCS), 2013.

11. Program Chair: CCF China National Computer Congress (CNCC), 2013.

12. General Co-Chair: The 12th IEEE International Symposium on Signal Processing and Information Technology (ISSPIT), 2012.

13. Co-Chair: IEEE INFOCOM Achievement Award Committee, 2011.

14. Program Co-Chair: The 30th IEEE Conference on Computer Communications (INFOCOM), 2011.

15. Panel Co-Chair: The 29th IEEE Conference on Computer Communications (INFOCOM), 2010.

16. Program Co-Chair: International Conference on Distributed Computing and Networking (ICDCN) 2010.

17. General Co-Chair: The 6th IEEE/IFIP International Symposium on Trusted Computing and Communications (TrustCom), 2010.

18. Award Chair: The 5th International Conference on Networking Architecture and Storages (NAS), 2010.

19. General Co-Chair: The 3rd International Workshop on Sensor Networks (SN 2010), in conjunction with IEEE ICDCS, 2010.

20. General Co-Chair: International Conference on Distributed Computing in Sensor Systems (DCOSS), 2009.

21. General Chair: The 1st ACM International Workshop on Medical-grade Wireless Networks (WiMD 2009) (in conjunction with MobiHoc 2009), 2009.

22. Panel Co-Chair: ACM Annual International Conference on Mobile Computing and Networking (MobiCom), 2009.

23. Area TPC Chair: The 27th, 28th, 29th, 31st, 32th, 33th, and 34th IEEE Conference on Computer Communications (INFOCOM), 2008, 2009, 2010, 2012, 2013, 2014, and 2015.

24. Area TPC Chair: The 8th and 11th Annual IEEE Communications Society Conference on Sensor, Mesh and Ad Hoc Communications and Networks (SECON), 2011 and 2014.

25. General Chair: The 22nd IEEE International International Parallel & Distributed Processing Symposium (IPDPS), 2008.

26. Executive Vice Program Chair: The 28th IEEE International Conference on Distributed Computing Systems (ICDCS), 2008.

27. Panel Co-Chairs: The 5th Annual IEEE Communications Society Conference on Sensor, Mesh and Ad hoc Communications and Networks (SECON), 2008.

28. General Co-Chair: International Conference on Wireless Algorithms, Systems and Applications (WASA), 2007.

29. General Co-Chair: The 5th International Symposium on Parallel and Distributed Processing and Applications (ISPA’07), 2007.

30. General Vice Chair: The 21st IEEE International International Parallel & Distributed Processing Symposium (IPDPS), 2007.

31. Program Co-Chair: Symposium on Advances in Internet, ChinaCom, 2007.

32. General Chair: The 3rd IEEE International Conference on Mobile Ad-Hoc and Sensor Systems (MASS), 2006.

33. Program Co-Chair: International Workshop on Networking, Architecture, and Storage, 2006.

34. General Vice Chair: The 2nd IEEE International Conference on Distributed Computing Sensor Systems (DCOSS), 2006.

35. General Chair: International Workshop on Wireless and Sensor Networks Security (WSNS’05) (in conjunction with MASS), 2005.

36. Program Co-Chair: The 1st International Workshop on Security and Mobility in Peer-to-Peer Systems (MPPS) (in conjunction with ICDCS), 2005.

37. Program Co-Chair: The 1st International Workshop on Ad-Hoc, Sensor and P2P Networks (AHSP) (in conjunction with ISADS), 2005.

38. Program Co-Chair: International Conference on Mobile Ad-Hoc and Sensor Networks (MSN), 2005.

39. Program Vice Chair: International Conference on Computer Networks and Mobile Computing (ICC-NMC), 2005.

40. Program Chair: The 1st IEEE International Conference on Mobile Ad-Hoc and Sensor Systems (MASS), 2004.

41. General Co-Chair: The 2nd International Workshop on Mobile Distributed Computing (MDS), 2004.

42. Workshops Chair: The IEEE International Conference on Distributed Computing Systems (ICDCS), 2003.

43. General Co-Chair: The 7th International Conference on Computer Science and Informatics (ICS), 2003.

44. Program Co-Chair: The 4th and 5th Workshops on Advances in Parallel and Distributed Computational Models (in conjunction with IPDPS), 2002 and 2003.

45. Program Vice Chair: The 21st IEEE International Conference on Distributed Computing Systems (ICDCS), 2001.

46. Program Co-Chair: Workshop on Wireless Networks and Mobile Computing 2001 (in conjunction with ICPP 2001), 2001.

47. Program Vice Chair: The 29th International Conference on Parallel Processing (ICPP), 2000.

48. Program Co-Chair: The 12th ISCA International Conference on Parallel and Distributed Computing Systems (PDCS), 1999.

49. Local Arrangements Chair: The 5th IEEE International Symposium on High Performance Computer Architecture (HPCA), 1999.

50. Advisory Committee: The 18th and 21st International Conference on Advanced Information Networking and Applications (AINA), 2004 and 2007.

51. Advisory Committee: The 13th and 14th ISCA International Conference on Parallel and Distributed Computing Systems (PDCS), 2000 and 2001.

52. Advisory Committee: The 20th, 21st, and 22nd IASTED International Conference on Parallel and Distributed Computing Systems, 1999, 2000, and 2001.

53. Steering Committee Chair: The IEEE International Workshop on Wireless Security and Privacy (WiS-P).

54. Steering Committee Chair: The IEEE International Workshop on Wireless and Sensor Networks Security (WSNS).

55. Steering Committee: The 1st and 2nd Workshop on Mission-Oriented Wireless Sensor Networking (MiSeNet).

56. Steering Committee: The IEEE International Parallel & Distributed Processing Symposium (IPDPS)

57. Steering Committee: The 6th Workshop on Advances in Parallel and Distributed Computational Models (in conjunction with IPDPS).

58. Steering Committee: Workshop on Advances in Parallel and Distributed Computational Models (APD-CM).

59. Steering Committee: International Symposium on Parallel and Distributed Processing and Application (ISPA).

60. Steering Committee: The IEEE International Conference on Mobile Ad-Hoc and Sensor Systems (MASS).

61. Steering Committee: The IEEE Symposium on Reliable Distributed Systems (SRDS).

62. Steering Committee: The IEEE International Parallel & Distributed Processing Symposium (IPDPS).

63. Steering Committee: The ACM Annual International Workshop on Mission-Oriented Wireless Sensor Networking (NISENET).

64. Steering Committee: The Workshop on the Internet of Things Applications.

6.4 Conference Organization (program committee/session chair)

1. The 27th and 31st IEEE International Symposium on Reliable Distributed Systems (SRDS), 2008 and 2012 (TPC member).

2. The 1st, 8th and 9th IEEE IEEE International Conference on Distributed Computing in Sensor Systems (DCOSS), 2005, 2012, 2013 and 2015 (TPC member)

Workshop on Security and Reliability in Transparent Computing (SRTC), 2013, (Program Chair).

3. The 30th Military Communications Conference (MILCOM), 2011 (TPC member).

4. The 7th International Conference on Mobile Ad-hoc and Sensor Networks (MSN), 2011 (TPC member).

5. The 16th, 18th, 20th, 21st, 24th, 30th, 32nd, and 34th International Conference on Distributed Computing Systems (ICDCS), 1996, 1998, 2000, 2001, 2004, 2010, 2012, 2014, and 2015 (Program Committee).

6. The 15th, 16th, 17th, 18th, 19th, 20th and 21st IEEE International Conference on Network Protocols (ICNP), 2007, 2008, 2009, 2010, 2011, 2012 and 2013 (Program Committee).

7. The 5th, 6th and 10th Annual IEEE Communications Society Conference on Sensor, Mesh and Ad Hoc Communications and Networks (SECON), 2008, 2009 and 2013 (Program Committee).

8. The 3rd and 4th NSF/TCPP Workshop on Parallel and Distributed Computing Education (EduPar), 2013 and 2014 (Program Committee).

9. The 3rd International Workshop on Delay/Disruption-Tolerant Mobile Networks (IEEE DTMN), 2008 (Program Committee).

10. The 9th IEEE International Symposium on a World of Wireless, Mobile and Multimedia Networks (WoWMoM), 2008 (Program Committee).

11. The 8th International Conference on Algorithms and Architectures for Parallel Processing (ICA3PP), June 2008 (Program Committee).

12. The 3rd and 4th IEEE International Workshop on Sensor Networks and Systems for Pervasive Computing (PerSeNS), 2007 and 2008.

13. The 2nd ACM Workshop on Sensor Actor Networks (SANET 2008), in conjunction with ACM MobiCom, 2008 (Program Committee).

14. The 1st ACM International Workshop on Integrated Heterogeneous Sensor Networks (HeterSenet), in conjunction with ACM MobiHoc, 2008 (Program Committee).

15. The ACM International Symposium on Mobile Ad-Hoc Networking and Computing (MobiHoc), 2006 and 2007 (Program Committee).

16. International Conference on Parallel Processing (ICPP), 2001, 2002, 2004, 2006, 2007 and 2010 (Program Committee).

17. International Conference on Mobile Ad-hoc Sensor Systems (MASS), 2007, 2008, 2009, 2010, 2011, 2012, 2013, and 2015(Program Committee).

18. The 5th Annual IEEE International Conference on Pervasive Computing and Communications (Per-Com), 2007 (Program Committee).

19. The 2006 IEEE International Symposium on Modeling, Analysis, and Simulation of Computers and Telecommunication Systems (MASCOTS), 2006 (Program Committee).

20. The 1st ACM International Workshop on Performance Monitoring and Measurement of Heterogeneous Wireless and Wired Networks (PM2HW2H’2006), 2006 (Program Committee).

21. International Workshop on Networking, Architecture, and Storage, 2006 (Session Chair).

22. The 2nd International Conference on Mobile Ad-hoc Sensor Systems (MASS), 2005 (Session Chair).

23. The IEEE International Parallel and Distributed Processing Symposium (IPDPS), 2005 (Program Committee).

24. The 24th Annual Joint Conference of the IEEE Computer and Communications (INFOCOM), 2005 (Program Committee).

25. The 61st IEEE Vehicular Technology Conference (VTC), 2005 (Program Committee).

26. The 12th Annual International Conference on High Performance Computing (HiPC), 2005 (Program Committee).

27. International Conference on Parallel Processing (ICPP), 2000, 2001, 2002 and 2005 (Session Chair).

28. The IEEE 23th and 25th International Conference on Distributed Systems Conference (ICDCS), 2003 and 2005(Session Chair).

29. The 1st International Workshop on Sensor Networks and Systems for Pervasive Computing (PerSeN-S), 2005 (Program Committee).

30. The 1st IEEE Symposium on World of Wireless, Mobile and Multimedia Networks (WoWMoM), 2005 (Program Committee).

31. The 19th IEEE International Conference on Advanced Information Networking and Applications (AINA), 2005 (Program Committee).

32. International Workshop on Mobile Distributed Computing (MDC), 2005 (Program Committee).

33. The 1st International Workshop on System and Networking for Smart Objects (SaNSO), 2005 (Program Committee).

34. International Workshop of Ubiquitous Smart Worlds (USW), 2005 (Program Committee).

35. The 11th International Conference on Parallel and Distributed Systems (ICPADS), 2005 (Program Committee).

36. The 4th International Workshop Performance Modeling, Evaluation, and Optimization of Parallel and Distributed Systems, 2005 (Program Committee).

37. The 5th IEEE Workshop on Algorithms for Wireless, Mobile, Ad Hoc and Sensor Networks, 2005 (Program Committee).

38. The 2nd International Workshop on Wireless Ad Hoc Networking (WWAN), 2005 (Program Committee).

39. The IEEE International Workshop on Strategies for Energy Efficiency in Ad Hoc and Sensor Networks (IWSEEASN), 2005 (Program Committee).

40. International Workshop on Information Fusion and Dissemination in Wireless Sensor Networks, 2005 (Program Committee).

41. The Aizu Workshop on Peer-to-Peer Workshops, March 2004 (Guest Speaker).

42. The HKUST Workshop on Network Research, 2004 (Guest Speaker).

43. International Conference on Parallel and Distributed Processing and Applications (ISPA), 2004 (Session Chair).

44. International Conference on High Performance Computing and Applications (HPCA), 2004 (Session Chair).

45. The 1st International Conference on Embedded Software and Systems (ICESS), 2004 (Program Committee).

46. The 2004 International Conference on Embedded and Ubiquitous Computing (EUC), 2004 (Program Committee).

47. Euro-Par, 2004 (Program Committee).

48. The 23rd Annual Joint Conference of the IEEE Computer and Communications (INFOCOM), 2004 (Program Committee).

49. The IEEE Global Telecommunications Conference (GLOBECOM), 2004 (Program Committee).

50. International Conference on Parallel and Distributed Computing and Networks (PDCN), 2004 (Program Committee).

51. The 17th International Conference on Parallel and Distributed Computing Systems (PDCS), 2004 (Program Committee).

52. International Workshop on Wireless Ad hoc Networking (WWAN), 2004 (Program Committee).

53. The 2nd International Workshop on Mobile Distributed Computing (MDC), 2004 (Program Committee).

54. The 4th International Workshop on Algorithms for Wireless, Mobile, Ad Hoc and Sensor Networks (WMAN), 2004 (Program Committee).

55. International Workshop on Performance Modeling, Evaluation, and Optimization of Parallel and Distributed Systems (PME)(PDS), 2004 (Program Committee).

56. The Chinese NSF Workshop on Advance in Computer Science, 2003 (Guest Speaker).

57. The 2nd International Workshop on Grid and Cooperative Computing (GCC 2003), 2003 (Session Chair).

58. The 10th Annual International Symposium on High Performance Computing (HPCA), 2003 (Program Committee).

59. The 2003 High Performance and Large Scale Computing Conference, 2003 (Program Committee).

60. The 2003 International Workshop on Mobile Distributed Computing, 2003 (Program Committee).

61. The 3rd International Workshop on Internet Computing and E-Commerce, 2003 (Program Committee).

62. Workshop on Advances in Parallel and Distributed Computational Models, 2003 (Session Chair).

63. Workshop on Massive Parallel Processing, 2003 (Session Chair).

64. The 3rd International Conference on Communications and Computing, 2002 (Program Committee).

65. The 2002 International Conference on Parallel and Distributed Systems (ICPADS), 2002 (Program Committee).

66. The IASTED International Conference on Networks, Parallel and Distributed Processing and Applications, 2002 (Program Committee).

67. The 4th International Conference on Formal Engineering Methods, 2002 (Program Committee).

68. Workshop on Wireless Networks and Mobile Computing, 2001 (Session Chair).

69. The 14th ISCA International Conference on Parallel and Distributed Computing Systems (PDCS), 2001 (Session Chair).

70. The 14th ISCA International Conference on Parallel and Distributed Computing Systems (PDCS), 2001 (Program Committee).

71. The 2nd International Conference on Parallel and Distributed Computing, Applications and Technologies (PDCAT), 2001 (Program Committee).

72. The 2001 International Conference on Parallel and Distributed Systems (ICPADS), 2001 (Program Committee).

73. The 1st International Workshop on Parallel and Distributed Computing Issues in Wireless Networks and Mobile Computing (in conjunction with IPDPS 2001), 2001 (Program Committee).

74. The 3rd Workshop on High Performance Scientific and Engineering Computing with Application (HPSECA-01) (in conjunction with ICPP), 2001 (Program Committee).

75. Workshop on Advances in Parallel and Distributed Computational Models (in conjunction with IPDP-S), 2001 (Program Committee).

76. The 13th ISCA International Conference on Parallel and Distributed Computing Systems (PDCS), 2000 (Program Committee).

77. The 4th International Conference on Algorithms and Architectures for Parallel Processing, 2000 (Program Committee).

78. The 4th International Conference on Algorithms and Architectures for Parallel Processing, 2000 (Session Chairman).

79. The 1st International Conference on Parallel and Distributed Computing, Applications and Technologies (PDCAT), 2000 (Program Committee).

80. The 2000 International Conference on Parallel and Distributed Systems (ICPADS), 2000 (Program Committee).

81. The 31st Southeastern Conference on Combinatorics, Graph Theory and Computing, 2000 (Session Chairman).

82. The 4th International Conference on Algorithms and Architectures for Parallel Processing, 2000 (Session Chair).

83. The 5th International Symposium on High Performance Computer Architecture (HPCA), 1999 (Local Arrangements Chairman).

84. The 12th ISCA International Conference on Parallel and Distributed Computing Systems (PDCS), 1999 (Program Chairman).

85. International Conference on Parallel Processing (ICPP), 1999 (Program Member).

86. International Conference on Computer Applications in Industry and Engineering, 1999 (Program Member).

87. The 14th ISCA International Conference on Computers and Their Applications, 1999 (Program Member).

88. The 30 Southeastern Conference on Combinatorics, Graph Theory and Computing, 1999 (Session Chairman).

89. The 5th International Conference for Young Computer Scientists, 1999 (Program Member).

90. The 29th Southeastern Conference on Combinatorics, Graph Theory and Computing, 1998 (Session Chairman).

91. The 1998 International Conference on Parallel and Distributed Computing and Networks, 1998 (Session Chairman).

92. The 13th International Conference on Computers and Their Applications, 1998 (Program Committee).

93. The Annual IEEE Workshop on Fault-Tolerant Parallel and Distributed Systems, 1998 (Program Committee).

94. The 10th ISCA International Conference on Parallel and Distributed Computing Systems, 1997 (Session Chairman).

95. The 28th Southeastern Conference on Combinatorics, Graph Theory and Computing, 1997 (Session Chairman).

96. The 9th ISCA International Conference on Computer Application in Industry and Engineering, 1996 (Program Committee).

97. The 9th ISCA International Conference on Computer Application in Industry and Engineering, 1996 (Session Chairman).

98. The 8th International Conference on Parallel and Distributed Computing and Systems, 1996 (Program Committee).

99. The 27th Southeastern Conference on Combinatorics, Graph Theory and Computing, 1996 (Session Chairman).

100. The 1996 International Workshop on Computer-Aided Design, Test, and Evaluation for Dependability, 1996 (Program Committee).

101. The 16th International Conference on Distributed Computing Systems (ICDCS), 1996 (Session Chairman).

102. The 26th Southeastern Conference on Combinatorics, Graph Theory and Computing, 1995 (Session Chairman).

103. The 4th International Conference on Computer Communications and Networks, 1995 (Program Committee).

104. The 7th International Conference on Parallel and Distributed Computing Systems, 1994 (Session Chairman).

105. International Conference on Parallel Processing (ICPP), 1994 (Session Chairman).

106. The IEEE Southeastcon’94, 1994 (Session Organizer and Chairman)

107. The 25th Southeastern Conference on Combinatorics, Graph Theory and Computing, 1994 (Session Chairman).

108. The 3rd International Conference for Young Computer Scientist, 1993 (Program Committee).

109. The 3rd International Conference for Young Computer Scientist, 1993 (Session Chairman).

110. The 24th Southeastern Conference on Combinatorics, Graph Theory and Computing, 1993 (Session Chairman).

111. The 23rd Southeastern Conference on Combinatorics, Graph Theory and Computing, 1992 (Session Chairman).

112. The 23rd Pittsburgh Modeling and Simulation Conference, 1992 (Session Organizer and Chairman).

113. The IEEE Southeastcon ’91, 1991, (Session Organizer and Chairman).

114. The 3rd ISMM International Conference on Parallel and Distributed Computing and System, 1990 (Session Chairman).

115. 1989 ACM Southeast Regional Conference, 1989 (Program Committee).

6.5 NSF-related Activities

1. 2013 NSF CAREER Proposal Writing Workshop, March 15, 2013.

2. Organizer of NSF Wireless Networks PI meeting, Chicago, July 2007.

3. Organizer of NSF Workshops • Wireless Security, April 2008.

• Mobility in Wireless Networks: Challenges and Opportunities, July 2007.

• Bridging the Gap: Wireless Networking Technologies and Advances at the Physical Layer, August 2007.

• Participant of NSF Future Internet Architecture Summit, Nov. 2009.

• Participant of TCPP/NSF Curriculum Standards on Parallel & Distributed Computing, Jan.

2010.

4. Outreach/Site Visit (01/2007 - 12/2008)

• Univ. of California at Los Angeles, Northeastern Univ., MIT, Harvard, George Washington Univ., Rutgers Univ., Univ. of California at Riverside, Univ. of Southern California, Indiana Univ. - Purdue Univ. Indianapolis, George Mason Univ., Univ. of Delaware, Univ. of Texas at Arlington, Stanford, and Univ. of California at Berkeley.

6.6 Professional Referee Service

Proposals: National Science Foundation: China, Cyprus, Taiwan, and United States. Universities: City University of Hong Kong and Hong Kong Polytechnic University. Tenure/Promotion Reviews:

American University in Cario, Georgia State University, Kent State University, Nanjing University, Nanyuan Technological University of Singapore, Old Dominion University, Simon Fraser University, SUNY at Stony Brook, University of Alabama, and Wayne State University.

Journal Papers:

IEEE Computer IEEE Software IEEE Parallel and Distributed Technology IEEE Transactions on Software Engineering IEEE Transactions on Computers IEEE Transactions on Parallel and Distributed Systems IEEE Transactions on Mobile Computing ACM Transactions on Knowledge and Data Engineering ACM Transactions on Computer Systems ACM Computing Reviews The Computer Journal Distributed Computing Discrete Applied Mathematics Computer Networks and ISDN Systems International Journal of Mini and Micro Computers International Journal of Parallel and Distributed Systems and Networks International Journal of Computers and Their Applications International Journal of Modeling and Simulation Journal of Parallel and Distributed Computing Journal of Systems and Software Information Sciences: An International Journal Computer System’s Science and Engineering Concurrency – Practice and Experience Software – Practice and Experience Journal of Information Resources Management International Journal of Applied Intelligence Computers & Electrical Engineering Journal of Integrated Computer-Aided Engineering

Conferences:

ACM International Symposium on Mobile Ad-Hoc Networking and Computing (MobiHoc) International Conference of the IEEE Computer and Communications (INFOCOM) IEEE Global Telecommunications Conference (GLOBECOM) IEEE International Conference on Mobile Ad-hoc and Sensor Systems (MASS) IEEE International Conference on Communications (ICC) IEEE Semiannual Vehicular Technology Conference (VTC) IEEE International Conference on Distributed Computing in Sensor Systems (DCOSS) International Symposium on a World of Wireless, Mobile and Multimedia Networks (WoWMoM)

The 21st IEEE International Conference on Distributed Computing Systems 2001 International Conference on Parallel Processing IASTED International Conference on Parallel and Distributed Computing Systems 2001 Workshop on Advances in Parallel and Distributed Computational Models The 14th ISCA International Conference on Parallel and Distributed Systems Workshop on Wireless Networks and Mobile Computing (in conjunction with ICPP 2001) The 3rd Workshop on High Performance Scientific and Engineering Computing with Application (HPSECA-01) (in conjunction with ICPP 2001) The 4th International Conference on Algorithms and Architectures for Parallel Processing The 2nd International Conference on Parallel and Distributed Computing, Applications and Technologies The 1st International Workshop on Parallel and Distributed Computing Issues in Wireless Networks and Mobile Computing (in conjunction with IPDPS 2001) The 13th ISCA International Conference on Parallel and Distributed Computing Systems The First International Conference on Parallel and Distributed Computing, Applications and Technologies 2000 Proc. of International Parallel and Distributed Processing Symposium 2000 International Conference on Parallel Processing The 29th International Symposium on Fault-Tolerant Computing (FTCS-29) 1999 International Conference on Parallel Processing 1999 ISCA International Conference on Parallel and Distributed Computing Systems The 5th IEEE International Symposium on High-Performance Computer Architecture 1998 International Conference on Parallel and Distributed Computing and Networks 1998 International Conference on Distributed Computing Systems 1998 International Parallel Processing Symposium and International Symposium on Parallel and Distributed Processing 1997 ISCA International Conference on Parallel and Distributed Computing Systems 1997 IASTED International Conference on Parallel and Distributed Computing and Systems 1997 ISCA International Conference on Computer Applications in Industry and Engineering 1997 International Conference on Parallel Processing 1997 International Parallel and Processing Symposium The 7th International Symposium on Parallel and Distributed Processing 1996 IASTED International Conference on Parallel and Distributed Computing and Systems 1996 International Conference on Distributed Systems 1996 International Workshop on Computer-Aided Design, Test, and Evaluation for Dependability The 3rd International Conference for Young Computer Scientists 1996 International Conference on Parallel Processing The 6th International Symposium on Parallel and Distributed Processing 1995 International Conference on Parallel Processing 1994 International Conference on Parallel Processing 1993 International Conference on Parallel Processing The 25th International Symposium on Fault-Tolerant Computing (FTCS-25) The 23rd International Symposium on Fault-Tolerant Computing (FTCS-23) The 11th Symposium on Reliable Distributed System (SRDS-11)

The 12th Symposium on Reliable Distributed System (SRDS-12) 1992 International Parallel Processing Symposium 1991 International Conference on Parallel Processing The 3rd IEEE Symposium on Parallel and Distributed Systems 1990 PARBASE The 5th IFIP WG 11.3 Working Conference on Database Security Hawaii International Conference on System Science (HICSS-27) The 13th International Computer Software and Applications Conference (COMPSAC ’89) The 5th ISMM Conference on Parallel and Distributed Computing Systems.

Published Reviews of Books and Articles:

“Designing Tree-Based Barrier Synchronization on 2D Mesh Networks,” by J. -S. Yang and C. -T. King, Computing Reviews, Vol. 40, No. 2, February 1999, 115-116.

“Hierarchical Interconnection Structure for Field Programmable Gate Arrays,” by Y. -T. Lai and P. -T. Wang, Computing Reviews, Vol. 39, No. 2, February 1998, 113.

“A Hybrid Distributed Mutual Exclusion Algorithm” by Y-I Chang, Computing Reviews, Vol. 38, No. 10, October 1997, 506.

“A Bounded First-in, First-Enabled Solution to the l-exclusion Problem ” by A. Yehuda et al, Computing Reviews, Vol. 36, No. 10, October 1995, 531.

“Network Programming in C” by N. Barry, Computing Reviews, Vol. 32, No. 10, 1991, 490-491.

“Dependability of Resilient Computers” by T. Anderson, IEEE Computer, Vol. 23, No. 8, 1990, 126.

6.7 IEEE Computer Society

• Chairman: IEEE Technical Committee on Distributed Processing (12/2005-12/2012)

• Secretary and Treasurer: IEEE Technical Committee on Distributed Processing (08/1999-12/2005)

• IEEE Fellow Evaluation Committee (2008, 2009, 2010, 2014)

6.8 Local Community Service

• Chairman: IEEE Computer Society, Philadelphia Chapter (5/2013 - present)

• The 5th Annual Statewide Florida SECME Olympiad (1990)

• Florida PE preparation classes (1990)

6.9 University Service

Task Force, Temple University University Online Education (09/2012-present) Freeh Report (09/2012-present)

Dept. of Computer Science and Engineering, Florida Atlantic University Department Personnel Committee (09/2000-12/2006) Graduate and Research Committee (09/1993-09/1995, 09/1997-12/2006) Advisor of CE Graduate Students (09/1994-present) TA/GA Committee (09/1991-09/1997; Chair, 09/2002-12/2006) Director of CSE Graduate Programs (09/1997-09/2002) Executive Committee (09/1997-09/2002) IEEE Student Advisor (09/1995-09/2001) Research Committee (Chair, 09/1996-09/1997) ACM Student Advisor (09/1994-09/1995) Laboratory and Equipment Committee (09/1992-09/1993) Coordinator of Computer Engineer Ph.D Qualifying Examination (02/1991-09/1992)

College of Engineering, Florida Atlantic University College Personnel Committee (09/2001-12/2006, chair: 11/2005-12/2006) Graduate Studies (09/1993-09/1994) Math Liaison (09/1990-09/1991)

Florida Atlantic University University Tenure and Promotion Committee (11/2005-12/2006)

7 Instructional Experience

7.1 Graduate Level Courses (12 courses, including 9 new courses)

1. Fault-Tolerant System Design – Florida Atlantic University

2. Concurrent Processing – Florida Atlantic University

3. Analysis of Algorithms – Florida Atlantic University

4. Ad Hoc Networks – Florida Atlantic University (new course) and Temple University

5. Distributed Algorithms – Florida Atlantic University (new course) and Temple University

6. Internet Routing Protocols – Florida Atlantic University (new course)

7. Design and Analysis of Parallel Algorithms – Florida Atlantic University (new course)

8. Distributed System Design – Florida Atlantic University (new course)

9. Formal Specification Methods – Florida Atlantic University (new course)

10. Introduction to Parallel Computers – Florida Atlantic University (new course)

11. Real-Time Software – Florida Atlantic University (new course)

12. Distributed Computing – Temple University (new course)

7.2 Undergraduate Level Courses (14)

1. Cooperative Eduation in CS – Temple University

2. Cooperative Eduation in IS&T – Temple University

3. Design and Analysis of Algorithms – Florida Atlantic University

4. Data Structure and Algorithm Design Using C – Florida Atlantic University

5. Introduction to Computers – Florida Atlantic University

6. Data Structure and Algorithm Design Using C++ – Florida Atlantic University

7. Introduction to Logic Design – Florida Atlantic University

8. Structured Computer Architecture – Florida Atlantic University

9. Switching Systems – Florida Atlantic University

10. Senior Project – Florida Atlantic University

11. Senior Seminar – Florida Atlantic University

12. Engineering Design II – Florida Atlantic University

13. Programming Language Pascal – Shanghai University of Science and Technology

14. Programming Methodologies – Shanghai University of Science and Technology

7.3 Theses and Dissertations

7.3.1 Ph.D. Dissertations (21 completed and 7 in progress)

1. Y. Dai, “Spectrum Management and Cross-layer Protocol Design in Cognitive Readio Networks,” (Winter 2014). (Software Engineer, LinkedIn Inc.)

2. W. Jiang, “Researches on Personalized Trust Evaluation in Online Social Networks,” (Co-adviser, Summer 2014). (Assistant Professor, Hunan University, P. R. China)

3. Y. Wang, “Social Feature based Routing in Delay Tolerant Networks,” (Summer 2013). (Assistant Professor, Kettering University, Michigan, USA)

4. Q. Liu, “Security in Cloud Computing with Multi-User Data Sharing,” (Co-adviser, Winter 2012). (Assistant Professor, Hunan University, P. R. China)

5. Y. Zhao, “Efficient Content-Based Publish/SubscribeSystems for Dynamic and Large-Scale Networked Applications,” (Summer 2012). (Software Engineer, Google Inc.)

6. L. Cong, “Design and Implementation of Efficient Routing Protocols in Delay Tolerant Networks,” (Summer 2009). (Assistant Professor, Sun Yat-Sen University, P. R. China)

7. F. Li, “Uncertainty-aware Reputation System in Mobile Networks: Analysis and Applications,” (Summer 20009). (Assistant Professor at Indiana University-Purdue University Indianapolis)

8. A. Srinivasan, “Secured Broadcasting in Mobile Ad-Hoc Networks,” (Summer 2008).

(Assistant Professor at Bloomburg University of Pennsylvania, Pennsylvania)

9. M. Lu, “Incentive Mechanisms in Mobile Ad-Hoc Networks,” (Summer 2008)

(Assistant Professor at Central South China University, P. R. China)

10. S. Yang, “Connected Dominating Set (CDS) in Wireless Ad Hoc Networks: Variations with Applications,” (Fall 2007). (Assistant Professor at Purdue University at Calumet)

11. B. Wu, “Key Management in Secured Ad-Hoc Networks,” (Fall 2006).

(Assistant Professor at Fayetteville State University, North Carolina)

12. X. Li, “Efficient Searching in Unstructured Peer-to-Peer Networks,” (Fall 2006).

(Assistant Professor at University of North Carolina at Pembroke, North Carolina)

13. I. Jawhar, “Quality of Service Support in TDMA-Based Mobile Ad-Hoc Networks,” (Summer 2005). (Tenured Associate Professor at United Arab Emirates University, UAE)

14. F. Dai, “Localized Broadcasting in Ad-Hoc Wireless Networks,” (Summer 2005).

(Senior SOE, Microsoft Corp.)

15. W. Lou, “Broadcasting in Ad-Hoc Wireless Networks,” (Summer 2004).

(Assistant Professor at Hong Kong Polytechnic University, Hong Kong)

16. Z. Jiang, “Fault-Tolerant Routing in Mesh-Connected Networks with Dynamic Faults,” (Summer 2002). (Tenured Associate Professor at West Chester University, Pennsylvania)

17. X. Chen, “Fault-Tolerant Routing in 2-D and 3-D Meshes,” (Summer 1999).

(Tenured Associate Professor at Texas State University, Texas)

18. P. Goyal, “Cube-Connected-Cubes Multicomputers,” (Fall 1998).

19. K. Huang, “The Balanced Hypercube and Its Applications,” (Summer 1997).

20. Y. Wang, “A Uniform Hardware and Software Fault-Tolerant Model,” (Fall 1995).

21. L. Brown, “Fault-Tolerant Distributed Shared Memories,” (Fall 1993).

22. W. Chang, “Security and Privacy Issues in Social Information-assisted Application Design,” (in progress).

23. P. Ostovari, “Network Coding in Wireless Networks,” (in progress).

24. D. Li, “Multi-core Processing and Its Application in Cellular Phones,” (in progress).

25. H. Zheng, “Opportunistic Routing/Scheduling in Social Networks” (in progress).

26. N. Wang,“Delay Tolerant Network Applications” (in progress).

27. Y. Chen,“Cost-Efficient Resource Scheduling in Data Center Networks” (in progress).

28. J. Shang,“Spectrum Sharing and Access Fairness” (in progress).

7.3.2 MS Theses (35 completed) 1. Z. Wan, “Scheduling Policies for Cloud Computing,” (Summer 2015).

2. A. Blaisse, “Better Selection of Virtual Machines for A Mapreduce Environment,” (Summer 2015).

3. K. Li, “Resource Allocation Schemes in Cloud Systems,” (Summer 2014).

4. X. Li, “Channel Switching Control Policy for Wireless Mesh Networks,” (Summer 2012).

5. R. Thanawala, “Efficient Resource Discovery Technique in a Mobile Ad Hoc Network,” (Fall 2008).

6. V. Ponce, “Improving Cooperation in P2P Networks: A Social Network Approach,” (Fall 2007).

7. B. Qian, “Multi-Path Intelligent Virtual Mobile Nodes for Ad Hoc Networks,” (Fall 2007).

8. T. Anantvalee, “Reputation-Based System for Encouraging the Cooperation of Nodes in Mobile Ad Hoc Networks,” (Fall 2006).

9. J. Huang, “A Class-Based Search System in Unstructured Peer-to-Peer Networks,” (Summer 2006).

10. T. Xu, “Quorum-Based IP Address Autoconfiguration in Mobile Ad Hoc Networks,” (Summer 2006).

11. Y. Wang, “Label Routing Protocol: A New Cross-Layer Protocol for Multi-Hop Ad Hoc Wireless Networks,” (Spring 2006).

12. J. Zhang, “XYZ: A Scalable, Partially Centralized Lookup Service for Large-Scale Peer-to-Peer Systems,” (Summer 2005).

13. Q. Cheng, “Using Classification and Regression Tree to Detect Hematology Abnormalities,” (Fall 2004).

14. J. Wood, “Active Networking and Innovation: An Evaluation of Active Networking as a Driver of Accelerated Internet Innovation,” (Summer 2004).

15. C. Huang, “Enhanced Ticket-Based QoS Routing Protocols in Ad Hoc Networks,” (Fall 2003).

16. Q. Dai, “Energy Efficient Routing in Ad Hoc Wireless Networks,” (Summer 2003).

17. R. Khumanthem, “Mobile Agent and Its Application in Distributed Systems,” (Summer 2003).

18. B. Wu, “Power-Aware Broadcasting and Activity Scheduling in Ad Hoc Wireless Networks Using Connected Dominating Sets,” (Fall 2002).

19. K. Du, “Data Dissemination in Sensor Networks,” (Summer 2002).

20. M. Gao, “Dominating-Set-Based Routing in Ad Hoc Wireless Networks,” (Spring 2001).

21. J. Ibriq, “A Reduced Overhead Routing Protocol for Ad Hoc Wireless Networks,” (Fall 2000).

22. H. Li, “Routing in Mobile Ad-Hoc Wireless Networks,” (Fall 1999).

23. S. Cang, “Time-Step Optimal Broadcasting in Mesh Networks with Minimum Total Communication Distance,” (Spring 1999).

24. H. Fu, “Load Balance of a Fire Model in a Network of Workstations,” (Fall 1998).

25. H. Huang, “Parallel Algorithms for Domain Load Balance,” (Fall 1997).

26. A. C. Wynn, “The Effect of Compression on Performance in a Demand Paging Operating System,” (Summer 1997).

27. Y. Luo, “Embedding Binomial Trees in Faulty Hypercube Multiprocessors,” (Fall 1996).

28. J. Pan, “Static and Dynamic Load Balance Strategies for a Class of Applications,” (Summer 1996).

29. H. Qian, “Multitriangle: A New Interconnection Network,” (Summer 1995).

30. W. Huang, “Stewart Platform Accuracy Enhancement,” (Summer 1995).

31. H. More, “Load Balancing on Multiprocessor Systems,” (Summer 1993).

32. K. Yao, “Fault-Tolerant Multicasting in Hypercube Multicomputers,” (Spring 1993).

33. M. H. Fernandez, “Group Structures in Object-Oriented Database Authorization,” (Fall 1992).

34. M. Gangadhar, “Fault-Tolerant Scheduling for Multiprocessor Systems,” (Summer 1992).

35. M. Zhang, “Design and Modeling of Hybrid Software Fault-Tolerant Systems,” (Spring 1992).

8 Professional Recognition

• ACM Distinguished Lecturer, 7/2009 - 7/2013.

• CCF Distinguished Lecturer, 7/2013 - present.

• EMC-Endowed Visiting Chair Professor, Department of Computer Science, Tsinghua University, China, 2009-2011.

• IV-Endowed Visiting Chair Professor, National Lab for Information Science and Technology, Tsinghua University, China, 2011-present.

• IEEE Computer Society Fellow Evaluation Committee Member, 2008-2009 and 2009-2010.

• Outstanding Overseas Achievement Award, China Computer Federation (CCF), China, 2011.

• Outstanding Overseas Young Researcher Award, National Natural Science Foundation of China (NS-FC), China, 2006.

• Fulbright Senior Specialist, Council for International Exchange of Scholars, 12/2006-12/2011.

• University Researcher of the Year Award at the Rank of Professor: Florida Atlantic University, 2006-2007.

• International Advisory Board: School of Electronics, Information and Electrical Engineering (SEIEE), Shanghai Jiaotong University, Shanghai, China, 2002 - 2005.

• Guest Professor: Shanghai Jiaotong University, Shanghai, China

• AP Professor: Shenzhen Institutes of Advanced Technology, Chinese Academic of Sciences

• Guest Professor: National University of Defense Technology, Changsa, China.

• Guest Professor: Central South University, Changsa, China.

• Guest Professor: Shanghai University, Shanghai, China.

• Concurrent Professor: Nanjing University, Nanjing, China.

• IEEE Computer Society and ICDCS’03 Outstanding Service Award, May 2003.

• University Researcher of the Year Award at the Rank of Professor: Florida Atlantic University, 2001-2002.

• IEEE Distinguished Visitor: Computer Society, September 1998 - August 2001.

• Member of US delegation: Sino-American Technology and Engineering Conference, 1999.

• Outstanding Achievements Award: International Association of Science and Technology for Development (IASTED), 1998.

• University Researcher of the Year Award at the Rank of Associate Professor: Florida Atlantic University, 1996-1997.

• Marquis Who’s Who in Science and Engineering.

• Who’s Who in Asian American.